### v1

# Closing the Gap on the Sample Complexity of 1-Identification

Abstract
1-identification is a fundamental multi-armed bandit formulation on pure exploration. An agent aims to determine whether there exists a qualified arm whose mean reward is not less than a known threshold $\mu_{0}$, or to output None if it believes such an arm does not exist. The agent needs to guarantee its output is correct with probability at least $1-\delta$, while making expected total pulling times $\mathbb{E}\tau$ as small as possible. We work on 1-identification with two main contributions. (1) We utilize an optimization formulation to derive a new lower bound of $\mathbb{E}\tau$, when there is at least one qualified arm. (2) We design a new algorithm, deriving tight upper bounds whose gap to lower bounds are up to a polynomial of logarithm factor across all problem instance. Our result complements the analysis of $\mathbb{E}\tau$ when there are multiple qualified arms, which is an open problem left by history literature.
Keywords 1-identification, Fixed Confidence, Pure Exploration
1 Introduction
1.1 Model & Problem Formulation
We study the 1-identification problem [^1] This name origins from Katz-Samuels and Jamieson (2020), Li and Cheung (2025). In Degenne and Koolen (2019), it is also called Any Low Problem., a variant of the best arm identification problem. An instance of 1-identification consists of a tuple of four elements: $\nu=([K],\rho=\{\rho_{a}\}_{a\in[K]},\mu_{0},\delta)$. The set $[K]$ denotes the arm set $\{1,2,\cdots,K\}$, and we assume $K\geq 2$. For each $a\in[K]$, $\rho_{a}$ is the reward probability distribution associated with arm $a$. The reward distribution $\rho_{a}$ is not known to the agent for each $a$. By pulling arm $a$ once, the agent receives a random reward distributed as $\rho_{a}$. We denote $r_{a}^{\nu}:=\mathbb{E}_{R\sim\rho_{a}}[R]$ as the mean reward of arm $a$ under reward distribution $\rho_{a}$. We assume the random noise $R_{a}-r_{a}^{\nu}$ to be 1-sub-Gaussian. The real number $\mu_{0}$ is a threshold known to the agent. The agent’s main goal is to predict whether there is an arm $a\in[K]$ with $r_{a}^{\nu}\geq\mu_{0}$, and to identify such an arm in the positive case, while outputting “None” in the negative case. The quantity $\delta\in(0,1)$ is the tolerant parameter on the probability of a wrong prediction.
The agent’s pulling strategy is specified by a sampling rule $\pi=\{\pi_{t}\}_{t=1}^{\infty}$, a stopping time $\tau$ and a recommendation rule $\hat{a}\in[K]\cup\{\textsf{None}\}$. When a pulling strategy is applied on a 1-identification problem instance with reward distribution $\rho$, the strategy generates a history $\{U_{0}\}\cup\{A_{t},X_{t},U_{t}\}_{t=1}^{\tau}$, where $A_{t}\in[K]$ denotes the arm pulled at $t$, $X_{t}\sim\rho_{A_{t}}$ denotes the reward received, and $U_{t}\sim\text{Unif}(0,1)$ is an internal random seed. Arm $A_{t}\in[K]$ is chosen contingent upon the history $H(t-1)=\{U_{0}\}\cup\{(A_{s},X_{s},U_{s})\}^{t-1}_{s=1}$, via the function $\pi_{t}(H(t-1))$. The agent stops pulling at the end of time step $\tau$, where $\tau$ is a stopping time with respect to the filtration $\{\sigma(H(t))\}^{\infty}_{t=1}$. After time step $\tau$, the agent outputs the prediction $\hat{a}\in[K]\cup\{\textsf{None}\}$. We require that $\hat{a}$ is measurable with respect to $\sigma(H(\tau))$, i.e. determined by the information $H(\tau)$. Outputting $\hat{a}\in[K]$ means that the agent concludes with arm $\hat{a}$ satisfying $r_{\hat{a}}^{\nu}\geq\mu_{0}$, while outputting $\hat{a}=\textsf{None}$ means that the agent concludes with no arm having mean reward $\geq\mu_{0}$. We say that $\nu$ is a positive instance if $\max_{a}r_{a}^{\nu}>\mu_{0}$, and say that $\nu$ is a negative instance if $\max_{a}r_{a}^{\nu}<\mu_{0}$.
For $\Delta>0$, we define $\mathcal{S}^{\text{pos}}_{\Delta}=\{\nu:\max_{a}r_{a}^{\nu}-\mu_{0}\geq\Delta\}$, $\mathcal{S}^{\text{neg}}_{\Delta}=\{\nu:\mu_{0}-\max_{a}r_{a}^{\nu}\geq\Delta\}$, and take $\mathcal{S}^{\text{pos}}=\cup_{\Delta>0}\mathcal{S}^{\text{pos}}_{\Delta},\mathcal{S}^{\text{neg}}=\cup_{\Delta>0}\mathcal{S}^{\text{neg}}_{\Delta}$. We define $i^{*}(\nu)$ as the set of correct answers. For $\nu\in\mathcal{S}^{\text{pos}}_{\Delta}$, we have $i^{*}(\nu)=\{a:r_{a}^{\nu}\geq\mu_{0}\}$, while for instance $\nu\in\mathcal{S}^{\text{neg}}_{\Delta}$, we have $i^{*}(\nu)=\{\textsf{None}\}$. We don’t consider instance whose $\max r_{a}^{\nu}=\mu_{0}$ and mainly focus on analyzing PAC algorithms, with the following definitions.
Definition 1.
A pulling strategy is $\delta$-PAC, if for any $\delta\in(0,1),\nu\in\mathcal{S}^{\text{pos}}\cup\mathcal{S}^{\text{neg}}$, it satisfies $\Pr_{\nu}(\tau<+\infty,\hat{a}\in i^{*}(\nu))>1-\delta$.
Appendix A.3 in Li and Cheung (2025) shows that a $\delta$-PAC algorithm can still suffer from $\mathbb{E}_{\nu}\tau=+\infty$ for some $\nu\in\mathcal{S}^{\text{pos}}$. Thus, we further introduce the following definition, which can be considered “Stronger $\delta$-PAC”.
Definition 2.
A pulling strategy is $(\Delta,\delta)$-PAC, if it is $\delta$-PAC, and for any $\Delta,\delta>0$, we have $\sup_{\nu\in\mathcal{S}^{\text{pos}}_{\Delta}\cup\mathcal{S}^{\text{neg}}_{\Delta}}\mathbb{E}_{\nu}\tau<+\infty$.
[…]
1.2 Lit Review on 1-identification
Before illustrating our main contribution, we first review the weakness in the history literature and we start the review by unifying the notations. In the following paper, we consider a reward vector $\{\mu_{a}\}_{a=1}^{K}\subset[0,1]^{K}$ satisfies that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$. Also, we consider $\mu_{0}\in[0,1]$ and define $\Delta_{i,j}=|\mu_{i}-\mu_{j}|$ for $\forall i,j\in[K]\cup\{0\}$. For an instance $\nu$, we assume $r_{a}^{\nu}=\mu_{a}$ without further explanation. In addition, we introduce different complexity terms occurred in the history result.
$\begin{split}H_{1}^{\text{neg}}=&\sum_{a=1}^{K}\frac{2}{\Delta^{2}_{0,a}},H^{\text{low}}_{1}=\sum_{a:\mu_{a}<\mu_{0}}\frac{2}{\Delta_{1,a}^{2}},H=\frac{2}{\Delta^{2}_{0,1}}\\ H_{1}=&\sum_{a=2}^{K}\frac{2}{\Delta^{2}_{1,a}},H_{0}=\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta^{2}_{0,a}}\\ H^{\text{u}}=&\frac{(\frac{K}{m}-1)}{\Delta^{2}},\Delta:=\min_{a:\mu_{a}>\mu_{0}}\Delta_{a,0},m:=\arg\max\{a:\mu_{a}>\mu_{0}\}\\ H(j)=&\frac{1}{j}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{j,a}^{2},\Delta_{a,0}^{2}\}}\end{split}$ (1)
The discussion on the sample complexity focuses on two cases: instance $\nu$ is negative or positive. There has been abundant research achieving the tight approximation of sample complexity on a negative instance, which is similar to the Best Arm Identification. Regarding the lower bound of $\mathbb{E}\tau$, Theorem 5.4 in Li and Cheung (2025) asserts $\mathbb{E}\tau\geq\Omega(H_{1}^{\text{neg}}\log\frac{1}{\delta})$, if we apply any $\delta$-PAC algorithm on an instance $\nu$. Meanwhile, Algorithm HDoC in Kano et al. (2017), APGAI in Jourdan and Réda (2023), SEE in Li and Cheung (2025), are able to guarantee $\mathbb{E}\tau\leq O(H_{1}^{\text{neg}}(\log\frac{1}{\delta}+\log H_{1}^{\text{neg}}))$, ignoring some logarithm factor. In addition, Algorithm Sticky Track and Stop (S-TaS) in Degenne and Koolen (2019) is able to guarantee $\lim\sup\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\leq T^{*}(\nu)$, where $T^{*}(\nu)=H_{1}^{\text{neg}}$ if $\nu$ is negative and the reward follows Gaussian distribution with variance $1$.
Contrast with the negative instance, the complexity of positive instance remains a mystery, as the proposed upper bounds for a positive instance suffer from different looseness. For simplicity, we consider a positive instance $\nu$ and denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and only consider reward following Gaussian distribution with variance $1$. We propose Table 1 to summarize the looseness in the current result, with imprecise formulation because of space limit. The actual upper bound existing in the history literature is not smaller than our presented formulation, and we can still observe the looseness in our presented formulation. Following the Appendix A.2 in Li and Cheung (2025), we put details of simplification on history upper bound in Appendix A.
Table 1: Comparison of non-asymptotic upper bounds on $\mathbb{E}\tau$ for a positive instance $\nu$. “LB” denotes Lower Bound. $\surd_{\text{cond}}$ means the upper or lower bounds are optimal under some specific instances. The definitions of different $H$’s are at the equation (1). We determine whether a sampling complexity upper bound is nearly optimal by comparing with the lower bound. An upper bound is nearly optimal iff the gap is up to a polynomial of $\log K,\{\log\frac{1}{\Delta_{a,0}^{2}}\}_{a=1}^{K}$.
Algorithm Non-asymptotic Bound Optimality
S-TaS, Degenne and Koolen (2019)
Poiani et al. (2025)
$\mathbb{E}\tau\leq O(\frac{\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\Delta_{0,1}^{4}})$ $\times$
HDoC, Kano et al. (2017) $\mathbb{E}\tau\leq O(\frac{\log\frac{K}{\delta}}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}})$ $\times$
lilHDoC, Tsai et al. (2024) $\mathbb{E}\tau\leq O(\frac{\log\frac{K}{\delta}}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}})$ $\times$
APGAI, Jourdan and Réda (2023) $\mathbb{E}\tau\leq O(H_{0}(\log\frac{K}{\delta}))$ $\times$
FWER-FWPD
Katz-Samuels and Jamieson (2020)
$\mathbb{E}\tau\leq O\Big(H^{\text{u}}\log\frac{1}{\delta}\log\Big(H^{\text{u}}\log\frac{1}{\delta}\Big)\Big)$ $\times$
SEE, Li and Cheung (2025) $\mathbb{E}\tau\leq O(\frac{\log\frac{1}{\delta}}{\Delta_{0,1}^{2}}+H(1)\log\frac{K}{\Delta_{0,1}^{2}})$ $\surd_{\text{cond}}$
Parallel-SEE, This work $\mathbb{E}\tau\leq O\Big(\log^{4}K\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)\log\frac{1}{\Delta_{j,0}^{2}}\Big)$ $\surd$
LB, Katz-Samuels and Jamieson (2020) $\mathbb{E}\tau\geq\Omega(\frac{1}{m}H^{\text{low}}_{1}-\frac{1}{\Delta^{2}_{1,m+1}})$ $\times$
LB, Li and Cheung (2025) $\mathbb{E}\tau\geq\Omega(H\log\frac{1}{\delta}+\frac{1}{m}H^{\text{low}}_{1}-\frac{1}{\Delta^{2}_{1,m+1}})$ $\surd_{\text{cond}}$
LB, This work $\mathbb{E}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$ $\surd$
Degenne and Koolen (2019) propose algorithm S-TaS, which provides asymptotic guarantee for both negative and positive instances. S-TaS is able to guarantee $\lim\sup\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\leq\frac{2}{\Delta_{0,1}^{2}}$. Together with their proposed lower bound $\lim\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\geq\frac{2}{\Delta_{0,1}^{2}}$, we know S-TaS is asymptotically optimal. Degenne and Koolen (2019) leaves non-asymptotic analysis of S-TaS as a future direction, which is complemented by Poiani et al. (2025). Poiani et al. (2025) present the non-asymptotic upper bound in an implicit way, but we are able to derive a smaller but explicit formulation. The weakness of Poiani et al. (2025) comes from the $\delta$-independent part $\frac{K^{2}}{\Delta_{0,1}^{4}}$, which is inconsistent with any lower bound.
Algorithm HDoC in Kano et al. (2017) proves an upper bound not smaller than $C^{\text{HDoC}}_{\text{pos}}:=\frac{\log(K/\delta)}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}}$, where $\epsilon=\min\Big\{\min\limits_{a\in[K]}\Delta_{0,a},\min\limits_{a\in[K-1]}\frac{\Delta_{a,a+1}}{2}\Big\}$. The suboptimality of $C^{\text{HDoC}}_{\text{pos}}$ consists of two parts. The first part is $H_{1}\log\log\frac{1}{\delta}$, with comments in Kano et al. (2017): “Therefore, in the case that $\ldots$ $K=\Omega(\frac{\log\frac{1}{\delta}}{\log\log\frac{1}{\delta}})$there still exists a gap between the lower bound in …and the upper bound in …”. The second part is $\frac{K}{\epsilon^{2}}$, which can be infinity if $\mu_{i}=\mu_{j}$ holds. for some $i,j\in[K]$. Based on HDoC, Tsai et al. (2024) propose Algorithm lilHDoC, which incorporate the LIL concentration event (Lemma 3 in Jamieson et al. (2014)) and add an extra warm-up stage. After adding these designs, the proposed algorithm lilHDoC still suffers from the same sub-optimality as HDoC.
Kano et al. (2017) design Algorithm APGAI to address Good Arm Identification, aiming to output all the arms whose reward is above a given threshold $\mu_{0}$. Though it is not the main focus, they also propose an upper bound for 1-identification. The weakness of upper bound $O(H_{0}(\log\frac{K}{\delta}))$ mainly comes from $H_{0}$. Easy to see $H_{0}$ is proportional to $K$, suggesting that the coefficient of $\log\frac{1}{\delta}$ is suboptimal compared to the others. What’s more, if there exists an arm $a$ such that $\mu_{a}=\mu_{0}$, we have $H_{0}=\infty$, suggesting that the upper bound is vacuous.
Katz-Samuels and Jamieson (2020) work on $\epsilon$-good identification and 1-identification simultaneously. They define a new concept Unverifiable Sample Complexity and prove upper bounds for it. Though Unverifiable Sample Complexity is different with Sample Complexity in this paper, we can still adapt their result, as discussed by Appendix A.1 in Li and Cheung (2025). The suboptimality of its upper bound consists of two points. First, its upper bound relies on $\log\frac{1}{\delta}\log\log\frac{1}{\delta}$ instead of $\log\frac{1}{\delta}$. Second, the complexity term $H^{u}$ relies on the minimum gap between a qualified mean reward and the threshold $\mu_{0}$, which can be much smaller than $\Delta_{0,1}$. In the case $m$ is much smaller than $K$ and $\Delta<\Delta_{0,1}$, this upper bound is much larger than their proposed lower bound.
Li and Cheung (2025) design Algorithm SEE, splitting the upper bound of $\mathbb{E}\tau$ into two parts which are $\delta$-dependent and $\delta$-independent. They prove both upper and lower bounds of $\mathbb{E}\tau$ with explicit formulation. They show when $m=1$, i.e. there is only one arm whose reward is greater than $\mu_{0}$, the gap between their proposed upper and lower bounds are up to a polynomial of $\log K,\log\frac{1}{\Delta_{0,1}^{2}}$, suggesting the optimality. However, they also comment $H(1)\log\frac{K}{\Delta_{0,1}^{2}}$, which is $\delta$-independent part, might be improved if $m>1$. They leave the possible improvement as an open direction.
1.3 Main Contribution
In this paper, we develop novel upper and lower bounds for the 1-identification problem. Regarding the lower bound, we conclude
Theorem 1.
For any $\delta$-PAC alg and any reward vector $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, any tolerance level $\delta$ such that $\delta<\min\{\frac{1}{10^{-8}},\frac{1}{64m^{2}}\}$, we can find a permutation $\sigma:[K]\rightarrow[K]$ and define a positive instance $\nu$ with $r_{\sigma(a)}^{\nu}=\mu_{a},a\in[K]$ with unit-variance Gaussian reward, such that $\mathbb{E}_{\text{alg},\nu}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$.
Regarding the upper bound, we merge and adapt the idea in Katz-Samuels and Jamieson (2020) and Li and Cheung (2025) and design a new algorithm Parallel Sequential Exploration Exploitation on Brackets(Algorithm 1, PSEEB). We can prove the following two theorems.
Theorem 2.
Algorithm 1 is $\delta$-PAC.
Theorem 3.
Apply Algorithm 1 to a 1-identification instance $\nu$, we have
$\mathbb{E}_{\nu}\tau\leq\begin{cases}O\left(\log K\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\log^{3}K(\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}})H(j)\right)&\nu\text{ is positive}\\ O\left((\log K)H_{1}^{\text{neg}}\log\frac{H_{1}^{\text{neg}}}{\delta}\right)&\nu\text{ is negative}\end{cases}$
combining Theorem 2 and 3, we know algorithm 1 is $(\Delta,\delta)$-PAC. What’s more, the gap between Theorem 1 and 3 is up to a polynomial of $(\log K,\log\{\frac{1}{\Delta_{0,a}^{2}}\}_{a=1}^{m})$, suggesting that our proposed algorithm 1 is nearly optimal for all 1-identification instances, and $\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)$ is an appropriate complexity measure. Compared to history literature, our proposed Algorithm 1 complement one of the open direction left in Li and Cheung (2025), by deriving a new upper bound for $\mathbb{E}\tau$ when there are multiple qualified arms.
1.4 Literature Review on Other Related Works
If we consider the given threshold $\mu_{0}$ as an arm with deterministic reward, we can address 1-identification by solving a Best Arm Identification(BAI) problem Even-Dar et al. (2002), Gabillon et al. (2012), Jamieson et al. (2013), Kalyanakrishnan et al. (2012), Karnin et al. (2013), Jamieson and Nowak (2014), Chen and Li (2015). However, this is an inefficient solution, suggested by the asymptotic result in Degenne and Koolen (2019), Garivier and Kaufmann (2016) and non-asymptotic result in Li and Cheung (2025), Kano et al. (2017). The reason is 1-identification doesn’t care about an exact comparison between arms with high confidence level, and we should not pull arms as frequently as BAI to distinguish two arms.
Besides 1-identification, there are other variants of pure exploration with multiple answers, as mentioned by Degenne and Koolen (2019). The agent needs to find all the arms Mason et al. (2020) in $\{a:\mu_{a}\geq\mu_{1}-\epsilon\}$ or only one arm Even-Dar et al. (2002), Gabillon et al. (2012), Kalyanakrishnan et al. (2012), Kaufmann and Kalyanakrishnan (2013), Katz-Samuels and Jamieson (2020). As discussed in Katz-Samuels and Jamieson (2020), if the gap $\mu_{1},\mu_{0}$ are known and $\mu_{1}>\mu_{0}$, by taking $\epsilon=\Delta_{0,1}$, the answer set of $\epsilon$-Good Arm Identification and 1-identification are the same. We can adopt same pulling strategy and recommendation rule. But it is unclear about the relationship if there are unknown. What’s more, Locatelli et al. (2016), Mukherjee et al. (2017) work on Thresholding Bandits, which aims to output a set of arms whose reward are above a given $\mu_{0}$, within fixed budget. The optimal set might not be unique under their proposed problem setting.
In this paper, we adopt the idea of parallel algorithm copies, which also exists in the history literature. Chen and Li (2015), Chen et al. (2017), Katz-Samuels and Jamieson (2020) all require a main algorithm to manage different algorithm copies. Katz-Samuels and Jamieson (2020) allocate $2^{\ell}$ time rounds to a copy each time with increasing index $\ell$. Chen and Li (2015), Chen et al. (2017) maintain an overall round index $t\in\mathbb{N}$ and further assign algorithm index to each copy. They execute an algorithm copy only when its algorithm index divide $t$. In our work, we apply round-robin procedure to execute each algorithm copy.
2 Lower Bound of Sample Complexity, for a Positive Instance
In this section, we continue to work on $\{\mu_{a}\}_{a=1}^{K}$ such that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$. We assume each reward distribution follows a unit-variance Gaussian distribution in our lower bound analysis. For an instance $\nu^{\prime}$ and a permutation $\sigma:[K]\rightarrow[K]$, represented as a bijection from $[K]$ to $[K]$, we define the instance $\sigma(\nu^{\prime})$ as $r^{\sigma(\nu^{\prime})}_{\sigma(a)}=r_{a}^{\nu^{\prime}}$.
We first introduce the definition of a symmetric algorithm.
Definition 3.
An algorithm $\widetilde{\text{alg}}$ is said to be symmetric if the following statement holds. For any instance $\nu$ with reward vector $\{\mu_{a}\}_{a=1}^{K}$, any sequence $\{a_{t}\}_{t=1}^{T}\in[K]^{T}$, natural number $T\geq 1$, and two different permutations $\sigma^{\prime},\sigma^{\prime\prime}$, the algorithm $\widetilde{\text{alg}}$ satisfies
$\Pr_{\sigma^{\prime}(\nu),\tilde{\text{alg}}}(\forall t\in[T],A_{t}=\sigma^{\prime}(a_{t}),\tau>T)=\Pr_{\sigma^{\prime\prime}(\nu),\tilde{\text{alg}}}(\forall t\in[T],A_{t}=\sigma^{\prime\prime}(a_{t}),\tau>T).$
Intuitively, under a symmetric algorithm $\widetilde{\text{alg}}$, the reward distribution of the pulled arm $A_{t}$ at time $t$ is invariant under any permutation on the arms’ indexes. In particular, for a symmetric algorithm $\widetilde{\text{alg}}$, we have $\mathbb{E}_{\sigma^{\prime}(\nu),\tilde{\text{alg}}}\tau=\mathbb{E}_{\sigma^{\prime\prime}(\nu),\tilde{\text{alg}}}\tau$ for any two permutations $\sigma^{\prime},\sigma^{\prime\prime}$. A similar definition also appears in Definition 12 of Simchowitz et al. (2017), which further requires the distribution of output $\hat{a}$ is also invariant regarding the distribution. This extra requirement is unnecessary in our analysis. By the following lemma, it suffices to consider only symmetric algorithms for establishing our lower bound.
Lemma 4.
Denote $S$ as the set of all permutations of $[K]$. Given any $\delta$-PAC algorithm alg, we can always find a $\delta$-PAC symmetric algorithm $\widetilde{\text{alg}}$, such that for any instance $\nu^{\prime}$, we have $\max\limits_{\sigma\in S}\mathbb{E}_{\text{alg},\sigma(\nu^{\prime})}\tau\geq\mathbb{E}_{\tilde{\text{alg}},\nu^{\prime}}\tau$.
Full proof of Lemma 4 is in Appendix B.1. We can shuffle all arms before applying any alg to create a symmetric algorithm $\widetilde{\text{alg}}$. Given Lemma 4, to prove Theorem 1, we suffice to prove the following Theorem.
Theorem 5.
Suppose $\delta<\min\{10^{-8},1/(64m^{2})\}$. For any $\delta$-PAC symmetric algorithm, and for any positive instance $\nu$ with $m\geq 1$ eligible arms, we have
$\mathbb{E}\tau\geq\Omega\left(\min_{1\leq j\leq m}\frac{\log(1/\delta)}{\Delta_{j,0}^{2}}+\frac{1/j}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$
The assumption $\delta<\min\{10^{-8},1/(64m^{2})\}$ is to ensure $\delta\cdot m^{2}$ is upper bounded by a non-trivial absolute constant. The upper bound eliminates the possibility of being $\delta$-PAC by outputting a uniformly chosen arm without any arm pulls.
Sketch Proof of Theorem 5.
In what follows, we present the main steps to prove Theorem 5. The full proof is in Appendix B.2, which consists of the statements and proofs of auxiliary lemmas. The main idea is to establish Lemma 6,
Lemma 6.
Assume $\delta<10^{-8}$, $\frac{1}{8m}\geq\sqrt{\delta}$, we can assert $\mathbb{E}_{\nu}\tau$ is lower bounded by the $\frac{1}{3200}$ times optimal value of the following optimization problem
$\min_{v,\{p_{j}\}_{j=1}^{m}}\quad$ $v$ (2)
s.t. $0\leq p_{j}\leq 1\quad\forall j\in[m]$ (C1)
$v\geq 0$ (C2)
$v\geq\sum_{j=1}^{m}\frac{p_{j}\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}$ (C3)
$\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}$ (C4)
$v\geq\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}\quad\forall j\in[m].$ (C5)
If Lemma 6 holds, we first prove the above optimization problem (2) is convex. Then, we turn to construct a feasible solution to the dual problem, and prove the objective value of this feasible solution is lower bounded by $\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$. By the strong duality, we complete the proof of Theorem 5.
Denote event $\mathcal{E}_{j}=\left\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right\}$. To prove Lemma 6, we are going to show taking $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ for $j\in[m]$ is a feasible solution to (2). The overall idea is to apply the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)) to prove the relationship between $\Pr_{\nu}(\mathcal{E}_{j})$ and $\Pr_{\nu^{\prime}}(\mathcal{E}_{j})$ for some alternative instances $\nu^{\prime}$, further the relationship between $3200\mathbb{E}_{\nu}\tau$ and $\{\Pr_{\nu}(\mathcal{E}_{j})\}_{j=1}^{m}$. We prove the feasibility of $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ by going through the constraint (C1), (C2), (C3), (C4) (C5).
For (C1), (C2), they are trivial. As the expected pulling times must be non-negative and the probability must be in the interval $[0,1]$.
For (C3), notice that $\mathds{1}(\tau=+\infty)+\mathds{1}(\tau<+\infty,\hat{a}\in[K]\cup\{\textsf{None}\})=1$, we have $\sum_{j=1}^{m}\mathds{1}(\hat{a}=j)\leq 1$, further $\sum_{j=1}^{m}\mathds{1}(\mathcal{E}_{j})\leq 1$. This observation concludes
$3200\mathbb{E}\tau\geq$ $\sum_{j=1}^{m}\mathbb{E}\tau\mathds{1}(\mathcal{E}_{j})\geq\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr_{\nu}(\mathcal{E}_{j}).$
For (C4), we start by proving the inequality (3).
$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])\geq 1-5\delta.$ (3)
To prove (3), we consider two cases, $\{a:\mu_{a}=\mu_{0}\}$ is empty or not. The $\delta$-PAC requirement suggests that (3) must hold for both cases.
Then, we are ready to prove (C4) by contradiction. Denote $\bar{\mathcal{E}_{j}}=\{N_{j_{0}}(\tau)<\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}},\hat{a}=j_{0}\}$. If $\sum_{j=1}^{m}\Pr_{\nu}(\mathcal{E}_{j})<\frac{1}{2}$, together with (3), we can find a $j_{0}\in[m]$ such that
$\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)\geq\frac{1}{8m}\geq\sqrt{\delta}.$ (4)
We consider an alternative instance $\nu^{\prime}_{j_{0}}(\epsilon)$ with $r_{a^{\prime}}^{\nu^{\prime}_{j_{0}}}=\begin{cases}\mu_{0}-\epsilon&a^{\prime}=j_{0}\\ r_{a^{\prime}}^{\nu}&else\end{cases}$. The $\delta$-PAC requirement of algorithm conclude
$\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})<\delta$ (5)
Combining (5),(4), we have $\log\frac{\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)}{\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})}\geq\frac{1}{2}\log\frac{1}{\delta}$ holds for all $\epsilon>0$. Meanwhile, Lemma 8, which utilizes the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)), suggests that
$\log\frac{\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)}{\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})}$
$\leq$ $\sqrt{2\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}\log\frac{1}{\sqrt{\delta}}}+\sqrt{2\pi\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}}+\frac{1}{2}\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}$ (6)
Taking $\epsilon\rightarrow 0$, we can show the right hand side of (2) is less than $\frac{1}{2}\log\frac{1}{\delta}$ for $\delta<10^{-8}$, which leads to a contradiction. Lemma 9 merges all the above ideas and conclude $\sum_{j=1}^{m}\Pr_{\nu}(\mathcal{E}_{j})\geq\frac{1}{2}$.
For (C5), we first prove Lemma 11, asserting that
$\mathbb{E}_{\nu}N_{a}(\tau)\geq\frac{1}{4}T_{j}(a)\Pr_{\nu}\left(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right).$ (7)
for each $j\in[m],a\in[K]$, where $T_{j}(a)=\lceil\frac{1}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\Big(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}\Big)}\rceil$. If Lemma 11 holds, we can sum up (7) for each $a\in[K]$ and use the fact that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$, concluding (C5) holds. Then, the remaining work is to prove (7). In the following, we present the sketch idea, which highly relies on the definition of $T_{j}(a)$.
For $n\in\mathbb{N}$, denote
$S_{j,a}^{n}=$ $\{t:N_{j}(t)+N_{a}(t)=n,A_{t+1}\in\{j,a\}\}$
$\tau_{n}^{j}(a)=$ $\min S_{j,a}^{n}$
For any sample path $\{U_{s},A_{s},X_{s}\}_{s=1}^{\tau}$, $S_{j,a}^{n}$ contains at most one element, which is the round that an agent is going to pull arm $a$ or $j$. We suffice to focus on these rounds. In the case that $S_{j,a}^{n}=\emptyset$, we take $\tau_{n}^{j}(a)=+\infty$, which means $\tau_{n}^{j}(a)\geq\tau$. Since $A_{t+1}$ only depends history $H_{t}$, we know $\tau_{n}^{j}(a)$ is a stopping time. Regarding $\tau_{n}^{j}(a)$, we are going to prove two inequalities
$\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)\geq\Pr_{\nu}(\mathcal{E}_{j}),\forall n\leq T_{j}(a).$ (8)
$\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau)\geq\frac{1}{4}\Pr_{\nu}(\tau_{n}^{j}(a)<\tau),\forall n\leq T_{j}(a).$ (9)
If (8) (9) both hold, from the fact that we only pull arm a at $\tau_{n}^{j}(a)+1$, we can conclude
$\mathbb{E}_{\nu}N_{a}(\tau)\geq$ $\sum_{n=0}^{T_{j}(a)}\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<+\infty)$
$\geq$ $\sum_{n=0}^{T_{j}(a)}\frac{1}{4}\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$
$\geq$ $\frac{1}{4}\sum_{n=0}^{T_{j}(a)}\Pr_{\nu}(\mathcal{E}_{j})$
$=$ $\frac{1}{4}T_{j}(a)\Pr_{\nu}(\mathcal{E}_{j}).$
The second step is by (9) and the third step is by (8). Thus, to prove (7), we suffice to show (8) (9) are both correct.
For (8), we notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,0}^{2}}$, further $T_{j}(a)<\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}$. That means for $n\leq T_{j}(a)$, we have $\mathcal{E}_{j}\Rightarrow\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Rightarrow\{\tau_{n}^{j}(a)<\tau\}$, which suggests (8).
For (9), we introduce an alternative instance $\nu_{j,a}$, with $r_{a^{\prime}}^{\nu_{j,a}}=\begin{cases}r_{a}^{\nu}&a^{\prime}=j\\ r_{j}^{\nu}&a^{\prime}=a\\ r_{a^{\prime}}^{\nu}&else\end{cases}$. $\nu_{j,a}$ swaps the position of arm $a$ and $j$ compared to the instance $\nu$. Denote $\tilde{\mathcal{E}}^{aj}(n)=\{\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{j}(n)=\{A_{\tau_{n}^{j}(a)+1}=j,\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{a}(n)=\{A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau\}$. Notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,a}^{2}}$, further $T_{j}(a)\leq\frac{O(1)}{\Delta_{j,a}^{2}(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})})}$, we can use Lemma 7 and 10, which apply the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)), to conclude
$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)$ (10)
where $\text{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y}$. To complete this step, we have to put $1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}$ in the denominator of $T_{j}(a)$. (10) suggests that the algorithm is unable to distinguish $\nu$ and $\nu_{ja}$ at the round $t$ such that $N_{a}(t)+N_{j}(t)\leq T_{j}(a)$. For a symmetric algorithm, we have $\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))$, which means
$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)$ (11)
By the Pinsker’s Inequality $kl(x,y)\geq 2(x-y)^{2}$ and $\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j})+\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a})=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj})$, we can derive
$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\geq\frac{1}{4}\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a).$
The above inequality suggests that we we have proved (9), further the correctness of (C5). ∎
3 Upper Bound on the Complexity of 1-identification
In this section, we present our proposed Algorithm PSEEB with rough analysis.
3.1 Algorithm
1Input: Threshold $\mu_{0}$, Error tolerance level $\delta$, Arm number $K$;
2
Randomly generate a permutation of $\sigma:[K]\rightarrow[K]$, take $B_{b}=\{\sigma(i)\}_{i=1}^{\min\{2^{b-1},K\}}$ for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$;
// $\sigma(i)$ denotes the arm index in position $i$
3
4Create algorithm copies $\{\text{alg}_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ by specifying $\text{alg}_{b}$ through calling Algorithm 2 with $B\leftarrow B_{b}$, $\mu_{0}\leftarrow\mu_{0}$, $\delta\leftarrow\frac{\delta}{\lceil\log_{2}K\rceil+1}$, $C\leftarrow 1.01$, $K\leftarrow K$. Samples across bracket are not shared;
5 for round index $t=1,2,\cdots$ do
6 for Bracket Index $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$ do
7      Execute $\text{alg}_{b}$ for one time round;
8    if $\text{alg}_{b}$ terminates with output $\hat{a}\in[K]\cup\{\textsf{None}\}$ then
9         Terminate the whole algorithm and return $\hat{a}$.
10      end if
11
12   end for
13
14 end for
Algorithm 1 Parallel Sequential Exploration Exploitation on Brackets(PSEEB)
We propose Algorithm 1 to solve the 1-identification problem, which merges and adapts the idea in Katz-Samuels and Jamieson (2020) and Li and Cheung (2025). Katz-Samuels and Jamieson (2020) create each bracket by a unique and independent random shuffle, and a corresponding algorithm copy to pull arms in this bracket. In contrast to them, we specify all the brackets from a single random permutation. Algorithm 1 generates a random permutation of $[K]$ before the start of the pulling process. Based on the random permutation, Algorithm 1 create $\lceil\log_{2}K\rceil+1$ brackets $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$, satisfying $B_{1}\subset B_{2}\subset\cdots\subset B_{\lceil\log_{2}K\rceil+1}=[K]$. For a bracket index $1\leq b\leq\lceil\log_{2}K\rceil$, we have $|B_{b}|=2^{b}$, and $|B_{\lceil\log_{2}K\rceil+1}|=\min\{2^{\lceil\log_{2}K\rceil},K\}$. Our approach is to guarantee $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ follow an increasing order, which is crucial to our Lemma 31. Lemma 31 asserts that with at least some probability, the first several brackets contain at least 1 qualified arm, if there are multiple qualified arms.
Then, Algorithm 1 calls Algorithm 2, defined in Appendix C, as an oracle to set up $\lceil\log_{2}K\rceil+1$ algorithm copies. Algorithm 2 is adapted from the Algorithm 2,3,4,5 in Li and Cheung (2025), named as Sequential Exploration Exploitation(SEE). Algorithm copy $b\in[\lceil\log_{2}K\rceil+1]$ takes bracket $B_{b}$ as an input parameter and only pull arms in $B_{b}$. To fit in the bracketing idea, we modify its budget allocation and stopping rules. Detailed discussion of Algorithm 2 is in Appendix C.
Algorithm 1 manages all the algorithm copies by a round-robin manner. In each round, Algorithm 1 execute one of the algorithm copies with one arm pull and check whether the copy terminates with an output $\hat{a}\in[K]\cup\{\textsf{None}\}$. By the design of Algorithm 2, $\text{alg}_{\lceil\log_{2}K\rceil+1}$ might terminate with taking output $\hat{a}\in[K]\cup\{\textsf{None}\}$. While for bracket $b$ whose size is smaller than $K$, $\text{alg}_{b}$ cannot output None but keep pulling arms, even if the $\text{alg}_{b}$ finds that $\max_{a\in B_{b}}\mu_{a}<\mu_{0}$. The output strategy differs with the algorithm SEE in Li and Cheung (2025).
Once there exists an algorithm copy terminates with an output, Algorithm 1 will adopt the output as the final output and terminate the overall sampling process. We require each algorithm copy to pull arms in a parallel manner without sharing samples. Denote $\tau$ as the total length of algorithm 1, and $\tau(B_{b})$ as the total length of copy $\text{alg}_{b}$. We allow $\tau,\tau(B_{b})$ to take infinities. Since any termination of algorithm copy results in the termination of Algorithm 1, we have $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$, further $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{b})$.
Generally speaking, for a positive instance $\nu$ with $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\mu_{K}$, we wish a bracket $B$ with size smaller than $K$ contains a qualified arm $j\in[m]$. Then, we wish the algorithm copy on this bracket $B$ is able to spend less pulling times to output arm $j$. While for a negative instance $\nu$ with $\mu_{0}>\mu_{1}\geq\mu_{2}\geq\cdots\mu_{K}$, it is possible that all the algorithm copies whose index $b\leq\lceil\log_{2}K\rceil$ never terminates. We only wish the copy $\lceil\log_{2}K\rceil+1$ is able to terminate with finite sampling and output $\hat{a}=\textsf{None}$ when encountering a negative instance.
3.2 Performance Guarantee of Algorithm 1
In the following of this section, we present the sketch proof of Theorem 2 and 3. We use Figure 1 in Appendix C.3 to summarize the overall steps.
Sketch Proof of Theorem 2.
The first step is prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)=0$, whose main idea is very similar to the proof of Theorem 5.2 in Li and Cheung (2025). Same as Li and Cheung (2025), denote $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ as the minimum phase index such that the concentration events of copy $b$ hold, we can prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)\leq\Pr(\max\{\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1},\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}\}=+\infty)=0$.
Given the first step, we are able to conclude $\Pr\Big(\tau<+\infty\Big)=1$, by the fact that $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$. Then, to prove Algorithm 1 is $\delta$-PAC, we suffice to show $\Pr_{\nu}\left(\hat{a}=\textsf{None}\text{ or }\{\hat{a}\in[K],\mu_{\hat{a}}<\mu_{0}\}\right)<\delta$ for a positive instance $\nu$, and $\Pr_{\nu}\left(\hat{a}\in[K]\right)<\delta$ for a negative instance $\nu$.
The proof for both targets is similar. Recall we deliver $\frac{\delta}{\lceil\log_{2}K\rceil+1}$ to each algorithm copy in the Line 1 o Algorithm 1. We are able to prove $\Pr(\text{alg}_{b}\text{ outputs incorrect }\hat{a})<\frac{O(1)\delta}{\lceil\log_{2}K\rceil+1}$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$. Taking the union bound, we complete the proof. ∎
Sketch Proof of Theorem 3.
We split the proof into two cases, which are positive and negative $\nu$.
If the instance $\nu$ is negative, we focus on analyzing $\tau(B_{\lceil\log_{2}K\rceil+1})$, which is very similar to the negative case of Theorem 5.3 in Li and Cheung (2025). By the fact that $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{\lceil\log_{2}K\rceil+1})$, we complete the proof.
If the instance $\nu$ is positive, denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$. Denote $\tilde{T}_{j}(a)=\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, $\tilde{T}_{j}(a;k)=\frac{\log\frac{4K}{\delta_{k}}+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, we are going to show $\mathbb{E}\tau\leq O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\tilde{T}_{j}(a)\right)$ holds for all $j\in[m]$. What’s more, the constant hidden by the Big O notation is independent of $j$. For any given $j\in[m]$, we introduce $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$, which is the minimum bracket index that contains at least one arm not worse than $\mu_{j}$. As we adopt a random permutation to create brackets, $b_{j}$ is a random variable, and we have $\mathbb{E}\tau=\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\mathbb{E}\tau\mathds{1}(b_{j}=b)$. Then, we turn to analyze $O(\log K)\mathbb{E}\tau(B_{b})\mathds{1}(b_{j}=b)$, which is the upper bound of $\mathbb{E}\tau\mathds{1}(b_{j}=b)$.
Denote $\tau^{\text{ee}}(B_{b}),\tau^{\text{et}}(B_{b})$ as the total length of exploration and exploitation periods of algorithm copy $\text{alg}_{b}$ separately. For any arm $a\in B_{b}$, denote $N_{a}^{\text{ee},B_{b}}$ as the pulling times of arm $a$ in the exploration period. Not hard to see $\tau(B_{b})=\tau^{\text{ee}}(B_{b})+\tau^{\text{et}}(B_{b})$. We are going to prove
$\mathbb{E}\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)\Pr(b_{j}=b),$ (12)
$\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)\leq\sum_{a=1}^{K}O(1)\tilde{T}_{j}(a)\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}+O(\log K)|B_{b}|\Pr(b_{j}=b).$ (13)
If (12)(13) are both correct, we can sum up (12)(13) for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$. Lemma 31, 33, 34 prove upper bounds for $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}|B_{b}|\Pr(b_{j}=b)$ and $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}$, which complete the proof of Theorem 3.
To prove (12), we first show the algorithm must terminate no later than the end of phase $L_{b}^{\text{pos}}=\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{O(1)(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$, if $b_{j}=b$. Notice that $\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}=O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$, and Lemma 17 shows $\Pr(\kappa^{\text{ee}}_{b}\geq k)\leq\frac{\pi^{2}\delta_{k-1}}{6}$, $\Pr(\kappa^{\text{et}}_{b}\geq k)\leq\frac{\pi^{2}\delta}{6\alpha_{k-1}}$, we can take expectation on both sides of $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}\mathds{1}(b_{j}=b)$ to complete the proof of (12).
To prove (13), we first show for each $a\in B_{b}$, we have
$N_{a}^{\text{ee},B_{b}}\mathds{1}(b_{j}=b)\leq O(\tilde{T}_{j}(a;L_{b}^{\text{pos}}))(C+1)^{2}\mathds{1}(b_{j}=b).$ (14)
The intuition is as follows. Given the fact that $j\in B_{b}$, once the pulling times of arm $a$ is higher than $O(\tilde{T}_{j}(a;L_{b}^{\text{pos}}))(C+1)^{2}$, copy $\text{alg}_{b}$ can assert either $\text{UCB}_{a}<\mu_{j}$ or $\text{LCB}_{a}>\mu_{0}$, with a constant confidence level. In both cases, $\text{alg}_{b}$ stops pulling arm $a$ in the exploration period. By the facts that $\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)=(|Q|+\sum_{a\in B_{b}}N_{a}^{\text{ee},B_{b}})\mathds{1}(b_{j}=b)$ and elements in $B_{b_{j}}$ are independent of $\kappa^{\text{ee}}_{b_{j}},\kappa^{\text{et}}_{b_{j}}$, we can sum up (14) for $a\in B_{b}$ and derive
$\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)\leq\sum_{a=1}^{K}O(1)\tilde{T}_{j}(a)\Pr(a\in B_{b},b_{j}=b)+O(\log K)|B_{b}|\Pr(b_{j}=b).$
Since $\Pr(a\in B_{b},b_{j}=b)\leq\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}$, and $\Pr(a\in B_{b})=\frac{|B_{b}|}{K}$, we can conclude (13). ∎
Appendix A Detailed Literature Review
A.1 Simplification of History Result
In this section, we only consider Gaussian reward with unit variance. We discuss more details of the upper bounds presented in Table 1. We simplify the result in the history literature with smaller but explicit formulations, for the comparison in Table 1.
A.1.1 Theorem 2 in Poiani et al. (2025)
Theorem 2 in Poiani et al. (2025) presents the Non-asymptotic analysis for algorithm S-TaS. For an instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$, $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and , they define $\neg a=\{\nu^{\prime}:r_{a}^{\nu^{\prime}}<\mu_{0}\}$ for $a\in[m]$, and
$i_{F}(\nu)=\arg\max_{a\in[K]}\sup_{w\in\Delta_{K}}\inf_{\nu^{\prime}\in\neg a}\sum_{a^{\prime}\in K}\frac{w_{a^{\prime}}(r_{a^{\prime}}^{\nu^{\prime}}-r_{a^{\prime}}^{\nu})^{2}}{2}.$
Then, Theorem 2 in Poiani et al. (2025) considers $\epsilon_{\nu}$ such that for all instance $\nu^{\prime}$ such that
$\|\{r_{a}^{\nu^{\prime}}\}_{a=1}^{K}-\{r_{a}^{\nu}\}_{a=1}^{K}\|_{\infty}\leq\epsilon_{\nu}\implies i_{F}(\nu^{\prime})\subset i_{F}(\nu)\cup([K]\setminus i^{*}(\nu)).$
Here $i^{*}(\nu)=[m]$. Let
$T_{\nu}=\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\sqrt{\frac{8D_{K}\log n}{\sqrt{\sqrt{n}+K^{2}}-2K}}\leq\epsilon_{\nu}\right\}\right\}$
where $D_{K}\geq 10$. The upper bound for $\mathbb{E}_{\nu}\tau$ is
$\mathbb{E}_{\nu}\tau\leq 10K^{4}+T_{0}(\delta)+\frac{\pi^{2}}{24}$
where
$T_{0}(\delta)=\inf\left\{t\in\mathbb{N}:\log(Ct^{2}/\delta)\leq(t-\sqrt{t}-1-T_{\nu})\frac{\Delta_{0,1}^{2}}{2}-g(t)\right\}$
and $C\geq e\sum_{t=1}^{+\infty}(\frac{e}{K})^{K}\frac{(\log^{2}(Ct^{2})\log(t))^{K}}{t^{2}}$, $g(t)=\tilde{O}(K^{2}\sqrt{D_{K}t}+\sqrt{D_{K}t^{\frac{3}{2}}})$.
To simplify the above result, we can prove $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\Delta_{0,1}^{4}}$ is a lower bound of $10K^{4}+T_{0}(\delta)+\frac{\pi^{2}}{24}$. From the definition of $T_{\nu}$, we know
$T_{\nu}\geq$ $\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\sqrt{\frac{8D_{K}\log n}{\sqrt{\sqrt{n}+K^{2}}-2K}}\leq\epsilon_{\nu}\right\}\right\}$
$\geq$ $\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\sqrt{\frac{8D_{K}}{\sqrt{\sqrt{n}+K^{2}}-2K}}\leq\epsilon_{\nu}\right\}\right\}$
$=$ $\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\frac{80}{\epsilon_{\nu}^{2}}+2K\leq\sqrt{\sqrt{n}+K^{2}}\right\}\right\}$
$\geq$ $\max\left\{10K^{4},\inf\left\{n\in\mathbb{N}:\frac{320K}{\epsilon_{\nu}^{2}}+4K^{2}\leq\sqrt{n}+K^{2}\right\}\right\}$
$\geq$ $\max\left\{10K^{4},\frac{90000K^{2}}{\epsilon_{\nu}^{4}}\right\}.$
We can further derive
$T_{0}(\delta)\geq$ $\inf\left\{t\in\mathbb{N}:\log(C/\delta)\leq\left(t-\max\left\{10K^{4},\frac{90000K^{2}}{\epsilon_{\nu}^{4}}\right\}\right)\frac{\Delta_{0,1}^{2}}{2}\right\}$
$=$ $\frac{2\log(C/\delta)}{\Delta_{0,1}^{2}}+\max\left\{10K^{4},\frac{90000K^{2}}{\epsilon_{\nu}^{4}}\right\}.$
Thus, we can conclude $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\epsilon_{\nu}^{4}}$ is smaller than the proposed upper bound in Poiani et al. (2025). The remaining work is to show $\epsilon_{\nu}\leq\Delta_{0,1}$. Denote $a^{\text{sub}}=\arg\max_{a:\mu_{a}<\mu_{1}}\mu_{a}$. We have
  • If $\mu_{a^{\text{sub}}}\geq\mu_{0}$, we assert $\epsilon_{\nu}\leq\Delta_{1,a^{\text{sub}}}<\Delta_{1,0}$. If not, we can set $\nu^{\prime}$ with $r_{a^{\text{sub}}}^{\nu^{\prime}}=\mu_{1}$ while keeping $r_{a}^{\nu^{\prime}}=r_{a}^{\nu^{\prime}}$ for $a\neq a^{\text{sub}}$. With this construction, we have $a^{\text{sub}}\in i_{F}(\nu^{\prime}),a^{\text{sub}}\notin\{1,m+1,m+2,\cdots,K\}$, contradicting to the requirement of $\epsilon_{\nu}$.
  • If $\mu_{a^{\text{sub}}}<\mu_{0}$, we assert $\epsilon_{\nu}\leq\Delta_{1,0}$. If $\epsilon_{\nu}>\Delta_{1,0}$, we can take $\nu^{\prime}$ with $r_{a}^{\nu^{\prime}}=\begin{cases}\mu_{1}-\epsilon_{\nu}&\text{if }\mu_{a}=\mu_{1}\\ \mu_{a}&\text{if }\mu_{a}<\mu_{1}\end{cases}$ for $a\in[K]$. Easy to validate $\|\{r_{a}^{\nu^{\prime}}\}_{a=1}^{K}-\{r_{a}^{\nu}\}_{a=1}^{K}\|_{\infty}\leq\epsilon_{\nu}$ and $i_{F}(\nu^{\prime})=\textsf{None}\notin i_{F}(\nu)\cup([K]\setminus i^{*}(\nu))$
Thus, we have shown $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\Delta_{0,1}^{4}}$ is a lower bound of $10K^{4}+T_{0}(\delta)+\frac{\pi^{2}}{24}$.
A.1.2 Theorem 3 in Kano et al. (2017)
Following Appendix A.2 in Li and Cheung (2025), we can also simplify the Theorem 3 in Kano et al. (2017). By applying the algorithm HDoC, for $\nu\in\mathcal{S}^{\text{pos}}$, we have
$\mathbb{E}_{\nu}\tau\leq n_{1}+\sum_{a=2}^{K}\left(\frac{\log(K\max_{j\in[K]}n_{j})}{2(\Delta_{1,a}-2\epsilon)^{2}}+\delta n_{a}\right)+\frac{K^{2-\frac{\epsilon^{2}}{(\min_{a\in[K]}\Delta_{0,a}-\epsilon)^{2}}}}{2\epsilon^{2}}+\frac{K(5+\log\frac{1}{2\epsilon^{2}})}{4\epsilon^{2}},$
where $n_{a}=\frac{1}{(\Delta_{0,a}-\epsilon)^{2}}\log\left(\frac{4\sqrt{K/\delta}}{(\Delta_{0,a}-\epsilon)^{2}}\log\frac{5\sqrt{K/\delta}}{(\Delta_{0,a}-\epsilon)^{2}}\right)$, $\epsilon=\min\Big\{\min\limits_{a\in[K]}\Delta_{0,a},\min\limits_{a\in[K-1]}\frac{\Delta_{a,a+1}}{2}\Big\}$. From direct calculation, we can observe
$n_{1}\geq\frac{\log\frac{K}{\delta}}{2\Delta_{0,1}^{2}},\quad\frac{K(5+\log\frac{1}{2\epsilon^{2}})}{4\epsilon^{2}}\geq\frac{K}{4\epsilon^{2}}$
$\log(K\max_{j\in[K]}n_{j})\geq\log(\frac{K}{2}\log\frac{1}{\delta})\geq\log\log\frac{1}{\delta}.$
From the above inequality, we can conclude
$\sum_{a=2}^{K}\frac{\log(K\max_{j\in[K]}n_{j})}{2(\Delta_{1,a}-2\epsilon)^{2}}$
$\geq$ $\sum_{a=2}^{K}\frac{\log\log\frac{1}{\delta}}{2(\Delta_{1,a}-2\epsilon)^{2}}$
$\geq$ $\sum_{a=2}^{K}\frac{\log\log\frac{1}{\delta}}{2\Delta_{1,a}^{2}}$
$\geq$ $\frac{H_{1}}{2}\log\log\frac{1}{\delta}.$
Thus, we can conclude
$n_{1}+\sum_{a=2}^{K}\left(\frac{\log(K\max_{j\in[K]}n_{j})}{2(\Delta_{1,a}-2\epsilon)^{2}}+\delta n_{a}\right)+\frac{K^{2-\frac{\epsilon^{2}}{(\min_{a\in[K]}\Delta_{0,a}-\epsilon)^{2}}}}{2\epsilon^{2}}+\frac{K(5+\log\frac{1}{2\epsilon^{2}})}{4\epsilon^{2}}$
$\geq$ $\frac{\log\frac{K}{\delta}}{2\Delta_{0,1}^{2}}+\frac{H_{1}}{2}\log\log\frac{1}{\delta}+\frac{K}{4\epsilon^{2}}.$
A.1.3 Thereom 2 in Jourdan and Réda (2023)
Jourdan and Réda (2023) do not explicitly provide upper bound for $\mathbb{E}\tau$. Therorem 2 in Jourdan and Réda (2023) suggests that for a positive instance $\nu$.
$\mathbb{E}_{\nu}\tau$ $\leq C^{\text{pos}}(\delta)+\frac{K\pi^{2}}{6}+1$
where
$C^{\text{pos}}(\delta)=$ $\sup\left\{t:t\leq\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)(\sqrt{c(t,\delta)}+\sqrt{3\log t})^{2}+D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})\right\}$
$2c(t,\delta)=$ $\bar{W}_{-1}\left(2\log\frac{K}{\delta}+4\log\log(e^{4}t)+\frac{1}{2}\right)$
$\bar{W}_{-1}(x)=$ $-W_{-1}(-e^{-x})\approx x+\log x.$
$W_{-1}$ is the negative branch of the Lambert W function, and $D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})$ is further defined in the Appendix F in Jourdan and Réda (2023), as
$D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})=$ $S_{\mu}+1-\frac{6\log S_{\mu}}{\Delta_{0,1}^{2}}$
$S_{\mu}=$ $h_{1}\left(12\sum_{a=1}^{K}\frac{1}{\Delta_{0,a}^{2}},n_{0}K+2m\right)$
$h_{1}(z,y)=$ $z\bar{W}_{-1}(\frac{y}{z}+\log z)$
$\bar{W}_{-1}(x)=$ $-W_{-1}(-e^{-x})$
From the definition, we know
$C^{\text{pos}}(\delta)\geq$ $\sup\left\{t:t\leq\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)c(t,\delta)\right\}$
$\geq$ $\sup\left\{t:t\leq\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)\left(\log\frac{K}{\delta}+2\log\log(e^{4}t)+\frac{1}{4}\right)\right\}$
$\geq$ $\sup\left\{t:t\leq\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)\log\frac{K}{\delta}+2\log 4\right\}$
$\geq$ $\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)\log\frac{K}{\delta}.$
Thus, we can conclude $\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)\log\frac{K}{\delta}$ is a lower bound of $C^{\text{pos}}(\delta)+\frac{K\pi^{2}}{6}+1$.
A.1.4 Theorem 8 in Katz-Samuels and Jamieson (2020)
As discussed in the Appendix A.1 in Li and Cheung (2025), proposed Algorithm FWER-FWPD only work on the positive instance with a different definition of $\mathbb{E}\tau$. But we can adapt their result. Apply Algorithm FWER-FWPD on a positive instance $\nu$, Theorem 8 in Katz-Samuels and Jamieson (2020) suggests that
$\mathbb{E}\tau\leq O\left(\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)\log\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)\right),$
where $\Delta=\min_{a:\mu_{a}>\mu_{0}}\Delta_{a,0}$.
Not hard to see
$\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)\log\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)$
$\geq$ $\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}\right)\log\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}\right).$
A.2 Notation Description
Table LABEL:tab:notation presents part of the notation description used in the following appendix.
Table 2: Notation Summary
Symbol Type Meaning
$K$ Integer Number of arms.
$a,i,j$ Integer in $[K]$ Arm Index
$\delta$ Real Tolerance level
$\hat{a}$ NA Output of Algorithm 1, 2, takes value in $K]\cup\{\textsf{None}\}$
$\hat{a}_{k}$ NA Rough guess of an exploration period in Algorithm 2, takes value in $K]\cup\{\textsf{None},\textsf{Not Complete}\}$
$k$ Integer Phase index in Algorithm 2
$\nu,\nu^{\prime},\nu^{\prime\prime},\tilde{\nu}$ NA An 1-identification instance
$\rho_{a}$ NA Reward distribution of arm $a$
$\sigma(\nu)$ NA Switch the arm positions in instance $\nu$ with a permutation $\sigma$
$\mu_{a}$ Real Mean value of arm $a$.
$m$ Integer Maximum arm index whose mean value is above $\mu_{0}$, we usually assume $\mu_{1}\geq\cdots\geq\mu_{m}\geq\mu_{0}>\mu_{m+1}\geq\cdots\geq\mu_{K}$
$r_{a}^{\nu}$ Real Mean value of arm $a$ instance $\nu$.
$\Delta_{i,j}$ Real Mean gap between arm index $i$ and $j$.
$A_{t}$ Integer in $[K]$ Action in round $t$
$B$ Subset of $[K]$ A bracket
$b,\tilde{b}$ Integer Bracket index in $\lceil\log_{2}K\rceil+1]$
$\tau,\tau(B)$ Random Variable $\tau$ is the termination round of Algorithm 1, $\tau(B)$ is the termination round of an Algorithm 2
$\tau_{k}^{\text{ee}}(B),\tau_{k}^{\text{et}}(B)$ Random Variable The total length of the exploration / exploitation periods at to the end of phase $k$, of Algorithm 2 with bracket $B$ as input
$\tau,\tau_{n},\tilde{\tau},\tau^{\prime},\tau_{n}^{j}(a)$ Random Variable Stopping times for analysis
$\mathcal{E}$ Random Event Random Event
$Z_{s},Z_{t}$ Random Variable Realized KL divergence, usually follows $N(\frac{\Delta^{2}}{2},\Delta^{2})$ as we only consider Gaussian distribution with variance $1$ as reward distribution in the proof lower bound
$D,D_{i}$ Random Variable $D:=\max_{1\leq t\leq n}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}}{2}\right)$. See the proof of Lemma 7
$N_{a}(t)$ Random Variable Pulling times of arm $a$ up to round $t$
$T_{j}(a)$ Real Value A pulling times threshold dependent on arm $a$ and $j$
$U(t,\delta)$ Real Value Half length of confidence interval with pulling times $t$
$\mathcal{H}^{ee},\mathcal{H}^{et},Q$ Random Set Collected history in Algorithm 2
$N_{a}^{\text{ee}}(\mathcal{H}^{ee}),N_{a}^{\text{et}}(\mathcal{H}^{et})$ Random Variable Pulling times of arm $a$ in history $\mathcal{H}^{ee},\mathcal{H}^{et}$
$N_{a}^{\text{ee}}(t)$ Random Variable Pulling times of arm $a$ at round $t$ of an exploration period
$N_{a}^{\text{et}}(t)$ Random Variable Pulling times of arm $a$ at round $t$ of an exploitation period
$\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee}),\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{et})$ Random Variable Empirical mean value of arm $a$ in history $\mathcal{H}^{ee},\mathcal{H}^{et}$
$\text{UCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta)$ Random Variable Upper Confidence Bound calculated from history $\mathcal{H}^{ee}$
$\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta)$ Random Variable Lower Confidence Bound calculated from history $\mathcal{H}^{ee}$
$\text{UCB}^{\text{et}}_{a}(\mathcal{H}^{et},\delta)$ Random Variable Upper Confidence Bound calculated from history $\mathcal{H}^{et}$
$\text{LCB}^{\text{et}}_{a}(\mathcal{H}^{et},\delta)$, Random Variable Lower Confidence Bound calculated from history $\mathcal{H}^{et}$
$C$ Real Value Fixed Constant In Algorithm 2, we take $C$ to any real number greater than 1
$t^{\text{ee}},t^{\text{et}}$ Integer Counter of the total pulling times in exploration and exploitation periods, used in Algorithm 2
$\delta_{k},\beta_{k},\alpha_{k}$ Real Value Hyperparameters of Algorithm 2 to control the pulling process
$X_{a,s}^{\text{ee}}$ Random Variable Realized reward by pulling arm $a$ at the $s$-th times in the exploration period
$X_{a,s}^{\text{et}}$ Random Variable Realized reward by pulling arm $a$ at the $s$-th times in the exploitation period
$\kappa^{\text{ee}}_{b},\kappa^{\text{ee}}_{B}$ Random Variable Minimum phase index such that the concentration event holds in the exploration period of algorithm copy $b$. If we only discuss a given bracket $B$, we do not emphasize its bracket index and may use $\kappa^{\text{ee}}_{B}$.
$\kappa^{\text{et}}_{b},\kappa^{\text{et}}_{B}$ Random Variable Minimum phase index such that the concentration event holds in the exploitation period of algorithm copy $b$. If we only discuss a given bracket $B$, we do not emphasize its bracket index and may use $\kappa^{\text{et}}_{B}$.
$\hat{\mu}^{ee}_{a,b}(t)$ Random Variable The empirical mean value of arm $a$ of the first $t$ collected reward in the exploration period of algorithm copy $b$
$\hat{\mu}^{et}_{a,b}(t)$ Random Variable The empirical mean value of arm $a$ of the first $t$ collected reward in the exploitation period of algorithm copy $b$
$\hat{\mu}_{a,t}$ Random Variable The empirical mean value of arm $a$ of the first $t$ collected reward. We use this notation if a specific bracket and a specific exploration/exploitation period is given, we may turn to use $\hat{\mu}_{a,t}$ instead of $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$.
Appendix B Proof of Lower Bound
B.1 Proof of Preparation
The first step is to prove Lemma 4, showing that we suffice to only consider symmetric algorithm
Proof of Lemma 4.
Given the alg, we define $\widetilde{\text{alg}}$ as the following. Shuffle all the arms randomly before the pulling process, and then apply the algorithm alg. We firstly show $\widetilde{\text{alg}}$ is indeed symmetric. For any pulling sequence $\{a_{t}\}_{t=1}^{T}\in[K]^{T}$ and any permutation $\sigma$, we have
$\Pr_{\widetilde{\text{alg}},\sigma(\nu^{\prime})}(\forall t\in[T],A_{t}=a,\tau>t)=\frac{\sum_{\tilde{\sigma}\in S}\Pr_{\text{alg},\tilde{\sigma}(\nu^{\prime})}(\forall t\in[T],A_{t}=a,\tau>t)}{K!},$
which means $\tilde{\text{alg}}$ fits in the definition. Then, we have
$\max_{\sigma\in S}\mathbb{E}_{\text{alg},\sigma(\nu^{\prime})}\tau$
$\geq$ $\frac{\sum_{\sigma\in S}\mathbb{E}_{\text{alg},\sigma(\nu^{\prime})}\tau}{K!}$
$\geq$ $\frac{\sum_{\sigma\in S}\sum_{t=1}^{+\infty}\Pr_{\text{alg},\sigma(\nu^{\prime})}(\tau\geq t)}{K!}$
$=$ $\mathbb{E}_{\widetilde{\text{alg}},\nu^{\prime}}\tau$
The last thing is to show $\widetilde{\text{alg}}$ is also $\delta$-PAC. Since alg is $\delta$-PAC, we know for any permutation $\sigma$, we have
$\Pr_{\text{alg},\sigma(\nu)}(\tau<+\infty,\hat{a}\in i^{*}(\sigma(\nu)))>1-\delta.$
That means
$\Pr_{\widetilde{\text{alg}},\sigma(\nu)}(\tau<+\infty,\hat{a}\in i^{*}(\sigma(\nu)))$
$=$ $\frac{\sum_{\sigma\in S}\Pr_{\text{alg},\sigma(\nu)}(\tau<+\infty,\hat{a}\in i^{*}(\sigma(\nu)))}{K!}$
$>$ $\frac{\sum_{\sigma\in S}1-\delta}{K!}$
$=$ $1-\delta.$
We can conclude $\widetilde{\text{alg}}$ is also $\delta$-PAC. ∎
The next step, which is also the main trick, is to the prove Lemma 7.
Lemma 7.
Given a pair $i,j\in[K],i\neq j$ and an arbitrary instance $\nu^{\prime}$, denote $\Delta=|r^{\nu^{\prime}}_{i}-r^{\nu^{\prime}}_{j}|$. Define
$\tau_{n}=\min\{t:N_{i}(t)+N_{j}(t)=n,A_{t+1}\in\{i,j\}\}.$
Apply a symmetric algorithm, we have
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}+$
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\frac{1}{2}n\Delta^{2}\Pr_{\nu^{\prime}}(\tau_{n}<\tau).$
An observation is given the realized reward and a symmetric algorithm, the set $\{t:N_{i}(t)+N_{j}(t)=n,A_{t+1}\in\{i,j\}\}=:S_{ij}(n)$ contains at most one element. If $|S_{ij}(n)|=1$, $\tau_{n}$ will be the unique element, and we have $\tau_{n}<\tau$ holds in this case. If $|S_{ij}(n)|=0$, we have $\tau_{n}=\infty$, meaning that $\tau_{n}\geq\tau$. Another observation is
$\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}+\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}$
is the kl-divergence between two Bernoulli distribution with probability $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)$, $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)$, and not hard to see $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)+\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)=1$.
Proof of Lemma 7.
Denote $\nu^{\prime\prime}$ as the instance which switch the position of arm $i$ and arm $j$ in $\nu^{\prime}$, i.e. $r_{i}^{\nu^{\prime}}=r_{j}^{\nu^{\prime\prime}},r_{j}^{\nu^{\prime}}=r_{i}^{\nu^{\prime\prime}}$ and $r_{a}^{\nu^{\prime}}=r_{a}^{\nu^{\prime\prime}},a\neq i,j$.
We first derive
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$ (15)
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$ (16)
where $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta^{2}}{2},\Delta^{2})$. Since these two inequalities are symmetric, we suffice to prove the first one. Take $\mathcal{E}=\{\tau_{n}<\tau,A_{\tau_{n}+1}=i\}$, from the Lemma 18 in Kaufmann et al. (2016), we have
$\Pr_{\nu^{\prime\prime}}(\mathcal{E})=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})+$
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\neg\mathcal{E}\right]\Pr_{\nu^{\prime}}(\neg\mathcal{E})$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})$
$\geq$ $\exp\left(\mathbb{E}_{\nu^{\prime}}\left[-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}|\mathcal{E}\right]\right)\Pr_{\nu^{\prime}}(\mathcal{E}),$
where $\hat{KL}_{i,s}:=\log\frac{\exp(-\frac{(X_{i,s}-r_{i}^{\nu^{\prime}})^{2}}{2})}{\exp(-\frac{(X_{i,s}-r_{i}^{\nu^{\prime\prime}})^{2}}{2})}=\frac{(2X_{i,s}-r_{i}^{\nu^{\prime}}-r_{j}^{\nu^{\prime}})(r_{i}^{\nu^{\prime}}-r_{j}^{\nu^{\prime}})}{2}$, $X_{i,s}\sim N(r_{i}^{\nu^{\prime}},1)$, $\hat{KL}_{j,s}:=\log\frac{\exp(-\frac{(X_{j,s}-r_{j}^{\nu^{\prime}})^{2}}{2})}{\exp(-\frac{(X_{j,s}-r_{j}^{\nu^{\prime\prime}})^{2}}{2})}=\frac{(2X_{j,s}-r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})}{2}$, $X_{j,s}\sim N(r_{j}^{\nu^{\prime}},1)$. The last line is by the Jensen’s Inequality. Following the last line, we have
$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}|\mathcal{E}\right]$
$=$ $\frac{\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)\right]}{\Pr_{\nu^{\prime}}(\mathcal{E})},$
and we can conclude $\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\geq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)\right]$. From the above calculation, we know $\hat{KL}_{i,s},\hat{KL}_{j,s}$ both follow $N(\frac{(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})^{2}}{2},(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}}))$, and we can rewrite the conclusion as
$\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right]$
where $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta^{2}}{2},\Delta^{2})$. The above inequality suggests that we have proved (15) and (16).
By the fact that the algorithm is symmetric, we can conclude $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)$ and $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)$. Sum up inequality (15) and (16), we get
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}+$
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}$
$=$ $\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}+$
$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right).$
The remaining work is to prove
$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\frac{1}{2}n\Delta^{2}\Pr_{\nu^{\prime}}(\tau_{n}<\tau).$
Denote $D:=\max_{1\leq t\leq n}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}}{2}\right)$, we have
$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\max_{1\leq t\leq n}\sum_{s=1}^{n}Z_{s}\right)$
$=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\max_{1\leq t\leq n}\sum_{s=1}^{n}Z_{s}-\frac{\Delta^{2}}{2}+\frac{\Delta^{2}}{2}\right)$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\max_{1\leq t\leq n}\sum_{s=1}^{n}Z_{s}-\frac{\Delta^{2}}{2}\right)+\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\frac{n\Delta^{2}}{2}$
$=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)D+\frac{n\Delta^{2}}{2}\Pr_{\nu^{\prime}}(\tau_{n}<\tau).$
The remaining work is to prove $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)D\leq\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)$. To simplify the notation, we take event $\mathcal{E}=\{\tau_{n}<\tau\}$ here and we can conduct the following calculation.
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D\right]$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D\right]$
$=$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathcal{E},D>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\Pr_{\nu^{\prime}}\Big(D>s\Big)\right\}ds$
From Lemma 32, we can conclude $\Pr_{\nu^{\prime}}\big(D\geq s\big)\leq\exp(-\frac{s^{2}}{2n\Delta^{2}})$. Plug in the inequality, we get
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D\right]\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\exp(-\frac{s^{2}}{2n\Delta^{2}})\right\}ds$
$=$ $\int_{0}^{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}\Pr_{\nu^{\prime}}(\mathcal{E})ds+\int_{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2n\Delta^{2}})ds$
$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\frac{\sqrt{2\pi}\sqrt{n\Delta^{2}}}{\sqrt{2\pi}}\int_{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2})ds$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n\Delta^{2}}\exp(-\frac{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}{2})$
$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\mathcal{E}).$
The second last step is by the equality $\int_{t}^{+\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{s^{2}}{2}}ds\leq e^{-\frac{t^{2}}{2}}$ for $t>0$. The reason is
$\int_{t}^{+\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{s^{2}}{2}}ds=\int_{0}^{+\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{(s+t)^{2}}{2}}ds\leq e^{-\frac{t^{2}}{2}}\int_{0}^{+\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{s^{2}}{2}}ds\leq e^{-\frac{t^{2}}{2}}.$
We complete the proof. ∎
Besides Lemma 7, we further introduce a Lemma which is similar, but focuses on the single side.
Lemma 8.
Consider two instances $\nu^{\prime}$, $\nu^{\prime\prime}$, which only differ in arm $a_{1},a_{2},\cdots,a_{M}$, i.e. $r^{\nu^{\prime}}_{a}\neq r^{\nu^{\prime\prime}}_{a}$ holds for all $a\in\{a_{1},a_{2},\cdots,a_{M}\}$ and $r^{\nu^{\prime}}_{a}=r^{\nu^{\prime\prime}}_{a}$ holds for all $a\notin\{a_{1},a_{2},\cdots,a_{M}\}$. Denote $\Delta_{i}=|r_{a_{i}}^{\nu^{\prime}}-r_{a_{i}}^{\nu^{\prime\prime}}|$, and $\tilde{\tau}$ is an arbitrary stopping time. If event $\mathcal{E}\in\mathcal{F}_{\tilde{\tau}}$ fulfills $\mathcal{E}\subset\{N_{a_{i}}(\tilde{\tau})\leq n_{i},\forall i\in[M]\}$ with some $\{n_{i}\}_{i=1}^{M}\in\mathbb{N}^{M}$, we can conclude
$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\sum_{i=1}^{M}\sqrt{2n_{i}\Delta_{i}^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}+\sqrt{2\pi n_{i}\Delta_{i}^{2}}+\frac{1}{2}n_{i}\Delta_{i}^{2}$
Proof.
From the Lemma 18 in Kaufmann et al. (2016), we have
$\Pr_{\nu^{\prime\prime}}(\mathcal{E})=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\mathcal{E})\exp\left(-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right)$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})+$
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right)|\neg\mathcal{E}\right]\Pr_{\nu^{\prime}}(\neg\mathcal{E})$
$=$ $\mathbb{E}_{\nu^{\prime}}\left[\exp\left(-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})$
$\geq$ $\exp\left(\mathbb{E}_{\nu^{\prime}}\left[-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}|\mathcal{E}\right]\right)\Pr_{\nu^{\prime}}(\mathcal{E}),$
where $\hat{KL}_{i,s}:=\log\frac{\exp(-\frac{(X_{i,s}-r_{a_{i}}^{\nu^{\prime}})^{2}}{2})}{\exp(-\frac{(X_{i,s}-r_{a_{i}}^{\nu^{\prime\prime}})^{2}}{2})}=\frac{(2X_{i,s}-r_{a_{i}}^{\nu^{\prime}}-r_{a_{i}}^{\nu^{\prime\prime}})(r_{a_{i}}^{\nu^{\prime}}-r_{a_{i}}^{\nu^{\prime\prime}})}{2}$, $X_{i,s}\sim N(r_{a_{i}}^{\nu^{\prime}},1)$. Following the last line, we have
$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}|\mathcal{E}\right]$
$=$ $\frac{\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right]}{\Pr_{\nu^{\prime}}(\mathcal{E})}.$
and we can conclude $\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\sum_{i=1}^{M}\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right]$. We turn to find an upper bound for each $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right]$. Since $\mathcal{E}\Rightarrow N_{a_{i}}(\tilde{\tau})<n_{i}$, we have
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right]$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\max_{1\leq t\leq n_{i}}\sum_{s=1}^{t}\hat{KL}_{i,s}-\frac{\Delta_{i}^{2}}{2}\right)\right]+\frac{n_{i}\Delta_{i}^{2}}{2}\Pr_{\nu^{\prime}}(\mathcal{E})$
Denote $D_{i}:=\max_{1\leq t\leq n_{i}}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}_{i}}{2}\right)$, $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta_{i}^{2}}{2},\Delta_{i}^{2})$, we have
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D_{i}\right]$
$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D_{i}\right]$
$=$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathds{1}(\mathcal{E})\mathds{1}(D_{i}\geq 0)D_{i}>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathcal{E},D_{i}>s\Big)ds$
$\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\Pr_{\nu^{\prime}}\Big(D_{i}>s\Big)\right\}ds$
From Lemma 32, we can conclude $\Pr_{\nu^{\prime}}\big(D_{i}\geq s\big)\leq\exp(-\frac{s^{2}}{2n_{i}\Delta^{2}_{i}})$. Plug in the inequality, we get
$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D_{i}\right]\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\exp(-\frac{s^{2}}{2n_{i}\Delta^{2}_{i}})\right\}ds$
$=$ $\int_{0}^{\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}\Pr_{\nu^{\prime}}(\mathcal{E})ds+\int_{\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2n_{i}\Delta^{2}})ds$
$=$ $\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\frac{\sqrt{2\pi}\sqrt{n_{i}\Delta^{2}_{i}}}{\sqrt{2\pi}}\int_{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2})ds$
$\leq$ $\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n_{i}\Delta^{2}_{i}}\exp(-\frac{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}{2})$
$=$ $\sqrt{2n_{i}\Delta^{2}_{i}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n_{i}\Delta^{2}_{i}}\Pr_{\nu^{\prime}}(\mathcal{E}).$
In summary, we have proved
$\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}$
$\leq$ $\sum_{i=1}^{M}\sqrt{2n_{i}\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n_{i}\Delta^{2}}\Pr_{\nu^{\prime}}(\mathcal{E})+\frac{n_{i}\Delta_{i}^{2}}{2}\Pr_{\nu^{\prime}}(\mathcal{E}),$
which is equivalent to the conclusion we need to prove. ∎
B.2 Proof of Theorem 5
Following the sketch proof in section 2, we prove the lemmas one by one. The first step is to build a “constant probability lower bound”.
Lemma 9.
Apply a $\delta$-PAC and symmetric algorithm to a positive instance $\nu$, If $\max_{i,j\in[K]}\Delta_{i,j}<1$ and $\delta<10^{-8}$, $\frac{1}{8m}\geq\sqrt{\delta}$, all holds, we have
$\sum_{j=1}^{m}\Pr_{\nu}\Big(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Big)\geq\frac{1}{2}$
Proof of Lemma 9.
Recall the notation $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, and the $\delta$-PAC assumption requires
$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m]\cup\{a:\mu_{a}=\mu_{0}\})>1-\delta$
[…]
We prove the claim by contradiction, if $Pr_{\nu}(\tau<+\infty,\hat{a}\in\{a:\mu_{a}=\mu_{0}\})\geq 4\delta$ holds, we can find a $T_{0}$, such that
$\Pr_{\nu}(\tau<T_{0},\hat{a}\in\{a:\mu_{a}=\mu_{0}\})\geq 2\delta.$ (17)
We define $\nu^{\prime}(\epsilon)$ with $r_{a}^{\nu^{\prime}(\epsilon)}=\begin{cases}\mu_{0}-\epsilon&\mu_{a}=\mu_{0}\\ r_{a}^{\nu}&\text{else}\end{cases}$, where $\epsilon$ is an arbitrary positive number. We know $i^{*}(\nu^{\prime}(\epsilon))=[m]$ and we have
$\Pr_{\nu^{\prime}(\epsilon)}(\tau<+\infty,\hat{a}\in[m])>1-\delta$
$\Rightarrow$ $\Pr_{\nu^{\prime}(\epsilon)}(\tau=+\infty\text{ or }\hat{a}\notin[m])<\delta$
[…]
$\log\frac{\Pr_{\nu}(\mathcal{E})}{\Pr_{\nu^{\prime}(\epsilon)}(\mathcal{E})}\leq$ $\sum_{a:\mu_{a}=\mu_{0}}\sqrt{2T_{0}\epsilon^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}+\sqrt{2\pi T_{0}\epsilon^{2}}+\frac{1}{2}T_{0}\epsilon^{2}$
$\log\frac{\Pr_{\nu}(\mathcal{E})}{\Pr_{\nu^{\prime}(\epsilon)}(\mathcal{E})}\geq$ $\log\frac{2\delta}{\delta}=\log 2$
[…]
$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m]\cup\{a:\mu_{a}=\mu_{0}\})>1-\delta$
[…]
$\Rightarrow$ $\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])>1-\delta-4\delta$
$\Rightarrow$ $\Pr_{\nu}(\tau=+\infty\text{ or }\hat{a}\notin[m])<5\delta$
$\Rightarrow$ $\Pr_{\nu}(\hat{a}\notin[m])<5\delta$
Now, we are ready to prove the Lemma. We prove the conclusion by contradiction. If the Lemma doesn’t hold, we have
$\Pr_{\nu}\Big(\exists j\in[m],\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Big)<\frac{1}{2}$
$\Leftrightarrow$ $\Pr_{\nu}\left(\cap_{j=1}^{m}\Big(\{\hat{a}\neq j\}\cup\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\right)\geq\frac{1}{2}$
$\Rightarrow$ $\Pr_{\nu}\left(\cap_{j=1}^{m}\{\hat{a}\neq j\}\right)+\Pr_{\nu}\left(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\right)\geq\frac{1}{2}$
$\Rightarrow$ $5\delta+\Pr_{\nu}\left(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\right)\geq\frac{1}{2}$
$\Rightarrow$ $\Pr_{\nu}\left(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\right)\geq\frac{1}{4}$
$\Rightarrow$ $\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}\in[m]\}\right)+\Pr_{\nu}\left(\{\hat{a}\notin[m]\}\right)\geq\frac{1}{4}$
$\Rightarrow$ $\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}\in[m]\}\right)+5\delta\geq\frac{1}{4}$
$\Rightarrow$ $\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}\in[m]\}\right)\geq\frac{1}{8}$
$\Rightarrow$ $\sum_{j^{\prime}=1}^{m}\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}=j^{\prime}\}\right)\geq\frac{1}{8}$
From the last step, we know there exists $j_{0}$, such that
$\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}=j_{0}\}\right)\geq\frac{1}{8m}\geq\sqrt{\delta}.$
We can consider instance $\nu_{j_{0}}^{(0)}$ whose reward of arm $j_{0}$ is $\mu_{0}-\epsilon$ for some $\epsilon>0$, others remain the same as $\nu$. By the Lemma 8, we have
$\log\frac{\sqrt{\delta}}{\delta}$
$\leq$ $\log\frac{\Pr_{\nu}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}=j_{0}\}\right)}{\Pr_{\nu_{j_{0}}^{(0)}}\left(\Big(\cap_{j=1}^{m}\{N_{j}(\tau)\leq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Big)\cap\{\hat{a}=j_{0}\}\right)}$
$\leq$ $\sqrt{2\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}+\sqrt{2\pi\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}}+\frac{1}{2}\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}$
$\leq$ $\sqrt{2\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}\log\frac{1}{\sqrt{\delta}}}+\sqrt{2\pi\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}}+\frac{1}{2}\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}}(\Delta_{j_{0},0}+\epsilon)^{2}.$
Taking $\epsilon\rightarrow 0$, the last line suggests that
$\frac{1}{2}\log\frac{1}{\delta}\leq\sqrt{\frac{\log\frac{1}{\delta}}{100}\log\frac{1}{\delta}}+\sqrt{2\pi\frac{\log\frac{1}{\delta}}{100}}+\frac{1}{2}\frac{\log\frac{1}{\delta}}{100},$
which leads to a contradiction, as $\delta<10^{-8}$. ∎
The next step is to prove a lemma analyzing the pulling behavior.
Lemma 10.
Assume $\delta<\frac{1}{e}$. For any $j\in[m]$, $a\in[K],a\neq j$, define
$\tau_{n}^{j}(a)=$ $\min\{t:N_{j}(t)+N_{a}(t)=n,A_{t+1}\in\{j,a\}\}$
$T_{j}(a)=$ $\lceil\frac{1}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}\right)}\rceil.$
Then, for $n=0,1,2,\cdots,T_{j}(a)-1$, we have
$\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=j)\log\frac{\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=j)}{\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=a)}+$
$\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=a)\log\frac{\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=a)}{\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=j)}$
$<$ $0.4\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$
Proof of Lemma 10.
The main idea is to utilize Lemma 7. From the definition, we know $T_{j}(a)\geq 1$. For $n=1,2,\cdots,T_{j}(a)-1$, we have
$\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}$
$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu}(N_{a}(\tau)+N_{j}(\tau)>n)}}$
$\leq$ $\sqrt{2\Big(\frac{2}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\log\frac{e}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}}\Big)\Delta_{a,j}^{2}\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}}$
$\leq$ $\sqrt{\frac{1}{50}}.$
The third step is by the fact that $\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\}\subset\{N_{a}(\tau)+N_{j}(\tau)>n\}$ holds for all $n=0,1,\cdots,T_{j}(a)-1$. The last step is by the fact that $\frac{\Delta_{a,j}^{2}}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq 1$. For the same reason, we have
$\sqrt{2\pi n\Delta^{2}}\leq$ $\sqrt{\frac{\pi}{50}}$
$\frac{n\Delta^{2}}{2}\leq$ $\frac{1}{200}.$
Not hard to validate that $\sqrt{\frac{1}{50}}+\sqrt{\frac{\pi}{50}}+\frac{1}{200}<0.4$. Applying Lemma 7, we prove the result. ∎
Based on the Lemma 10, we can prove a lower bound connecting $\mathbb{E}N_{A}(\tau)$ and $\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)$, which is the following Lemma.
Lemma 11.
Following the same notation as Lemma 10 and still assume $\delta<\frac{1}{e}$, we have
$\mathbb{E}_{\nu}N_{a}(\tau)\geq\frac{T_{j}(a)}{4}\Pr_{\nu}\left(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right)$
holds for all $j\in[m],a\in[K],a\neq j$.
Proof of Lemma 11.
We firstly prove a claim: Assume $\tau_{n+1}^{j}(a)<\tau$, for all $t$ such that $\tau_{n}+1^{j}(a)<t<\tau_{n+1}^{j}(a)+1$, we have $A_{t}\neq a$. We prove this claim by contradiction. Assume there exists $t$ such that $\tau_{n}^{j}(a)+1<t<\tau_{n+1}^{j}(a)+1$, $A_{t}=a$. Then, from the definition of $\tau_{n}^{j}(a)$, we know $N_{j}(\tau_{n}^{j}(a))+N_{a}(\tau_{n}^{j}(a))=n,N_{j}(\tau_{n}^{j}(a)+1)+N_{a}(\tau_{n}^{j}(a)+1)=n+1$. Since $t>\tau_{n}^{j}(a)+1$, which means $t-1\geq\tau_{n}^{j}(a)+1$, we can conclude
$N_{j}(t)+N_{a}(t)$
$=$ $\mathds{1}(A_{t}=j)+\mathds{1}(A_{t}=a)+N_{j}(t-1)+N_{a}(t-1)$
$=$ $1+N_{j}(t-1)+N_{a}(t-1)$
$\geq$ $1+N_{j}(\tau_{n}^{j}(a)+1)+N_{a}(\tau_{n}^{j}(a)+1)$
$=$ $n+2.$
However, $t\leq\tau_{n+1}^{j}(a)$ suggests that $N_{j}(t)+N_{a}(t)\leq N_{1}(\tau_{n+1}^{j}(a))+N_{a}(\tau_{n+1}^{j}(a))=n+1$, which leads to a contradiction. Then we can rewrite $\mathbb{E}_{\nu}N_{a}(\tau)$ as the following.
$\mathbb{E}_{\nu}N_{a}(\tau)$
$=$ $\mathbb{E}\sum_{t=1}^{+\infty}\mathds{1}(A_{t}=a)\mathds{1}(\tau\geq t)$
$=$ $\mathbb{E}\sum_{n=0}^{+\infty}\mathds{1}(A_{\tau_{n}^{j}(a)+1}=a)\mathds{1}(\tau>\tau_{n}^{j}(a))$
$\geq$ $\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a,\tau>\tau_{n}^{j}(a)\Big).$
From Lemma 10, we know
$kl\Bigg(\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big),1-\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)\Bigg)<0.4$
where $kl(x,y)=x\log\frac{x}{y}+(1-y)\log\frac{1-y}{1-x}$. From Pinsker’s Inequality, we can further conclude
$2\Bigg(1-2\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)\Bigg)^{2}<0.4$
$\Rightarrow$ $\frac{1-\sqrt{0.2}}{2}<\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)$
$\Rightarrow$ $\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a,\tau>\tau_{n}^{j}(a)\Big)>\frac{1}{4}\Pr_{\nu}\Big(\tau>\tau_{n}^{j}(a)\Big).$
The last line suggests that
$\mathbb{E}_{\nu}N_{a}(\tau)$
$\geq$ $\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a,\tau>\tau_{n}^{j}(a)\Big)$
$\geq$ $\frac{1}{4}\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(\tau>\tau_{n}^{j}(a)\Big)$
$\geq$ $\frac{1}{4}\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)$
$=$ $\frac{T_{j}(a)}{4}\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big),$
which is the conclusion we wish to prove. ∎
Given Lemma 11, we are able to take the optimal value of an optimization problem as the lower bound of $\mathbb{E}_{\nu}\tau$, which is Lemma 6.
Proof of Lemma 6.
For any $j\in[m]$, by Lemma 11, we have
$\mathbb{E}\tau$
$=$ $\mathbb{E}N_{j}(\tau)+\sum_{a\neq j}\mathbb{E}\sum_{n=0}^{+\infty}\mathds{1}(A_{\tau_{n}^{j}(a)+1}=a)\mathds{1}(\tau>\tau_{n}^{j}(a))$
$\geq$ $\mathbb{E}N_{j}(\tau)\mathds{1}(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}})+\sum_{a\neq j}\mathbb{E}\sum_{n=0}^{T_{j}(a)-1}\mathds{1}(A_{\tau_{n}^{j}(a)+1}=a)\mathds{1}(\tau>\tau_{n}^{j}(a))$
$\geq$ $\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}})+\sum_{a\neq j}\frac{1}{4}T_{j}(a)\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)$
$\geq$ $\frac{1}{100\Delta_{j,0}^{2}}\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)+\sum_{a\neq j}\frac{1}{4}T_{j}(a)\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)$
$\geq$ $\sum_{a=1}^{K}\frac{1}{4}T_{j}(a)\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big).$
Thus, we can conclude
$\mathbb{E}\tau$
$\geq$ $\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{1}{4}T_{j}(a)\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)$
$\geq$ $\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big)}{800\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}\right)}$ (18)
Meanwhile, we can also derive
$\mathbb{E}\tau\geq$ $\sum_{j=1}^{m}\mathbb{E}\tau\mathds{1}(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}})$
$\geq$ $\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr_{\nu}(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}})$ (19)
Merge (18) and (19), notice that Lemma 9 asserts
$\sum_{j=1}^{m}\Pr_{\nu}\Big(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Big)\geq\frac{1}{2},$
we know $\mathbb{E}\tau$ must be lower bounded by the optimal value of the following optimization problem.
$\begin{array}[]{cl}\min&\frac{1}{800}\max\left\{\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j},\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}\right\}\\ s.t.&\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}\\ &0\leq p_{j}\leq 1\end{array}$
Then, we turn to show $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$ holds for $j\in[m],a\in[K]$. If $\mu_{a}\geq\mu_{j}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. If $\mu_{j}>\mu_{a}>\mu_{0}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}=\frac{1}{\Delta_{j,0}^{2}}>\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. If $\mu_{0}\geq\mu_{a}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}=\frac{1}{\Delta_{j,a}^{2}}>\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$. Thus, we know $\mathbb{E}\tau$ must be lower bounded by the optimal value of the following optimization problem.
$\begin{array}[]{cl}\min&\frac{1}{3200}\max\left\{\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j},\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}\right\}\\ s.t.&\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}\\ &0\leq p_{j}\leq 1\end{array}$
Reformulate the above optimization problem, we can turn to solve
$\begin{array}[]{cl}\min&v\\ s.t.&v\geq\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j}\\ &v\geq\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)},\forall j\in[m]\\ &\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}\\ &0\leq p_{j}\leq 1\\ &v\geq 0.\end{array}$
Given Lemma 6, the last step is to prove a lower bound of the optimal value of (2).
Lemma 12.
Denote opt as the optimal value of problem (2). We can prove
$\text{opt}\geq\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right)$
Proof.
Proof of Lemma 12
We first show function $f(p)=\frac{p}{1+\log\frac{1}{p}}$ is convex for $p\in(0,1]$. Easy to calculate
$\frac{\partial\frac{p}{1+\log\frac{2}{p}}}{\partial p}=$ $\frac{1+\log\frac{2}{p}-p(-\frac{1}{p})}{(1+\log\frac{2}{p})^{2}}$
$=$ $\frac{1}{1+\log\frac{2}{p}}+\frac{1}{(1+\log\frac{2}{p})^{2}}$
$\frac{\partial^{2}\frac{p}{1+\log\frac{2}{p}}}{\partial p^{2}}=$ $\left(-\frac{1}{(1+\log\frac{2}{p})^{2}}-\frac{2}{(1+\log\frac{2}{p})^{3}}\right)\left(-\frac{1}{p}\right)$
$=$ $\left(\frac{1}{(1+\log\frac{2}{p})^{2}}+\frac{2}{(1+\log\frac{2}{p})^{3}}\right)\frac{1}{p}.$
It is straight forward to see $\frac{\partial^{2}\frac{p}{1+\log\frac{2}{p}}}{\partial p^{2}}>0$ holds for all $p\in(0,1]$, suggesting that function $\frac{p}{1+\log\frac{1}{p}}$ is convex, further indicating (2) is a convex optimization problem.
Introduce dual variables $\alpha\geq 0,\gamma\geq 0,\{\beta_{j}\}_{j=1}^{m}\in\mathbb{R}_{+}^{m}$ and Lagrangian function
$L(\alpha,\gamma,\{\beta_{j}\}_{j=1}^{m};\{p_{j}\}_{j=1}^{m},v)$
$=$ $v+\left(\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j}-v\right)\alpha+$
$\sum_{j=1}^{m}\left(\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}-v\right)\beta_{j}+(\frac{1}{2}-\sum_{j=1}^{m}p_{j})\gamma$
$=$ $v(1-\alpha-\sum_{j=1}^{m}\beta_{j})+\sum_{j=1}^{m}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j}+\beta_{j}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}-\gamma p_{j}\right)+\frac{\gamma}{2}.$
Take $\alpha^{0}=\frac{1}{2}$, $\beta_{j}^{0}=\frac{\frac{1}{2j}}{2\sum_{s=1}^{m}\frac{1}{2s}}$, $\gamma^{0}=\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\frac{1}{2}\cdot+\frac{\frac{1}{2j}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}}{(1+\log(m+1))(1+\log 200+\log m)}$. Since (2) is a convex optimization problem, by the strong duality, we can assert
$\text{opt}=$ $\max_{\alpha\geq 0,\gamma\geq 0,\{\beta_{j}\}_{j=1}^{m}\in\mathbb{R}_{+}^{m}}\min_{\{p_{j}\}_{j=1}^{m}\in[0,1]^{m},v\geq 0}L(\alpha,\gamma,\{\beta_{j}\};\{p_{j}\},v)$
$\geq$ $\min_{\{p_{j}\}_{j=1}^{m}\in[0,1]^{m},v\geq 0}L(\alpha^{0},\gamma^{0},\{\beta_{j}^{0}\};\{p_{j}\},v).$
In the following, we suffice to prove
$\min_{\{p_{j}\}\in[0,1]^{m},v\geq 0}L(\alpha^{0},\gamma^{0},\{\beta_{j}^{0}\}_{j=1}^{m};\{p_{j}\}_{j=1}^{m},v)$
$\geq$ $\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$
Before that, we prove a claim
$\min_{1\geq p_{j}\geq 0}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p_{j}+\frac{p_{j}}{1+\log\frac{2}{p_{j}}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p_{j}\right)\geq-\frac{\gamma^{0}}{100m}$ (20)
Define $\bar{L}_{j}(p)=\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p+\frac{p}{1+\log\frac{2}{p}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p$ and $p^{*}_{j}=\arg\min\limits_{1\geq p\geq 0}\bar{L}_{j}(p)$. Not hard to validate $\bar{L}_{j}(p)$ is a convex function. We prove the claim by discussing two cases.
If $\frac{\partial\bar{L}_{j}(p)}{\partial p}\geq 0$ holds for all $0\leq p\leq 1$, we have
$\min_{1\geq p\geq 0}\bar{L}_{j}(p)=\lim_{p\rightarrow 0}\bar{L}_{j}(p)=0>-\frac{\gamma^{0}}{100m},$
suggesting that the claim (20) is true.
If $\frac{\partial\bar{L}_{j}(p)}{\partial p}<0$ for some $p$, by the fact that $\bar{L}_{j}(p)$ is convex, we can assert there exists $p_{0}\in(0,1)$ such that $\bar{L}_{j}(p)$ is decreasing in the interval $(0,p_{0})$, and $\bar{L}_{j}(p^{*}_{j})<0$. Then, we can prove $p^{*}_{j}<\frac{1}{100m}$ by contradiction. If $p^{*}_{j}\geq\frac{1}{100m}$, we have
$\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p^{*}_{j}+\frac{p^{*}_{j}}{1+\log\frac{2}{p^{*}_{j}}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p^{*}_{j}$
$\geq$ $\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p^{*}_{j}+\frac{p^{*}_{j}}{1+\log 200+\log m}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p^{*}_{j}$
$=$ $p^{*}_{j}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}+\frac{\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}}{1+\log 200+\log m}-\gamma^{0}\right)$
$\geq$ $p^{*}_{j}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\frac{1}{2}+\frac{\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\frac{\frac{1}{2j}}{2\sum_{s=1}^{m}\frac{1}{2s}}}{1+\log 200+\log m}-\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\frac{1}{2}-\frac{\sum_{a=1}^{K}\frac{1}{2j\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}}{(1+\log 200+\log m)(1+\log(m+1))}\right)$
$>$ $0.$
The second last line is by the definition of $\gamma^{0}$. The last line is by the fact that $2\sum_{s=1}^{m}\frac{1}{2s}<1+\log(m+1)$ for $m\geq 1$. This inequality contradicts with the fact that $\bar{L}_{j}(p^{*}_{j})<0$. Thus, we can conclude
$\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p^{*}_{j}+\frac{p^{*}_{j}}{1+\log\frac{2}{p^{*}_{j}}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p^{*}_{j}$
$\geq$ $-\gamma^{0}p^{*}_{j}$
$\geq$ $-\frac{\gamma^{0}}{100m}.$
That means in the case $\frac{\partial\bar{L}_{j}(p)}{\partial p}<0$, we can still conclude the claim (20) is true.
Based on the (20), we have
$\min_{\{p_{j}\}_{j=1}^{m}\in[0,1]^{m},v\geq 0}L(\alpha^{0},\gamma^{0},\{\beta_{j}^{0}\};\{p_{j}\},v)$
$=$ $\min_{v\geq 0}v(1-\alpha^{0}-\sum_{j=1}^{m}\beta_{j}^{0})+$
$\sum_{j=1}^{m}\min_{1\geq p_{j}\geq 0}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j}+\beta_{j}^{0}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}-\gamma^{0}p_{j}\right)+\frac{\gamma^{0}}{2}$
$\geq$ $\sum_{j=1}^{m}-\frac{\gamma^{0}}{100m}+\frac{\gamma^{0}}{2}$
$\geq$ $\frac{\gamma^{0}}{4}.$
The second last line is by the fact that $1-\alpha^{0}-\sum_{j=1}^{m}\beta_{j}^{0}=1-\frac{1}{2}-\sum_{j=1}^{m}\frac{\frac{1}{2j}}{2\sum_{s=1}^{m}\frac{1}{2s}}=0$ and the claim (20). From the definition of $\gamma^{0}$, we know
$\min_{\{p_{j}\}\in[0,1]^{m},v\geq 0}L(\alpha^{0},\gamma^{0},\{\beta_{j}^{0}\};\{p_{j}\},v)$
$\geq$ $\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right),$
which also completes the proof. ∎
Appendix C Proof of Upper Bound
C.1 Rigorous Presentation of Algorithm
$U(t,\delta)=\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}.$ (21)
where $\lceil x\rceil:=\max\{x,1\}$
Before presenting the full statement of the algorithm, we first introduce some notations
$N_{a}^{\text{ee}}(\mathcal{H}^{ee})=\sum_{s=1}^{|\mathcal{H}^{ee}|}\mathds{1}(A_{s}(\mathcal{H}^{\text{ee}})=a),N_{a}^{\text{et}}(\mathcal{H}^{et})=\sum_{s=1}^{|\mathcal{H}^{et}|}\mathds{1}(A_{s}(\mathcal{H}^{\text{et}})=a)$ (22)
$\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee})=\frac{\sum_{s=1}^{|\mathcal{H}^{ee}|}\mathds{1}(A_{s}(\mathcal{H}^{\text{ee}})=a)X_{s}(\mathcal{H}^{\text{ee}})}{N_{a}^{\text{ee}}(\mathcal{H}^{ee})},$
$\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{et})=\frac{\sum_{s=1}^{|\mathcal{H}^{et}|}\mathds{1}(A_{s}(\mathcal{H}^{\text{et}})=a)X_{s}(\mathcal{H}^{\text{et}})}{N_{a}^{\text{et}}(\mathcal{H}^{et})}$
$\text{UCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta)=\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee})+U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta}{K})$
$\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta)=\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee})-C\cdot U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta}{K})$
$\text{UCB}^{\text{et}}_{a}(\mathcal{H}^{et},\delta)=\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{et})+U(N_{a}^{\text{et}}(\mathcal{H}^{ee}),\frac{\delta}{K})$
$\text{LCB}^{\text{et}}_{a}(\mathcal{H}^{et},\delta)=\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{et})-U(N_{a}^{\text{et}}(\mathcal{H}^{ee}),\frac{\delta}{K})$
where we consider $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$ as stacks that can add index to the stored item, and denote $A_{s}(\mathcal{H}^{\text{ee}}),X_{s}(\mathcal{H}^{\text{ee}})$ as the action and reward stored with index $s$ in $\mathcal{H}^{\text{ee}}$. Same idea applies to $A_{s}(\mathcal{H}^{\text{et}}),X_{s}(\mathcal{H}^{\text{et}})$. If $N_{a}^{\text{ee}}(\mathcal{H}^{ee})=0$ holds for some $a\in[K]$, we take $\text{UCB}^{\text{ee}}_{a}=+\infty,\text{LCB}^{\text{ee}}_{a}=$-$\infty$. Same idea applies to $N_{a}^{\text{et}}(\mathcal{H}^{et}),\text{UCB}^{\text{et}}_{a},\text{LCB}^{\text{et}}_{a}$. Given the above notation, we introduce algorithm 2, which is adapted from Li and Cheung (2025). The following analysis is very similar to the Section 4 and Appendix B in Li and Cheung (2025), but we still remain the full description and proof for completeness. In the following subsection, we discuss each part of the algorithm line by line and highlight our adaptation of history literature in Appendix C.1.5.
1Input: Bracket $B$, Threshold $\mu_{0}$, Tolerance level $\delta$, Tunable parameter $C>1$, Arm number $K$;
2
3Initialize: $\delta_{k}=\frac{1}{3^{k}},\beta_{k}=2^{k},\alpha_{k}=5^{k},\forall k\in\mathbb{N}$, History $\mathcal{H}^{\text{ee}}=\mathcal{H}^{\text{et}}=\emptyset$, Temporary Container $Q=\emptyset$. Introduce notations in (22), $t^{\text{ee}}=t^{\text{et}}=0$;
4 for phase index $k=1,2,\cdots$ do
 /* Phase $k$ starts */
5 while True do
6    if $t^{\text{ee}}\geq 1$ and $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})\geq\mu_{0}$ then
7         Take $\hat{a}_{k}\leftarrow A_{t}^{\text{ee}}$;
8       $\mathcal{H}^{\text{ee}}\leftarrow\mathcal{H}^{\text{ee}}\setminus\{(A^{\text{ee}}_{t},X)\}$, $N_{A^{\text{ee}}_{t}}^{\text{ee}}\leftarrow N_{A^{\text{ee}}_{t}}^{\text{ee}}-1$, $Q\leftarrow Q\cup\{(A^{\text{ee}}_{t},X)\}$, $\hat{a}\leftarrow A_{t}^{\text{ee}}$;
9       while $N_{\hat{a}}^{\text{et}}\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}-1$ do
10            Sample $X\sim\rho_{\hat{a}}$, $\mathcal{H}^{\text{et}}\leftarrow\mathcal{H}^{\text{et}}\cup\{(\hat{a},X)\}$, $N_{\hat{a}}^{\text{et}}\leftarrow N_{\hat{a}}^{\text{et}}+1$, $t^{\text{et}}\leftarrow t^{\text{et}}+1$;
11          if $\text{LCB}^{\text{et}}_{\hat{a}}(\mathcal{H}^{\text{et}},\frac{\delta}{\alpha_{k}})>\mu_{0}$ then
12               Return $\hat{a}$ as a qualified arm;
13
14            end if
15
16         end while
        Break;
       // Enter phase $k+1$
17
18      end if
19    if $\forall a\in B$, $\text{UCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k})\leq\mu_{0}$ then
20       if $\delta_{k}\leq\frac{\delta}{3}$ and $|B|=K$ then
21            Take $\hat{a}_{k}=\textsf{None}$ and return None;
          /* Only the bracket with size $K$ is allowed to output None */
22
23       else
            Take $\hat{a}_{k}=\textsf{Not Completed}$ and break;
          // Enter phase $k+1$
24
25         end if
26
27      end if
28    if $N_{a}^{\text{ee}}>(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1,\forall a\in B$ then
         Take $\hat{a}_{k}=\textsf{Not Completed}$ and break;
       // Enter phase $k+1$
29
30      end if
31      Determine $A^{\text{ee}}_{t}=\arg\max\limits_{a\in B,N_{a}^{\text{ee}}\leq(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1}\text{UCB}_{a}(\mathcal{H}^{\text{ee}},\delta_{k})$;
32    if $\exists(a,X)\in Q$ such that $a=A^{\text{ee}}_{t}$ then
33       $\mathcal{H}^{\text{ee}}\leftarrow\mathcal{H}^{\text{ee}}\cup\{(a,X)\}$, $N_{A^{\text{ee}}_{t}}^{\text{ee}}\leftarrow N_{A^{\text{ee}}_{t}}^{\text{ee}}+1$, $Q\leftarrow Q\setminus\{(a,X)\}$;
34
35    else
36         Sample $X\sim\rho_{A_{t}^{\text{ee}}}$, $\mathcal{H}^{\text{ee}}\leftarrow\mathcal{H}^{\text{ee}}\cup\{(A^{\text{ee}}_{t},X)\}$, $N_{A^{\text{ee}}_{t}}^{\text{ee}}\leftarrow N_{A^{\text{ee}}_{t}}^{\text{ee}}+1$, $t^{\text{ee}}\leftarrow t^{\text{ee}}+1$;
37
38      end if
39
40   end while
  Take $\tau^{\text{ee}}_{k}(B)\leftarrow t^{\text{ee}},\tau^{\text{et}}_{k}(B)\leftarrow t^{\text{et}}$;
 // Independent of algorithm design, only for analysis
 /* Phase $k$ ends */
41
42 end for
Algorithm 2 Sequential Exploration Exploitaion (SEE)
C.1.1 From Line 2 to 2: Input and Initialized Parameter
Algorithm 2 takes bracket $B$, threshold $\mu_{0}$, tolerance level $\delta$, arm number $K$, tunable parameter $C$ as input, and only pull arms in $B$. $\mu_{0},K$ are the same as what Algorithm 1 receives, while Algorithm 1 sets $\delta\leftarrow\frac{\delta}{\lceil\log_{2}K\rceil+1}$ when calling Algorithm 2. We take $C=1.01$ by default, but we can use any value greater than $1$.
Besides input parameter specified by the problem nature and Algorithm 1, 2 also initialize some parameters to manage the pulling process. Parameter $\{\delta_{k}=\frac{1}{3^{k}},\beta_{k}=2^{k},\alpha_{k}=5^{k}\}_{k=1}^{+\infty}$ are used to set up the tolerance level and budget during the pulling process. It is possible to specify other values for $\delta_{k},\beta_{k},\alpha_{k}$, see Appendix B.1 in Li and Cheung (2025). $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}},Q$ are used to store reward collected from pulling arms. As we have discussed above, $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$ are two stacks that can maintain the arrival order of reward of each arm $a\in B$. This requirement is to guarantee we are able to transfer the latest collected reward out of $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$. $Q$ is used to guarantee $\text{LCB}_{a}^{\text{ee}}$ cannot be greater than $\mu_{0}$ at the start of each phase, which will be discussed in section C.1.4.
C.1.2 Line 2 to Line 2: Overall Idea of Algorithm 2
At the start of Algorithm 2, it splits the arm distributions into two parallel and independent copies, one for exploration period, and another for exploitation period. Correspondingly, reward collected in the exploration periods are stored in $\mathcal{H}^{\text{ee}}$. Reward collected in the exploration periods, which is from Line 2 to Line 2, are stored in $\mathcal{H}^{\text{et}}$.
After specifying the parameters through Line 2 and 2, Algorithm 2 splits the pulling process into multiple phases, noted by the For loop in Line 2. In each phase $k$, Algorithm 2 consists of two parts. The first part is the conditions from Line 2 to Line 2. The second part, which is called Exploration Period, ranges from Line 2 to 2. Exploration Period adopts UCB rule in Jamieson and Nowak (2014) to pull arms. These two parts are sequentially executed in the While Loop at Line 2. Though the condition of the While Loop is always True, but the conditions in the first part is able to guarantee that the Algorithm 2 never gets stuck in a infinity loop. See the discussion in section C.1.3.
At a phase $k$, the exploration period utilizes $\delta_{k}$ as the tolerance level to apply UCB rule of pulling arms, while the exploitation period utilizes $\frac{\delta}{\alpha_{k}}$ as the tolerance level to calculate Upper and Lower Confidence Bounds and only pull a single arm. Once the tolerance level of both periods are below $\frac{\delta}{3}$ and $\hat{a}$ takes value in $K]\cup\{\textsf{None}\}$, Algorithm 2 will terminate. What’s more, Algorithm 2 also sets up a maximum pulling times for both periods. For the exploration period at phase $k$, the maximum pulling times for each arm $a\in B$ is $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$. For the exploitation period at phase $k$, the maximum pulling times of the single arm is $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$. Section C.1.3 and C.1.4 provide more details.
C.1.3 Line 2 to 2: Conditions on Phase Transition
There are three possible conditions to end the exploration period in a phase $k$. The first one is Line 2, judge whether the algorithm should enter the exploitation period. Intuitively speaking, the inequality $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})\geq\mu_{0}$ provides a recommended qualified arm with confidence $1-\Theta(\delta_{k})$. Since $A_{t}^{\text{ee}}$ is the action we take in the last round, this condition also guarantees $\text{LCB}^{\text{ee}}_{A_{t}^{\text{ee}}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ holds before the collecting the latest sample of $A_{t}^{\text{ee}}$. To further confirm whether arm $A_{t}^{\text{ee}}$ is qualified, we need to keep pulling arm $A_{t}^{\text{ee}}$ by collecting samples independent of the exploration periods. Line 2 suggests that the algorithm will terminate and output $A_{t}^{\text{ee}}$, if the LCB calculated by the tolerance level $\delta/\alpha_{k}$ is above $\mu_{0}$ during the exploitation process. Before entering the while loop (Line 2) in the exploitation period, Line 2 transfer the latest collected sample of $A_{t}^{\text{ee}}$ in the exploration period into $Q$. Usage of $Q$ is to guarantee Lemma 22, requiring $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. Similar to the reason described in the section 4.1 of Li and Cheung (2025), Lemma 22 is required to prove Lemma 23, further Theorem 3. We put more discussion of $Q$ to section C.1.4.
The second condition is Line 2, aiming to output None for negative case. To be specific, if the Upper Confidence Bound(UCB) in the exploration periods are all below $\mu_{0}$, Algorithm 2 will leave the exploration periods in phase $k$. Then, the condition in Line 2 determines the next period. If the bracket $B$ is indeed $[K]$ and the tolerance level $\delta_{k}<\frac{\delta}{3}$, Algorithm 2 will terminate the overall pulling process and output None. In other cases, Algorithm 2 will enter the exploration period in the next phase $k+1$. In the case $\max_{a\in B}\mu_{a}<\mu_{0}$ and $|B|<K$, it is possible that Algorithm 2 never break the for loop in Line 2.
The third condition is Line 2, which is correlated to the budget of the exploration period. For a phase $k$, Algorithm 2 sets up a maximum pulling times $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ for each arm $a\in B$. Recall we only consider picking $A^{\text{ee}}_{t}$ from those arms whose total pulling times from phase 1 to phase $k$ is below $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ in Line 2. The third condition is Line 2 is to enter the exploration period in phase $k+1$, if pulling times are no less than $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1$ holds for all $a\in B$. Line 2 and 2 also guarantee Algorithm 2 will not get stuck into a infinite loop of “While True” in Line 2.
C.1.4 Line 2 to 2: Exploration
In phase $k$, Algorithm 2 use tolerance level $\delta_{k}$ to calculate Upper Confidence Bound(UCB) according to (22) and determine the arm $A_{t}^{\text{ee}}$ by Line 2. Line 2 to 2 determine where to collect sample of arm $A_{t}^{\text{ee}}$. If $Q$ contains one sample of arm $A_{t}^{\text{ee}}$, Algorithm 2 will transfer this sample to $\mathcal{H}^{\text{ee}}$ without collecting a new sample from distribution $\nu_{A_{t}^{\text{ee}}}$. If not, Algorithm 2 samples a new random variable from $\nu_{A_{t}^{\text{ee}}}$.
As we mentioned in section C.1.3, usage of $Q$ is to guarantee Lemma 22, requiring $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. Someone may question whether we can directly abandon the latest collected sample of $A_{t}^{\text{ee}}$ in $\mathcal{H}$, which seems also able to guarantee $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$. We cannot to do so. By transferring samples into or out $Q$, we are able to derive
Lemma 13.
Given an algorithm copy alg taking bracket $B$ as input. Condition on reward sequence in the exploration period $\{X_{a,s}^{\text{ee}}\}_{s=1}^{+\infty}$, for each $a\in B$, we can assert $\hat{\mu}_{a}(\mathcal{H}^{\text{ee}})\in\{\frac{\sum_{s=1}^{t}X_{a,s}^{\text{ee}}}{t}\}_{t=1}^{+\infty}$ throughout the pulling process.
The reason is we do not change the arrival order of $\{X_{a,s}^{\text{ee}}\}_{s=1}^{+\infty}$ for each $a\in B$. To analyze Algorithm 2, we need to introduce
$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\right\}$
$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\right\}.$
where $\lceil x\rceil^{+}=\max\{x,1\}$. $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately. More details are in section C.2. Lemma 13 suggests that for phase $k\geq\kappa^{\text{ee}}_{b}$, we can guarantee $\text{LCB}_{a}^{\text{ee}}<\mu_{a}<\text{UCB}_{a}^{\text{ee}}$ holds for all $a\in B_{b}$, which further helps to prove upper bounds of the pulling times for all $a\in B_{b}$. Simply dropping the latest collected samples are unable to guarantee 13, making it much harder to conduct the analysis.
Besides, we also care about the size of $Q$ during the pulling process, as we can observe the following Lemmas.
Lemma 14.
Given an algorithm copy alg, throughout the pulling process, we have
$t^{\text{ee}}=|\mathcal{H}^{\text{ee}}|+|Q|$
always hold.
The reason is all the collected samples in the exploration periods are either in $\mathcal{H}^{\text{ee}}$ or $Q$. Notice that we only transfer sample from $\mathcal{H}^{\text{ee}}$ to $Q$ in Line 2. In this case, $A_{t}^{\text{ee}}$ just get pulled in the last round of exploration. Meanwhile, once $Q$ contains a sample for an arm $a\in B$, condition in 2 guarantees Algorithm 2 will never collect new samples of arm $a$ in the exploration period. This property implies following.
Lemma 15.
Consider an algorithm copy alg taking bracket $B$ as input, we have $|Q|\leq|B|$ throughout the pulling process
At the end of this subsection, we present a Lemma similar to 13, but for the exploitation period.
Lemma 16.
Given an algorithm copy alg taking bracket $B$ as input. Condition on $\{X_{a,s}^{\text{et}}\}_{s=1}^{+\infty}$, for each $a\in B$, we can assert $\hat{\mu}_{a}(\mathcal{H}^{\text{et}})\in\{\frac{\sum_{s=1}^{t}X_{a,s}^{\text{et}}}{t}\}_{t=1}^{+\infty}$ throughout the pulling process.
This is correct, as we collect the samples in the exploitation period sequentially in $\mathcal{H}^{\text{et}}$.
C.1.5 Adaptation of History Literature
Compared to the Algorithm SEE in Li and Cheung (2025), our main adaptation consists of two points. The first adaptation is to put the stopping conditions (Line 2 to 2) before the exploration periods (Line 2 to 2). This design is to guarantee Lemma 29, guaranteeing Algorithm 2 will not pull arm $a$ whose $\text{UCB}_{a}^{\text{ee}}$ is below $\mu_{0}$.
The second point is for each phase $k$, we set up a maximum pulling times of each arm in the exploration period, instead of maximum pulling times of the total pulling times (reference line index is to be added). The reason is the arm in a bracket $B$ is randomized, it is hard to analyze the required pulling times that a bracket need to output a $\hat{a}_{k}$.
C.2 Properties of Algorithm 2
This subsections follows the Appendix B in Li and Cheung (2025), which are nearly the same. Here we remain the full proof for completeness, as we have modified some implementation details and proof.
In Algorithm 1, we create $\lceil\log_{2}K\rceil+1$ independent copies of Algorithm 2. For each copy index $b\in[\lceil\log_{2}K\rceil+1]$, we further create two independent sampling copies for exploration and exploitation periods separately. We define
$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\right\}$
$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\right\}.$
where $\lceil x\rceil^{+}=\max\{x,1\}$. $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately. The above definition of $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ are conditioned on the realized reward of $\{X_{a,s}^{\text{ee}}\}_{a=1,s=1}^{K,+\infty}$, $\{X_{a,s}^{\text{et}}\}_{a=1,s=1}^{K,+\infty}$ collected by the algorithm copy $b$. $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ are the minimum phase index such that concentration event holds. Since $B_{b}$ is determined by an exogenous random permutation, we know $B_{b}$ is independent with $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$.
In the following, we will analyze the properties of Algorithm 2, a given bracket $B$. For convenience, given a bracket $B$ with index $b$, we denote $\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B}$ as corresponding $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$, without emphasizing the index $b$. Further, we denote $\tau_{k}^{\text{ee}}(B),\tau_{k}^{\text{et}}(B)$ as the value of $t^{\text{ee}},t^{\text{et}}$ at the end of phase $k$, i.e. each time we enter Line 2 in Algorithm 2.
Lemma 17.
For any $B$, we have $\Pr(\kappa^{\text{ee}}_{B}\geq k)\leq\frac{\pi^{2}}{6}\delta_{k-1}$, $\Pr(\kappa^{\text{ee}}_{B}\geq k)\leq\frac{\pi^{2}}{6}\frac{\delta}{\alpha_{k-1}}$
Proof of Lemma 17.
By the Lemma 35, we have
$\Pr(\kappa^{\text{ee}}_{B}\geq k)\leq$ $\Pr\left(\exists t,\exists a\in[K],|\hat{\mu}^{ee}_{a}(t)-r^{\nu}_{a}|\geq\sqrt{\frac{4\log\frac{2K}{\delta_{k-1}}+8\log\log(2t)}{t}}\right)$
$\leq$ $\sum_{a=1}^{K}\Pr\left(\exists t,|\hat{\mu}^{ee}_{a}(t)-r^{\nu}_{a}|\geq\sqrt{\frac{4\log\frac{2K}{\delta_{k-1}}+8\log\log(2t)}{t}}\right)$
$\leq$ $\sum_{a=1}^{K}\frac{\pi^{2}\delta_{k-1}}{6K}$
$=$ $\frac{\pi^{2}\delta_{k-1}}{6}$
and
$\Pr(\kappa^{\text{et}}_{B}\geq k)\leq$ $\Pr\left(\exists t,\exists a\in[K],|\hat{\mu}^{et}_{a}(t)-r^{\nu}_{a}|\geq\sqrt{\frac{4\log\frac{2K\alpha_{k-1}}{\delta}+8\log\log(2t)}{t}}\right)$
$\leq$ $\sum_{a=1}^{K}\Pr\left(\exists t,|\hat{\mu}^{ee}_{a}(t)-r^{\nu}_{a}|\geq\sqrt{\frac{4\log\frac{2K\alpha_{k-1}}{\delta}+8\log\log(2t)}{t}}\right)$
$\leq$ $\sum_{a=1}^{K}\frac{\pi^{2}\delta}{6K}$
$=$ $\frac{\pi^{2}\delta}{6\alpha_{k-1}}.$
Lemma 18.
Consider an Algorithm copy $b$ with bracket $B$, we have $\Pr(\text{alg copy }b\text{ outputs }\hat{a}\text{ such that }\mu_{\hat{a}}\leq\mu_{0})\leq\frac{\pi^{2}}{6}\frac{\delta}{5}$.
Proof of Lemma 18.
By the Line 2 to 2, we know
$\Pr(\text{alg copy }b\text{ outputs }\hat{a}\text{ such that }\mu_{\hat{a}}<\mu_{0})$
$\leq$ $\Pr\left(\exists a,s.t.\mu_{a}<\mu_{0},t\in\mathbb{N},\hat{\mu}_{a}^{\text{et}}(t)-\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{1}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\geq\mu_{0}\right)$
$\leq$ $\sum_{a:\mu_{a}<\mu_{0}}\Pr\left(t\in\mathbb{N},\hat{\mu}_{a}^{\text{et}}(t)-\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{1}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\geq\mu_{a}\right)$
$\leq$ $\sum_{a:\mu_{a}\leq\mu_{0}}\frac{\pi^{2}}{6}\frac{\delta}{\alpha_{1}K}$
$\leq$ $\frac{\pi^{2}}{6}\frac{\delta}{5}.$
The second last line is by the Lemma 35. ∎
Lemma 19.
Assume $\mu_{1}>\mu_{0}$. Consider an Algorithm copy $\lceil\log_{2}K\rceil+1$, whose allocated bracket is $[K]$. We have $\Pr(\text{alg copy }\lceil\log_{2}K\rceil+1\text{ outputs }\textsf{None})\leq\frac{\pi^{2}}{6}\frac{\delta}{3}$.
Proof of Lemma 19.
By the condition in Line 2, we know
$\Pr(\text{alg copy }\lceil\log_{2}K\rceil+1\text{ outputs }\textsf{None})$
$\leq$ $\Pr(\exists t\in\mathbb{N},\delta_{k}\leq\frac{\delta}{3},\hat{\mu}_{1}^{\text{ee}}(t)+\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\leq\mu_{0})$
$\leq$ $\Pr(\exists t\in\mathbb{N},\hat{\mu}_{1}^{\text{ee}}(t)+\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta/3}}}{t}\leq\mu_{0})$
$\leq$ $\frac{\pi^{2}}{6}\frac{\delta}{3}.$
The last line is by the Lemma 35. ∎
Lemma 20.
For any $B$, and any $a\in B$, we have $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\leq(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$.
Proof of Lemma 20.
We use induction on $k\in\mathbb{N}$. For $k=1$, by the Line 2 in Algorithm 2, we konw
$N_{a}^{\text{ee},B}\Big(\tau_{1}^{\text{ee}}(B)\Big)\leq(C+1)^{2}\beta_{1}\log\frac{4K}{\delta_{1}}.$
If the conclusion holds for $k$, then for $k+1$, we can still conclude
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}\right\}$
$\leq$ $\max\left\{(C+1)^{2}\beta_{k-}\log\frac{4K}{\delta_{k-1}},(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}\right\}$
$=$ $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}.$
The first step is still caused by the Line 2 in Algorithm 2. The second step is from the foundation. The last line shows that the induction holds. ∎
Lemma 21.
For any $B$, $\tau_{k}^{\text{ee}}(B)\leq|B|(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$, $\tau_{k}^{\text{et}}(B)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$
Proof of Lemma 21.
Regarding $\tau_{k}^{\text{ee}}(B)\leq|B|(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$, we can directly conclude this from Lemma 20.
Regarding $\tau_{k}^{\text{et}}(B)\leq\sum_{k^{\prime}=1}^{k}\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k^{\prime}}\log\frac{4K\alpha_{k^{\prime}}}{\delta}$, we suffice to show $\tau_{k}^{\text{et}}(B)-\tau_{k-1}^{\text{et}}(B)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$ holds for any phase index $k$. By direct calculation, we have
$\tau_{k}^{\text{et}}(B)-\tau_{k-1}^{\text{et}}(B)$
$=$ $\sum_{a=1}^{K}N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k}^{\text{et}}(B)\Big)-N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k-1}^{\text{et}}(B)\Big)\mathds{1}(\hat{a}_{k}=a)$
$\leq$ $\sum_{a=1}^{K}N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k}^{\text{et}}(B)\Big)\mathds{1}(\hat{a}_{k}=a)$
$\leq$ $\sum_{a=1}^{K}\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}\mathds{1}(\hat{a}_{k}=a)$
$\leq$ $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}.$
The second last line is by the loop condition in Line 2 of Algorithm 2. ∎
Lemma 22.
For any $B$, at the start of each phase $k$, i.e. each time we enter Line 2 in Algorithm 2, we have $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ holds for all $a\in B$.
Proof of Lemma 22.
We apply induction on $k$. For $k=1$, this is obviously true, as we denote the $\text{LCB}_{a}^{\text{ee}}$ as -$\infty$ if the pulling times of arm $a$ is 0.
If the conclusion holds for the phase k1k-1, we prove the case of $k$ by discussing all the possible values of $\hat{a}_{k-1}$. Before the start of phase $k$, there are three possible scenarios for the value of $\hat{a}_{k-1}$. Phase k1k-1 may end with taking $\hat{a}_{k-1}$ takes value at the Line 2, 2 or 2.
If $\hat{a}_{k-1}$ takes Not Completed at the Line 2, we have we $\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k})<\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$, forall $a\in[K]$.
If $\hat{a}_{k-1}$ takes Not Completed at the Line 2, We can also assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\in[K]$. Prove by contradiction. Assume there exists $a\in[K]$ such that $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})\geq\mu_{0}$ at the end of phase k1k-1. From the induction, since $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$ at the start of phase k1k-1, the algorithm would enter Line 2 and $\hat{a}_{k-1}$ cannot take value Not Completed. Thus, we can assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\in[K]$ holds at the end of phase k1k-1. Further, at the start of phase $k$, we have $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$.
The remaining work is to prove the conclusion in the case that $\hat{a}_{k-1}\in[K]$ at the Line 2. We can first assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\neq\hat{a}_{k-1}$. The reason is from the induction, $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$ holds for all $a$ at the start of phase k1k-1. Once there exists an arm $a$ whose LCB is above $\mu_{0}$ after an arm pull, the algorithm will enter Line 2, suggesting that there is at most one arm whose LCB is above $\mu_{0}$.
Then, we turn to analyze $\text{LCB}_{\hat{a}_{k-1}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})$ at the start of phase $k$. Before the execution of Line 2, the following inequalities must hold
$\frac{X+\sum_{s=1}^{N_{\hat{a}_{k-1}}^{\text{ee}}-1}X_{\hat{a}_{k-1},s}}{N_{\hat{a}_{k-1}}^{\text{ee}}}-C\cdot U\left(N_{\hat{a}_{k-1}}^{\text{ee}},\frac{\delta_{k-1}}{K}\right)>\mu_{0}$
$\frac{\sum_{s=1}^{N_{\hat{a}_{k-1}}^{\text{ee}}-1}X_{\hat{a}_{k-1},s}}{N_{\hat{a}_{k-1}}^{\text{ee}}-1}-C\cdot U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k-1}}{K}\right)\leq\mu_{0}$
Since we put the last collected sample $X$ into $Q$ in Line 2, we can assert $\hat{\mu}_{\hat{a}_{k-1}}$ before the start of phase $k$ must be $\frac{\sum_{s=1}^{N_{\hat{a}_{k-1}}^{\text{ee}}-1}X_{\hat{a}_{k-1},s}}{N_{\hat{a}_{k-1}}^{\text{ee}}-1}$. Since $U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k}}{K}\right)>U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k-1}}{K}\right)$, we know before the start of phase $k$, $\text{LCB}_{\hat{a}_{k-1}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ must hold.
Following lemma suggests that the exploration period is able to output correct answer, given some concentration event.
Lemma 23.
For $B$ such that $\max_{a\in B}\mu_{a}\geq\mu_{j}$ for some $\mu_{j}>\mu_{0}$, and phase index $k\geq\kappa^{\text{ee}}_{B}$, Algorithm 2 either takes $\hat{a}_{k}\in[K]$ such that $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$ or takes $\hat{a}_{k}=\textsf{Not Completed}$
Proof of Lemma 23.
From the assumption $k\geq\kappa^{\text{ee}}_{B}$, by the Lemma 13, we know
$\hat{\mu}_{a}(\mathcal{H}^{\text{ee}})-U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})\leq\mu_{a}\leq\hat{\mu}_{a}(\mathcal{H}^{\text{ee}})+U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$ (23)
holds for $a\in B$ through out the phase $k$.
We first show that Algorithm 2 never enters Line 2. Denote $a^{*}(B)=\arg\max_{a\in B}\mu_{a}$. From the assumption, we know $\mu_{a^{*}(B)}>\mu_{0}$. Together with the definition $\text{UCB}^{\text{ee}}_{a}(\mathcal{H}^{ee},\delta_{k})=\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{ee})+U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$, we know $\text{UCB}^{\text{ee}}_{a^{*}(B)}(\mathcal{H}^{ee},\delta_{k})>\mu_{0}$ holds through out the phase $k$, suggesting that the condition in Line 2 never holds. Meanwhile, to quit the While loop in the Line 2, Algorithm 2 must enter one of the Lines in 2, 2, 2. Then, we can conclude $\hat{a}_{k}$ only takes value in $K]\cup\{\textsf{Not Completed}\}$.
The remaining work is to prove if $\hat{a}_{k}\in[K]$, we have $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$. For simplicity, we denote $\tau_{k}^{\text{ee}}$ as $\tau_{k}^{\text{ee}}(B)$, and prove by contradiction. If $\hat{a}\in[K]$ and $\mu_{\hat{a}}<\omega\mu_{j}+(1-\omega)\mu_{0}<\omega\mu_{a^{*}(B)}+(1-\omega)\mu_{0}$, we have
$A_{\tau^{\text{ee}}_{k}}^{\text{ee}}=\hat{a}_{k},\hat{a}_{k}=\arg\max_{i\in B}\hat{\mu}_{i,N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)}+U(N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})$
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}$
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1,\frac{\delta_{k}}{K})<\mu_{0}.$
where $\hat{\mu}_{i,N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)}$ denotes the empirical mean value of arm i of the first $N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)$ samples. The third line is by the Lemma 22, suggesting $\text{LCB}_{\hat{a}_{k}}^{\text{ee}}<\mu_{0}$ must hold before the latest collection of the sample.
As we take $U(t,\frac{\delta}{K})=\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}t\rceil,1\}}\log\frac{2K(\lceil\log_{2}t\rceil)^{2}}{\delta}}}{t}$, we get
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}$
$\Leftrightarrow$ $\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}>\mu_{0}$
[…]
$\Rightarrow$ $\omega\mu_{a^{*}(B)}+(1-\omega)\mu_{0}-(C-1)\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}>\mu_{0}$
$\Leftrightarrow$ $\omega(\mu_{a^{*}(B)}-\mu_{0})>(C-1)\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}$
$\Leftrightarrow$ $\frac{2\omega(\mu_{a^{*}(B)}-\mu_{0})}{C-1}>\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}.$
By $\omega=\frac{C-1}{C+3}$, we have $\frac{(1-\omega)}{2}=\frac{1-\frac{C-1}{C+3}}{2}=\frac{4}{2(C+3)}=\frac{2\omega}{C-1}$, which means
$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}>\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}.$
Notice that
$\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}$
$\leq$ $\frac{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}$
$\leq$ $\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})},$
we can conclude
$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}\geq\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}.$
On the other hands,
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}+U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
[…]
$\Leftarrow$ $\omega\mu_{1}+(1-\omega)\mu_{0}+2U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
$\Leftrightarrow$ $\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\leq(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})$
$\Leftrightarrow$ $\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\leq\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}.$
In summary, we can conclude
$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau)}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}$
$\Rightarrow$ $\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}(\tau^{\text{ee}}-1)}+U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
[…]
$\Rightarrow$ $A_{\tau^{\text{ee}}_{k}}^{\text{ee}}\neq\hat{a}_{k}.$
The last line suggests a contradiction. We can conclude $\mu_{\hat{a}}\geq\frac{C-1}{C+3}\mu_{j}+\frac{4}{C+3}\mu_{0}$. ∎
Lemma 24.
For $B$ such that $\max_{a^{\prime}\in B}\mu_{a^{\prime}}\geq\mu_{j}$ for some $\mu_{j}>\mu_{0}$, phase index $k\geq\kappa^{\text{ee}}_{B}$, for each arm $a\in B$, we have
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right\}$
Proof of Lemma 24.
We first prove an intermediate conclusion, which is for each arm $a\in B$, we have
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$ (24)
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right\}$
hols for for any phase index $k\geq\kappa^{\text{ee}}_{B}$. If the above conclusion holds, we can use induction to conclude
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\Bigg\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k-1}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$
$\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\Bigg\}$
$=$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right\},$
which proves the Lemma. Thus, we suffice to prove the intermediate conclusion (24).
From the assumption $k\geq\kappa^{\text{ee}}_{B}$, we know (23) still holds. Fix the arm $a\in B$, we prove the claim by discussing the value of $\mu_{a}$. We consider two cases, $\mu_{a}\geq\frac{\mu_{j}+\mu_{0}}{2}$ or $\mu_{a}<\frac{\mu_{j}+\mu_{0}}{2}$.
For arm $a$ such that $\mu_{a}\geq\frac{\mu_{j}+\mu_{0}}{2}$ holds, we suffice to prove
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{1}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\right\}.$
We can conduct direct calculation, as
$t\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
[…]
$\Leftrightarrow$ $\mu_{a}-(C+1)\sqrt{\frac{4\log\frac{2K(\log_{2}(2t))^{2}}{\delta_{k}}}{t}}>\mu_{0}$
$\Rightarrow$ $\mu_{a}-(C+1)\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}>\mu_{0}$
[…]
By the Lemma 22, we know at the start of phase $k$, $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$, suggesting that $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)\leq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$. Then, we can assert
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)<\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1.$
If not, the algorithm will output $\hat{a}_{k}=a$ before $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$ holds.
For arm $a$ such that $\mu_{a}<\frac{\mu_{j}+\mu_{0}}{2}$ holds, we suffice to prove
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}\right\}.$
We can also conduct direct calculation, with
$t\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}$
[…]
$\stackrel{{\scriptstyle C>1}}{{\Rightarrow}}$ $\mu_{a}+2\sqrt{\frac{4\log\frac{2K(\log_{2}(2t))^{2}}{\delta_{k}}}{t}}\leq\mu_{j}$
$\Rightarrow$ $\mu_{a}+2\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\leq\mu_{j}$
[…]
If $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)\geq\frac{28(C+1)^{2}\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}+1$, we know $\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})<\mu_{j}\leq\mu_{a^{*}(B)}\leq\text{UCB}_{a^{*}(B)}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})$ holds through out the phase $k$, and arm $a$ will never get pulled. We can assert $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)=N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)$.
If $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)<\frac{28(C+1)^{2}\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}+1$, we can assert $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\leq\frac{29(C+1)^{2}\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a,j}^{2}}\right)}{\Delta_{a,j}^{2}}$ as the arm $a$ never gets pulled once $\text{UCB}_{a}^{\text{ee}}\leq\mu_{a^{*}(B)}\leq\text{UCB}_{a^{*}(B)}^{\text{ee}}$ holds. We complete the proof. ∎
Lemma 25.
For $B$ such that $\max_{a^{\prime}\in B}\mu_{a^{\prime}}\geq\mu_{j}$, and phase index $k\geq\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{112(C+1)^{2}}{\Delta_{j,0}^{2}}\rceil\}$, Algorithm 2 will take $\hat{a}_{k}\in[K]$ such that $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$.
Proof of Lemma 25.
From the proof of Lemma 23, we know Algorithm 2 will never meet the condition in Line 2.
Meanwhile, from the calculation in the proof of Lemma 24, we know
$t\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a^{*}(B),0}^{2}}\right)}{\Delta_{a^{*}(B),0}^{2}}$
$\Rightarrow$ $\hat{\mu}_{a^{*}(B),t}-C\cdot\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}>\mu_{0}$
and
$k\geq\lceil\log_{2}\frac{112(C+1)^{2}}{\Delta_{j,0}^{2}}\rceil$
$\Rightarrow$ $\beta_{k}\geq\frac{112(C+1)^{2}}{\Delta_{j,0}^{2}},\log\frac{1}{\delta_{k}}\geq\log 3\log_{2}\frac{112(C+1)^{2}}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $\frac{1}{2}\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{56\log\frac{2K}{\delta_{k}}}{\Delta_{j,0}^{2}},\frac{1}{2}\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{56\log\log\frac{24(C+1)^{2}}{\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{56\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{j,0}^{2}}\right)}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{j,0}^{2}}\right)}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1\geq\frac{28(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\Delta_{a^{*}(B),0}^{2}}\right)}{\Delta_{a^{*}(B),0}^{2}},$
which means the Algorithm will first fulfill the condition in Line 2 before fulfilling the condition in Line 2. This observation concludes $\hat{a}_{k}\in[K]$. By the Lemma 23, we can further conclude $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$. ∎
Lemma 26.
For $B$ such that $\max_{a^{\prime}\in B}\mu_{a^{\prime}}\geq\mu_{j}$, and phase index $k\geq\max\{\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$. If Algorithm 2 enters Line 2 with $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$, Algorithm 2 will terminate and output $\hat{a}_{k}$.
Proof of Lemma 26.
From the condition $k\geq\kappa^{\text{et}}_{B}$, by the Lemma 16, we know
$\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{\text{et}})-U(N_{a}^{\text{et}}(\mathcal{H}^{et}),\frac{\delta}{K\alpha_{k}})\leq\mu_{a}\leq\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{\text{et}})+U(N_{a}^{\text{et}}(\mathcal{H}^{et}),\frac{\delta}{K\alpha_{k}})$
holds for $a\in B$ through out the phase $k$.
By similar calculation in the proof of Lemma 24, we know
$t\geq\frac{112\left(\log\frac{2K\alpha_{k}}{\delta}+\log\log\frac{96}{\omega^{2}\Delta_{j,0}^{2}}\right)}{\omega^{2}\Delta_{j,0}^{2}}$
$\Rightarrow$ $t\geq\frac{112\left(\log\frac{2K\alpha_{k}}{\delta}+\log\log\frac{96}{\Delta_{\hat{a}_{k},0}^{2}}\right)}{\Delta_{\hat{a}_{k},0}^{2}}$
$\Rightarrow$ $\hat{\mu}_{\hat{a}_{k},t}^{\text{et}}-\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}>\mu_{0}.$
Here $\hat{\mu}_{\hat{a}_{k},t}^{\text{et}}$ denotes the empirical mean of first $t$ samples collected in the exploitation period. Notice that
$k\geq\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil$
$\Rightarrow$ $\beta_{k}\geq\frac{448}{\Delta_{j,0}^{2}},\log\alpha_{k}\geq\log\frac{96(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\log 5$
$\Rightarrow$ $\beta_{k}\log\frac{4K\alpha_{k}}{\delta}\geq\frac{448\log\frac{2K\alpha_{k}}{\delta}}{\Delta_{j,0}^{2}},\beta_{k}\log\frac{4K\alpha_{k}}{\delta}\geq\frac{448\log\log\frac{96}{\omega^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $\beta_{k}\log\frac{4K\alpha_{k}}{\delta}\geq\frac{224\left(\log\frac{2K\alpha_{k}}{\delta}+\log\log\frac{96}{\omega^{2}\Delta_{j,0}^{2}}\right)}{\Delta_{j,0}^{2}}$
$\Rightarrow$ $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}-1\geq\frac{112\left(\log\frac{2K\alpha_{k}}{\delta}+\log\log\frac{96}{\omega^{2}\Delta_{j,0}^{2}}\right)}{\omega^{2}\Delta_{j,0}^{2}}.$
The last line suggests that Algorithm 2 would output $\hat{a}_{k}$ before meeting the stop condition $N_{\hat{a}}^{\text{et}}>\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}-1$ in Line 2. ∎
Lemma 27.
For $B$ such that $\max_{a^{\prime}\in B}\mu_{a^{\prime}}\geq\mu_{j}$, Algorithm 2 must terminate no later than the end of phase $\max\{\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$.
Proof of Lemma 27.
This is a direct conclusion from Lemma 25, 26. ∎
Lemma 28.
Consider the largest bracket $B$, i.e. $B=[K]$. If $\max_{a^{\prime}\in B}\mu_{a^{\prime}}=\max_{a^{\prime}\in[K]}\mu_{a^{\prime}}<\mu_{0}$, then for phase index $k\geq\kappa^{\text{ee}}_{B}$, Algorithm 2 either takes $\hat{a}_{k}=\textsf{None}$ or takes $\hat{a}_{k}=\textsf{Not Completed}$
Proof of Lemma 28.
From the assumption $k\geq\kappa^{\text{ee}}_{B}$, by the Lemma 13, we know
$\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}})-U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})\leq\mu_{a}\leq\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}})+U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$ (25)
holds for $a\in B$ through out the phase $k$.
The analysis is similar to the Lemma 23. To quit the While loop in the Line 2, Algorithm 2 must enter one of the Lines in 2, 2, 2. By the assumption $\mu_{a}<\mu_{0}$ holds for all $a\in[K]$ and (28), we know
$\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k})$
$=$ $\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}})-C\cdot U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$
$\stackrel{{\scriptstyle C>1}}{{\leq}}$ $\hat{\mu}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}})-U(N_{a}^{\text{ee}}(\mathcal{H}^{ee}),\frac{\delta_{k}}{K})$
$\leq$ $\mu_{a}$
$<$ $\mu_{0},$
holds through out the phase $k$. The last line also suggests that the condition in Line 2 will never be satisfied in phase $k$. Thus, we can conclude $\hat{a}_{k}=\textsf{None}\text{ or }\textsf{Not Completed}$. ∎
Lemma 29.
For $B$ such that $|B|=K$ and $\max_{a^{\prime}\in B}\mu_{a^{\prime}}<\mu_{0}$, phase index $k\geq\kappa^{\text{ee}}_{B}$, for each arm $a\in B$, we have
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\cdot\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\right\}$
Proof of Lemma 29.
Similar to the idea of Lemma 24, we first turn to prove an intermediate conclusion
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$ (26)
$\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\right\}$
holds for all $k\geq\kappa^{\text{ee}}_{B}$ and $a\in B$. If the above conclusion holds, we can use induction to conclude
$N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
$\leq$ $\max\Bigg\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{113\left(\log\frac{4K}{\delta_{k-1}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}},$
$\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\Bigg\}$
$=$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big),\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}\right\}.$
which proves the Lemma. Thus, we suffice to prove the intermediate conclusion (26).
We can conduct direct calculation for each arm $a$.
$t\geq\frac{28\cdot 2^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
[…]
$\Leftrightarrow$ $\mu_{a}+2\sqrt{\frac{4\log\frac{2K(\log_{2}(2t))^{2}}{\delta_{k}}}{t}}\leq\mu_{0}$
$\Rightarrow$ $\mu_{a}+2\cdot\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}>\mu_{0}$
[…]
The last line is by the assumption $k\geq\kappa^{\text{ee}}_{B}$, which means (23) still holds.
If $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)\geq\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$, we know $\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})<\mu_{0}$ holds at the start of phase $k$. If $a=\arg\max_{a^{\prime}\in B}\text{UCB}_{a^{\prime}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})$ holds during the phase $k$, Algorithm 2 will turn to phase $k+1$ as the condition in the Line 2 holds. Thus, we can conclude Algorithm 2 will never pull arm $a$ during the phase $k$. That means if $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)\geq\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$, $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)=N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)$.
If $N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big)<\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$. We can also assert $\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})\leq\mu_{0}$ will hold if $N_{a}^{\text{ee},B}\geq\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$ holds in the futuer rounds. Similar to the above argument, if $\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})\leq\mu_{0}$, $a=\arg\max_{a^{\prime}}\text{UCB}_{a^{\prime}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{k}}{K})$ both holds, the algorithm would turn to phase $k+1$ or terminate because of meeting the condition in the Line 2. That means once $N_{a}^{\text{ee},B}\geq\frac{112\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}+1$ holds, Algorithm 2 will never pull arm $a$ in the remaining phase $k$. We can assert $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)\leq\frac{113\left(\log\frac{4K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$. ∎
Lemma 30.
For $B$ such that $|B|=K$ and $\max_{a^{\prime}\in B}\mu_{a^{\prime}}<\mu_{0}$, Algorithm 2 must terminate no later than the end of phase $\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}$.
Proof of Lemma 30.
From the calculation in the proof of Lemma 29, we know
$t\geq\frac{112\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
$\Rightarrow$ $\hat{\mu}_{a,t}+\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}>\mu_{0}$
for arm $a\in B=[K]$. Notice that for any $a\in[K]$, we have
$k\geq\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil$
$\stackrel{{\scriptstyle\mu_{0}>\mu_{1}\geq\mu_{a}}}{{\Rightarrow}}$ $\beta_{k}\geq\frac{226}{\Delta_{a,0}^{2}},\log\frac{1}{\delta_{k}}\geq\frac{96}{\Delta_{a,0}^{2}}\log 3$
$\Rightarrow$ $\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{226\log\frac{2K}{\delta_{k}}}{\Delta_{a,0}^{2}},\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{226\log\log\frac{96}{\Delta_{a,0}^{2}}}{\Delta_{a,0}^{2}}$
$\Rightarrow$ $\beta_{k}\log\frac{4K}{\delta_{k}}\geq\frac{113\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
$\stackrel{{\scriptstyle C>1}}{{\Rightarrow}}$ $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1\geq\frac{113\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{96}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}.$
That means at the end of phase $\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}$, we must have
$\text{UCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\frac{\delta_{\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\log_{3}\frac{3}{\delta}\}}}{K})\leq\mu_{0}$
holds for all $a\in[K]$. Also, not hard to see $\delta_{\max\{\kappa^{\text{ee}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}}\leq\delta_{\lceil\log_{3}\frac{3}{\delta}\rceil}<\frac{\delta}{3}$. We can conclude Algorithm 2 must terminate by the condition in Line 2, 2. ∎
C.3 Proof of Theorem 2 and 3
The full proof requires multiple lemmas to support, which motivates us to use figure 1 to summarize the overall ideas. An arrow from A to B means Lemma A is used in the proof of Lemma/Theorem B. All the Lemmas inside the orange box are the properties of Algorithm 2, which are mainly adapted from Li and Cheung (2025).
Figure 1: Overall Idea of Full Proof of Theorem 2 and 3
Before illustrating the proof of Theorem 2 and 3, we first introduce a Lemma for the random permutation, Line 1 of Algorithm 1.
Lemma 31.
Given a qualified arm index $j\in[m]$, denote bracket index $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$. We have
$\Pr(b_{j}\geq\tilde{b})\leq\frac{1}{\exp\left(\frac{j|B_{\tilde{b}-1}|}{K}\right)},\forall\tilde{b}\in[\lceil\log_{2}K\rceil+1]$
Here we denote $B_{0}=\emptyset$ and $|B_{0}|=0$. Since we take $|B_{\tilde{b}}|=\min\{2^{\tilde{b}-1},K\}$ for $\tilde{b}\in[\lceil\log_{2}K\rceil+1]$, we can rewrite the conclusion as
$\Pr(b_{j}\geq\tilde{b})\leq\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{\tilde{b}-2}\rfloor}{K}\right)},\forall\tilde{b}\in[\lceil\log_{2}K\rceil+1].$
Proof.
If $\tilde{b}=1$, the conclusion is $\Pr(b_{j}\geq 1)\leq 1$, which is obviously true. We focus on the case that $2\leq\tilde{b}\leq\lceil\log_{2}K\rceil+1$. We have
[…]
$=$ $\Pr\Big(\{\sigma(i)\}_{i=1}^{2^{\tilde{b}-2}}\cap[j]=\emptyset\Big).$
The second line is from the fact that $\tilde{b}\in[\lceil\log_{2}K\rceil+1]$, which concludes $2^{\tilde{b}-2}\leq 2^{\lceil\log_{2}K\rceil-1}\leq 2^{\log_{2}K}=K$. We can conclude
$\Pr\Big(\{\sigma(i)\}_{i=1}^{2^{\tilde{b}-2}}\cap[j]=\emptyset\Big)$
$=$ $\prod_{i=1}^{2^{\tilde{b}-2}}\frac{K-j-i+1}{K-i+1}$
$\leq$ $(\frac{K-j}{K})^{2^{\tilde{b}-2}}$
$=$ $\left(\left(1-\frac{j}{K}\right)^{\frac{K}{j}}\right)^{\frac{j\cdot 2^{\tilde{b}-2}}{K}}$
$\leq$ $\exp\left(-\frac{j\cdot 2^{\tilde{b}-2}}{K}\right).$
Given the above preparation, we are ready to prove the Theorem 2 and 3. We first prove Theorem 2.
Proof of Theorem 2.
We first show $\Pr_{\nu}(\tau<+\infty)=1$, by applying Algorithm 1. This conclusion holds no matter $\nu$ is positive or negative. Consider the bracket $\lceil\log_{2}K\rceil+1$, which is indeed $[K]$. From Lemma 30 and Lemma 27, we know
$\Pr_{\nu}(\tau=+\infty)$
$\leq$ $\Pr_{\nu}(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)$
$\leq$ $\Pr_{\nu}(\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1}=+\infty)+\Pr_{\nu}(\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}=+\infty)$
holds for any instance $\nu$. By the Lemma 17, we know $\Pr_{\nu}(\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1}=+\infty)=\Pr_{\nu}(\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}=+\infty)=0$. That means $\Pr_{\nu}(\tau=+\infty)=0$, further $\Pr_{\nu}(\tau<+\infty)=1$. This suggest that $\hat{a}$ is well defined with probability 1, and we can assert $\Pr_{\nu}(\hat{a}\in[K]\cup\{\textsf{None}\})=1$ if $\nu$ is positive or negative.
Then, we suffice to show $\Pr_{\nu}(\mu_{\hat{a}}\leq\mu_{0}\text{ or }\mu_{\hat{a}}=\textsf{None})<\delta$ for a positive instance $\nu$ and $\Pr_{\nu}(\hat{a}\in[K])<\delta$ for a negative instance $\nu$. If $\nu$ is positive, we get
$\Pr_{\nu}(\mu_{\hat{a}}\leq\mu_{0}\text{ or }\hat{a}=\textsf{None})$
$\leq$ $\Pr_{\nu}(\mu_{\hat{a}}\leq\mu_{0})+\Pr_{\nu}(\hat{a}=\textsf{None})$
$\leq$ $\Pr_{\nu}(\text{alg copy }\lceil\log_{2}K\rceil+1\text{ outputs }\hat{a}=\textsf{None})+$
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\Pr_{\nu}(\text{alg copy }b\text{ outputs }\hat{a},s.t.\mu_{\hat{a}}\leq\mu_{0})$
$\leq$ $\frac{\pi^{2}}{6}\frac{\delta}{3}+\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{\pi^{2}}{6\cdot 5}\frac{\delta}{\lceil\log_{2}K\rceil+1}$
$\leq$ $\delta.$
The second last line is by Lemma 18 and 19.
If $\nu$ is negative, we get
$\Pr_{\nu}(\hat{a}\in[K])$
$\leq$ $\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\Pr_{\nu}(\text{alg copy }b\text{ outputs }\hat{a},s.t.\mu_{\hat{a}}\leq\mu_{0})$
$\leq$ $\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{\pi^{2}}{6\cdot 5}\frac{\delta}{\lceil\log_{2}K\rceil+1}$
$<$ $\delta.$
The second last line is by Lemma 18. ∎
To prove Theorem 3, we split the proof into two cases, $\nu$ is positive or negative.
Proof of Theorem 3, when $\nu$ is positive.
Fix arbitrary $j\in[m]$ and follow the notation in Lemma 31. Denote $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$. We have $\tau=\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\tau\mathds{1}(b_{j}=b)$. Then, we can turn to find the upper bound for each $\tau\mathds{1}(b_{j}=b)$.
From the algorithm design, we know
$\tau\mathds{1}(b_{j}=b)\leq$ $(\lceil\log_{2}K\rceil+1)\tau(B_{b})\mathds{1}(b_{j}=b)$
$\leq$ $(\lceil\log_{2}K\rceil+1)\Big(\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)+\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\Big).$
Then, we are going to derive an upper bound for $\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)$ and $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\Big)$ separately.
For $\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)$, we have
$\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)$
[…]
$\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\Bigg)$
[…]
$\underbrace{\sum_{a=1}^{K}\frac{29(C+1)^{2}\left(\log\frac{4K\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}}{\delta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\mathds{1}(b_{j}=b,a\in B_{b})}_{\ddagger}.$
For $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)$, we have
$\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)$
[…]
$\leq$ $\underbrace{\frac{2(C+3)^{2}}{(C-1)^{2}}\mathds{1}(b_{j}=b)\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}|_{k=\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}_{\diamondsuit}.$
The last line is by the fact that $\frac{\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}}{\beta_{k-1}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k-1}}{\delta}}\geq 2$.
Then, to bound $\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)$ and $\mathbb{E}\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\Big)$, we suffice to consider the upper bound of the expectation of $\dagger,\ddagger,\diamondsuit$. For $\dagger$, we have
$\mathbb{E}\left[\mathds{1}(b_{j}=b)|B_{b}|\left(\beta_{\kappa^{\text{ee}}_{b}-1}\log\frac{4K}{\delta_{\kappa^{\text{ee}}_{b}-1}}+1\right)\right]$
$\leq$ $\Pr(b_{j}=b)|B_{b}|\mathbb{E}\left[\left(\beta_{\kappa^{\text{ee}}_{b}-1}\log\frac{4K}{\delta_{\kappa^{\text{ee}}_{b}-1}}+1\right)\right]$
$\leq$ $\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}\left(1+\sum_{k=1}^{+\infty}2^{k-1}\log(4K\cdot 3^{k-1})\frac{\pi^{2}}{6}\frac{1}{3^{k-1}}\right)$
$\leq$ $O(\log K)\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}.$
The second line is by the fact that $\kappa^{\text{ee}}_{b}$ is independent with $b_{j}$. The third line is by the Lemma 17, 31. For the term $\ddagger$, we have
$\mathbb{E}\left[\sum_{a=1}^{K}\frac{\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}+\log\log\frac{1}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\mathds{1}(b_{j}=b)\mathds{1}(a\in B_{b})\right]$
$\leq$ $\sum_{a=1}^{K}\frac{\log\log\frac{1}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\Pr(a\in B_{b},b_{j}=b)+$
$\sum_{a=1}^{K}\mathbb{E}\left[\frac{\log(4K)+(\kappa^{\text{ee}}_{b}+\kappa^{\text{et}}_{b}+\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil)\log 3}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right]\Pr(a\in B_{b},b_{j}=b)$
$\leq$ $\sum_{a=1}^{K}\frac{O\left((\log(4K)+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}})\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}+$
$\sum_{a=1}^{K}\frac{\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\cdot(\sum_{k=1}^{+\infty}k\frac{\pi^{2}}{6}\frac{1}{3^{k-1}}+\sum_{k=1}^{+\infty}k\frac{\pi^{2}}{6}\frac{\delta}{5^{k-1}}).$
The second step is by the fact that $\frac{1}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\leq\frac{4}{\Delta_{j,0}^{2}}$ holds for all $a,j\in[K]$ and the third step is by the Lemma 17.
From Lemma 31, we know $\Pr(b_{j}=b)\leq\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}$. What’s more, we know $\Pr(a\in B_{b})=\frac{|B_{b}|}{K}=\frac{\min\{2^{b-1},K\}}{K}$ holds for all $a\in[K]$. Thus, we can conclude
$\mathbb{E}\left[\sum_{a=1}^{K}\frac{\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\log_{2}\frac{\Theta(1)}{\Delta_{j,0}^{2}}\}}}+\log\log\frac{1}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\mathds{1}(b_{j}=b)\mathds{1}(a\in B_{b})\right]$
$\leq$ $\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\sum_{a=1}^{K}\frac{O\left((\log(4K)+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}})\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}+$
$\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\sum_{a=1}^{K}\frac{O(1)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$
$\leq$ $\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\sum_{a=1}^{K}\frac{O\left((\log(4K)+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}})\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}.$
For the expectation of $\diamondsuit$, we have
$\mathbb{E}\mathds{1}(b_{j}=b)\beta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}{\delta})$
$\leq$ $\Pr(b_{j}=b)O\left(\mathbb{E}\beta_{\kappa^{\text{ee}}_{b}}\log\frac{4K\alpha_{\kappa^{\text{ee}}_{b}}}{\delta}+\mathbb{E}\beta_{\kappa^{\text{et}}_{b}}\log\frac{4K\alpha_{\kappa^{\text{et}}_{b}}}{\delta}+\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$
$\leq$ $\Pr(b_{j}=b)O\left(\sum_{k=1}^{+\infty}\frac{\pi^{2}}{6}\frac{2^{k}}{3^{k}}\log\frac{4K\cdot 5^{k}}{\delta}+\sum_{k=1}^{+\infty}\frac{\pi^{2}}{6}\frac{2^{k}\delta}{5^{k}}\log\frac{4K\cdot 5^{k}}{\delta}+\frac{\log\frac{4K}{\delta}+\log_{2}\frac{\Theta(1)}{\Delta_{j,0}^{2}}\log 5}{\Delta_{j,0}^{2}}\right)$
$\leq$ $\Pr(b_{j}=b)O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$
Sum up all the upper bounds, we have
$\mathbb{E}\tau$
$\leq$ $O(\log K)\sum_{b=1}^{\lceil\log_{2}K\rceil+1}O(\log K)\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}+$
$O(\log K)\sum_{a=1}^{K}\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{O\left((\log(4K)+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}})\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}+$
$O(\log K)\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{(C+3)^{2}}{(C-1)^{2}}\Pr(b_{j}=b)O\left(\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$
By the Lemma 33, 34, we can conclude
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}\leq\frac{K}{j}\cdot O\Big((\log_{2}\frac{K}{j})+1\Big)$
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\leq O\Big(\frac{\log j}{j}\Big).$
Together with the fact $\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\Pr(b_{j}=b)=1$ and plug in default value $C=1.01$ specified by Algorithm 1, we have
$\mathbb{E}\tau$
$\leq$ $O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{K\log K\Big((\log_{2}\frac{K}{j})+1\Big)}{j}+\frac{\log j}{j}\sum_{a=1}^{K}\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)$
$\leq$ $O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log j\log K\Big((\log_{2}\frac{K}{j})+1\Big)}{j}\sum_{a=1}^{K}\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)$
$\leq$ $O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right).$
The second last line is by the assumption $\Delta_{i,j}<1$ for all $i,j\in[K]\cup\{0\}$. Notice the above upper bound holds for all $j\in[m]$, we can conclude
$\mathbb{E}\tau\leq O(\log K)\cdot O\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right).$
Proof of Theorem 3, when $\nu$ is negative.
Consider the bracket with size $K$, denoting this bracket as $B$. Similar to the proof of positive case, we split $\tau(B)=\tau^{\text{ee}}(B)+\tau^{\text{et}}(B)$ and prove an upper bound for $\mathbb{E}\tau^{\text{ee}}(B)$ and $\mathbb{E}\tau^{\text{et}}(B)$ separately. From Lemma 14, 29 and Lemma 30, we know
$\tau^{\text{ee}}(B)\leq$ $|Q|+\sum_{a=1}^{K}N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big)+\frac{O\left(\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}}}+\log\log\frac{1}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
$=$ $K+\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)+\sum_{a=1}^{K}\frac{O\left(\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}}}+\log\log\frac{1}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}.$
Taking expectation on both sides, we have
$\mathbb{E}\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)$
$\leq$ $\sum_{k=1}^{+\infty}|K|\beta_{k-1}\log\frac{4K}{\delta_{k-1}}\Pr(\kappa^{\text{ee}}_{B}=k)$
$\leq$ $\sum_{k=1}^{+\infty}|K|2^{k-1}\log(4K\cdot 3^{k-1})\frac{\pi^{2}}{6\cdot 3^{k-1}}$
$\leq$ $O(K\log K).$
The second last line is by the Lemma 17. In addition, we can also derive
$\mathbb{E}\log\frac{1}{\delta_{\max\{\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}}}$
$\leq$ $\mathbb{E}\log\frac{1}{\delta_{\kappa^{\text{ee}}_{B}}}+\mathbb{E}\log\frac{1}{\delta_{\kappa^{\text{et}}_{B}}}+\mathbb{E}\log\frac{1}{\delta_{\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil}}+\mathbb{E}\log\frac{1}{\delta_{\lceil\log_{3}\frac{3}{\delta}\rceil}}$
$\leq$ $\sum_{k=1}^{+\infty}k\log 3\Pr(\kappa^{\text{ee}}_{B}=k)+\sum_{k=1}^{+\infty}k\log 3\Pr(\kappa^{\text{et}}_{B}=k)+O\left(\log_{2}\frac{1}{\Delta_{1,0}^{2}}\right)+\log\frac{9}{\delta}$
$\leq$ $\sum_{k=1}^{+\infty}\frac{\pi^{2}}{6}\frac{k}{3^{k-1}}\log 3\Pr(\kappa^{\text{ee}}_{B}=k)+\sum_{k=1}^{+\infty}\frac{\pi^{2}}{6}\frac{k\delta}{5^{k-1}}\log 3+O\left(\log_{2}\frac{1}{\Delta_{1,0}^{2}}\right)+\log\frac{3}{\delta}$
$\leq$ $O\left(\log_{2}\frac{1}{\Delta_{1,0}^{2}}+\log\frac{1}{\delta}\right).$
The second last line is still by the Lemma 17. The above two inequalities suggests that
$\mathbb{E}\tau^{\text{ee}}(B)\leq$ $O(K\log K)+\sum_{a=1}^{K}O\left(\frac{\log\frac{12K(\lceil\log_{2}K\rceil+1)}{\delta}+\log\frac{1}{\Delta_{0,1}^{2}}}{\Delta_{a,0}^{2}}\right)$
$\leq$ $\sum_{a=1}^{K}O\left(\frac{\log\frac{12K(\lceil\log_{2}K\rceil+1)}{\delta}+\log\frac{1}{\Delta_{0,1}^{2}}}{\Delta_{a,0}^{2}}\right).$
Meanwhile, for $\mathbb{E}\tau^{\text{et}}(B)$, we can utilize Lemma 28 to conclude
$\mathbb{E}\tau^{\text{et}}(B)\leq$ $\mathbb{E}\sum_{k=1}^{\kappa^{\text{ee}}_{B}-1}O(1)\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}$
$\leq$ $\mathbb{E}O(1)\beta_{\kappa^{\text{ee}}_{B}-1}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{\kappa^{\text{ee}}_{B}-1}}{\delta}$
$\leq$ $O(1)\sum_{k=1}^{+\infty}\beta_{k-1}\frac{\pi^{2}}{6}\delta_{k-1}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k-1}}{\delta}$
$\leq$ $O\left(\log\frac{4K}{\delta}\right).$
Combining the upper bound of $\mathbb{E}\tau^{\text{ee}}(B)$ and $\mathbb{E}\tau^{\text{et}}(B)$, we have
$\mathbb{E}\tau\leq O(\log K)\mathbb{E}\tau(B)\leq O(\log K)O\left(\sum_{a=1}^{K}\frac{\log\frac{K}{\delta}+\log\frac{1}{\Delta_{0,1}^{2}}}{\Delta_{a,0}^{2}}\right)$
Appendix D Technical Lemma
This section is for the proof of inequalities.
Lemma 32.
Denote $\{Z_{s}\}_{s=1}^{+\infty}$ is a group of random variable fulfilling that $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(0,\Delta^{2})$ for some $\Delta>0$. Then, we have
$\Pr(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}>z)\leq\exp\left(-\frac{z^{2}}{2n\Delta^{2}}\right).$
Proof of Lemma 32.
Notice that for $\lambda=\frac{z}{n\Delta^{2}}$, we have
$\Pr\big(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}\geq s\big)$
$=$ $\Pr\big(\max_{1\leq t\leq n}\lambda\big(\sum_{s=1}^{t}Z_{s}\big)\geq\lambda s\big)$
$=$ $\Pr\left(\max_{1\leq t\leq n}\prod_{s=1}^{t}\exp\left(\lambda\big(Z_{s}\big)\right)\geq\exp\left(\lambda s\right)\right).$
Easy to see $\{\prod_{s=1}^{t}\exp\left(\lambda\big(Z_{s}\big)\right)\}_{t=1}^{+\infty}$ is a submartingale, as
$\mathbb{E}[\exp\left(\lambda\big(Z_{t}\big)\right)|z_{1},\cdots,z_{t-1}]$
$\stackrel{{\scriptstyle\text{Jensen's Inequality}}}{{\geq}}$ $\exp\left(\lambda\mathbb{E}[\big(Z_{t}\big)|Z_{1},\cdots,Z_{t-1}]\right)$
$=$ $\exp(0)=1.$
Then, we can conclude
$\Pr(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}>z)$
$\leq$ $\frac{\mathbb{E}_{\nu^{\prime}}\prod_{t=1}^{n}\exp(\lambda(Z_{t}))}{\exp(\lambda s)}$
$=$ $\frac{\exp(\frac{n\lambda^{2}\Delta^{2}}{2})}{\exp(\lambda s)}$
$=$ $\exp(-\frac{s^{2}}{2n\Delta^{2}}).$
Lemma 33.
For any $K\in\mathbb{N}$, $K\geq 2$, $j\in[K]$ we have
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}\leq\frac{K}{j}\cdot O\Big((\log_{2}\frac{K}{j})+1\Big)$
Proof.
We can conduct direct calculation.
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}$
$\stackrel{{\scriptstyle K\geq 2}}{{=}}$ $1+\frac{2}{\exp(\frac{j}{K})}+\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}$
$\leq$ $1+\frac{2}{\exp(\frac{j}{K})}+\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{2^{b-1}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$=$ $1+\frac{2}{\exp(\frac{j}{K})}+\frac{2K}{j}\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $1+\frac{2K}{j}+\frac{2K}{j}\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}.$
Then, we suffice to prove $\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}\leq O\left((\log_{2}\frac{K}{j})+1\right)$.
From the fact that $\frac{d\frac{x}{\exp(x)}}{dx}=e^{-x}-xe^{-x}=(1-x)e^{-x}$, we know $\frac{x}{\exp(x)}$ monotonically increases in the interval $(0,1)$ and decreases in the interval $(1,+\infty)$. We have
$\sum_{b=3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $\sum_{b=3}^{\lfloor\log_{2}\frac{K}{j}\rfloor+3}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}+\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $(\lfloor\log_{2}\frac{K}{j}\rfloor+1)\frac{x}{e^{x}}|_{x=1}+\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $(\lfloor\log_{2}\frac{K}{j}\rfloor+1)\frac{1}{e}+\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}.$
The second last line is by that function $\frac{x}{e^{x}}$ achieves its maximum at $x=1$. The remaining work is to prove $\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}\leq O(1)$.
Notice that for $b\geq\lceil\log_{2}\frac{K}{j}\rceil+3$, we have $\frac{j2^{b-2}}{K}\geq\frac{j2^{\log_{2}\frac{K}{j}+1}}{K}\geq 2$. Since we can conclude $\frac{x}{e^{x}}\geq 2\cdot\frac{2x}{e^{2x}}$ holds for all $x\geq 2$, we can conclude
$\frac{\frac{j\cdot 2^{b-1}}{K}}{\exp\left(\frac{j\cdot 2^{b-1}}{K}\right)}/\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}\leq\frac{1}{2}$
holds for all $b\geq\lceil\log_{2}\frac{K}{j}\rceil+3$. Thus, we have
$\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{\frac{j\cdot 2^{b-2}}{K}}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $\sum_{b=\lceil\log_{2}\frac{K}{j}\rceil+3}^{+\infty}2^{-(b-\lceil\log_{2}\frac{K}{j}\rceil-3)}\frac{\frac{j\cdot 2^{\lceil\log_{2}\frac{K}{j}\rceil+3-2}}{K}}{\exp\left(\frac{j\cdot 2^{\lceil\log_{2}\frac{K}{j}\rceil+3-2}}{K}\right)}$
$\leq$ $\frac{1}{e}\sum_{b=0}^{+\infty}2^{-b}$
$\leq$ $O(1).$
That means we complete the proof. ∎
Lemma 34.
For any $K\in\mathbb{N}$, $K\geq 2$, $j\in[K]$ we have
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}\leq O\Big(\frac{\log j}{j}\Big).$
Proof.
We can conduct direct calculation.
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}$
$\leq$ $\frac{1}{K}+\sum_{b=2}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)},\frac{2}{j}\frac{j\cdot 2^{b-2}}{K}\}$
$\leq$ $\frac{1}{K}+\sum_{b=2}^{\lceil\log_{2}\frac{K\log j}{j}\rceil+2}\frac{2}{j}\frac{j\cdot 2^{b-2}}{K}+\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}.$
We prove an upper bound for $\sum_{b=2}^{\lceil\log_{2}\frac{K\log j}{j}\rceil+2}\frac{2}{j}\frac{j\cdot 2^{b-2}}{K}$ and $\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$ separately.
First, we have
$\sum_{b=2}^{\lceil\log_{2}\frac{K\log j}{j}\rceil+2}\frac{2}{j}\frac{j\cdot 2^{b-2}}{K}$
$\leq$ $\frac{2}{j}\cdot 2\cdot\frac{j2^{b-2}}{K}|_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+2}$
$\leq$ $\frac{2}{j}\cdot 2\cdot\frac{j2^{1+\log_{2}\frac{K\log j}{j}}}{K}$
$=$ $\frac{8\log j}{j}.$
For $\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$, we first notice that when $b\geq\lceil\log_{2}\frac{K\log j}{j}\rceil+3$, we have
$\frac{j2^{b-2}}{K}\geq\frac{j2^{\lceil\log_{2}\frac{K\log j}{j}\rceil+3-2}}{K}\geq\frac{j2^{\log_{2}\frac{K\log j}{j}+1}}{K}\geq 2.$
Further, we can conclude for $b\geq\lceil\log_{2}\frac{K\log j}{j}\rceil+3$, we have $\frac{j2^{b-2}}{K}+1\leq\frac{j2^{b-2}}{K}\cdot 2=\frac{j2^{b-1}}{K}$. Thus, we have
$\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{\lceil\log_{2}K\rceil+1}\frac{1}{\exp\left(\frac{j\cdot 2^{b-2}}{K}\right)}$
$\leq$ $\sum_{b=\lceil\log_{2}\frac{K\log j}{j}\rceil+3}^{+\infty}\frac{1}{\exp(b-\lceil\log_{2}\frac{K\log j}{j}\rceil-3)}\frac{1}{\exp\left(\frac{j\cdot 2^{\lceil\log_{2}\frac{K\log j}{j}\rceil+3-2}}{K}\right)}$
$\leq$ $\frac{2}{\exp\left(\frac{j\cdot 2^{\lceil\log_{2}\frac{K\log j}{j}\rceil+3-2}}{K}\right)}$
$\leq$ $\frac{2}{\exp\left(\frac{j\cdot 2^{\log_{2}\frac{K\log j}{j}+1}}{K}\right)}$
$=$ $\frac{2}{\exp\left(\frac{j\cdot 2\cdot\frac{K\log j}{j}}{K}\right)}$
$=$ $\frac{2}{\exp\left(2\log j\right)}$
$<$ $\frac{2}{j}.$
In summary, we have proved
$\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)},\frac{\min\{2^{b-1},K\}}{K}\}$
$\leq$ $\frac{1}{K}+\frac{8\log j}{j}+\frac{2}{j}$
$\leq$ $O\Big(\frac{\log j}{j}\Big).$
Lemma 35 (Adapted Lemma 3 in Jamieson et al. (2014), Lemma D.3 in Li and Cheung (2025)).
Denote $\{X_{i}\}_{i=1}^{+\infty}$ as i.i.d $\sigma^{2}$-subgaussian random variable with true mean reward $\mu=0$. For any $\delta\in(0,1)$, we have
$\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2(\log_{2}2^{\lceil\log_{2}t\rceil^{+}})^{2}}{\delta}}\right)<\frac{\pi^{2}}{6}\delta.$
Or equivalently,
$\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}\right)<\frac{\pi^{2}}{6}\delta.$
Proof of Lemma 35.
Define $u_{k}=2^{k}$, $k\geq 1$. Define $x=\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}$, $S_{t}=\sum_{i=1}^{t}X_{i}$ and the event
$E_{k}=\left\{\max_{1\leq t\leq u_{k}}S_{t}>\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}\right\}\cup\left\{\min_{1\leq t\leq u_{k}}S_{t}<-\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}\right\},$
For $\lambda>0$, notice that
$\mathbb{E}\left[\exp(\lambda(\sum_{s=1}^{t}X_{s}))|X_{1},\cdots,X_{t-1}\right]$
$=$ $\exp(\lambda(\sum_{s=1}^{t-1}X_{s}))\mathbb{E}\exp(\lambda X_{t})$
$\geq$ $\exp(\lambda(\sum_{s=1}^{t-1}X_{s}))\exp(\mathbb{E}\lambda X_{t})$
$=$ $\exp(\lambda(\sum_{s=1}^{t-1}X_{s})).$
Take $\lambda=\frac{x}{u_{k}\sigma^{2}}$, we can conclude $\{\exp(\lambda S_{t})\}$ is a submartingale. Then,
$\Pr\left(\max_{1\leq t\leq u_{k}}S_{t}\geq x\right)$
$=$ $\Pr\left(\max_{1\leq t\leq u_{k}}\exp(\lambda S_{t})>\exp\left(\lambda x\right)\right)$
$\stackrel{{\scriptstyle*}}{{\leq}}$ $\frac{\mathbb{E}\exp(\lambda S_{u_{k}})}{\exp\left(\lambda x\right)}$
$\leq$ $\frac{\exp(\frac{u_{k}\lambda^{2}\sigma^{2}}{2})}{\exp\left(\lambda x\right)}$
$\stackrel{{\scriptstyle\lambda=\frac{x}{u_{k}\sigma^{2}}}}{{=}}$ $\exp(-\frac{x^{2}}{2u_{k}\sigma^{2}}).$
Step * is by the maximal inequality for the submartingale. Take $x=\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}$, we have
$\Pr\left(\max_{1\leq t\leq u_{k}}S_{t}\geq\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}\right)\leq\exp\left(-\log\frac{2(\log_{2}u_{k})^{2}}{\delta}\right)=\frac{\delta}{2(\log_{2}u_{k})^{2}}=\frac{\delta}{2k^{2}}$
For the part of $\Pr\left(\min_{1\leq t\leq u_{k}}S_{t}<-x\right)$, the proof is similar. We can conclude $\Pr(E_{k})\leq\frac{\delta}{k^{2}}$ and further $\Pr(\cup_{k=1}^{+\infty}E_{k})\leq\frac{\pi^{2}\delta}{6}$.
Thus,
$\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}\max\{2^{\lceil\log_{2}t\rceil},2\}\log\frac{2(\log_{2}\max\{2^{\lceil\log_{2}t\rceil},2\})^{2}}{\delta}}\right)$
$\leq$ $\Pr\left(\exists k,\max_{1\leq t^{\prime}\leq u_{k}}|\sum_{s=1}^{t^{\prime}}X_{s}|\geq\sqrt{2\sigma^{2}u_{k}\log\frac{2(\log_{2}u_{k})^{2}}{\delta}}\right)$
$\leq$ $\Pr(\cup_{k=1}^{+\infty}E_{k})\leq\frac{\pi^{2}\delta}{6}.$
Some Comments are as follows.
  • We can similarly prove $\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2\pi^{2}(\log_{2}2^{\lceil\log_{2}t\rceil^{+}})^{2}}{6\delta}}\right)<\delta$ holds for all $\delta\in(0,1)$.
  • Since $\lceil\log_{2}t\rceil^{+}\leq 1+\log_{2}t$, we have
    $\frac{\pi^{2}\delta}{6}\geq$ $\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{2\sigma^{2}2^{\max\{\lceil\log_{2}t\rceil,1\}}\log\frac{2(\log_{2}2^{\lceil\log_{2}t\rceil})^{2}}{\delta}}\right)$
    $\geq$ $\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{4\sigma^{2}t\log\frac{2(\log_{2}2t)^{2}}{\delta}}\right)$
Lemma 36 (Lemma D.2 in Li and Cheung (2025)).
For any $\Delta\in(0,1],K\geq 2,\delta\in(0,\frac{1}{2}],C\geq 1$, we can conclude
$t>\frac{28C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{16C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)\right)}{\Delta^{2}}$
$\Rightarrow$ $C\sqrt{\frac{4\log\frac{2K(\log_{2}2t)^{2}}{\delta}}{t}}<\Delta$
Proof.
By simple calculation, we can derive
$t>\frac{28C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{16C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)\right)}{\Delta^{2}}$
$\Leftrightarrow$ $2t>\frac{56C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)\right)}{\Delta^{2}}$
$\Leftrightarrow$ $2t>\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)\right)+32C^{2}\log(\frac{2K}{\delta})}{\Delta^{2}}$
$\stackrel{{\scriptstyle\log(x+y)\leq\log x+\log y,\forall x,y\geq 2}}{{\Rightarrow}}$ $2t>\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)+\frac{2K}{\delta}\right)}{\Delta^{2}}$
$\Rightarrow$ $2t>\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\left(\log\left(\frac{24C^{2}}{\Delta^{2}}\right)+\log\left(\log\frac{2K}{\delta}\right)\right)}{\Delta^{2}}$
$\Leftrightarrow$ $2t>\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}+\frac{32C^{2}\log\log\left(\frac{24C^{2}\log\frac{2K}{\delta}}{\Delta^{2}}\right)}{\Delta^{2}}$
[…]
$\Rightarrow$ $2t>\frac{8C^{2}\log\frac{2K}{\delta}+16C^{2}\log\log_{2}e+16C^{2}\log\log(2t)}{\Delta^{2}}$
$\Leftrightarrow$ $t>\frac{4C^{2}\log\frac{2K}{\delta}+8C^{2}\log\frac{\log(2t)}{\log 2}}{\Delta^{2}}$
$\Leftrightarrow$ $t>\frac{4C^{2}\log\frac{2K}{\delta}+8C^{2}\log(\log_{2}2t)}{\Delta^{2}}$
$\Leftrightarrow$ $C\sqrt{\frac{4\log\frac{2K(\log_{2}2t)^{2}}{\delta}}{t}}<\Delta.$
Lemma 37.
For any $b\geq a\geq 0$,
  • If $b\geq e^{2}$, we have $x\geq b+2a\log\log b\Rightarrow x\geq a\log\log(x)+b$.
  • If $b,a\geq e$, we have $e\leq x\leq b+a\log\log b\Rightarrow x<a\log\log(x)+b$.
The second inequality also implies $x\geq a\log\log(x)+b\Rightarrow x\geq b+a\log\log b$.
Proof.
We prove the first claim. Easy to see $\frac{d(x-a\log\log x-b)}{dx}=1-\frac{a}{x\log x}=\frac{x\log x-a}{x\log x}$. Take $x_{0}=b+2a\log\log b$, then $x_{0}\log x_{0}>x_{0}\log(e^{2}+2a\log\log e^{2})>x_{0}\log e^{2}=2x_{0}>a$. Thus $x-a\log\log(x)-b$ increases in the interval $(x_{0},+\infty)$. On the other hand, easy to check
$x_{0}-a\log\log x_{0}-b$
$=$ $b+2a\log\log b-a\log\log(b+2a\log\log b)-b$
$=$ $2a\log\log b-a\log\log(b+2a\log\log b)$
$=$ $a\left(2\log\log b-\log\log(b+2a\log\log b)\right)$
$=$ $a\left(\log(\log b)^{2}-\log\log(b+2a\log\log b)\right)$
$=$ $a\log\frac{(\log b)^{2}}{\log(b+2a\log\log b)}$
$\geq$ $a\log\frac{(\log b)^{2}}{\log(b+2b\log\log b)}$
$=$ $a\log\frac{(\log b)^{2}}{\log b+\log(1+2\log\log b)},$
Notice that
$(\log b)^{2}-\log b-\log(1+2\log\log b)$
$=$ $\log b(\log b-1)-\log(1+2\log\log b)$
$\stackrel{{\scriptstyle b\geq e^{2}}}{{\geq}}$ $\log b-\log(1+2\log\log b)$
$\geq$ $\log b-2\log\log b$
$\stackrel{{\scriptstyle x>2\log x,\forall x>0}}{{>}}$ $0,$
we can conclude $x_{0}-a\log\log x_{0}-b>0$ holds for all $x\geq b+2a\log\log b$.
Then we turn to prove the second claim. As $\frac{d(x-a\log\log x-b)}{dx}=1-\frac{a}{x\log x}=\frac{x\log x-a}{x\log x}$, we know there is at most 1 zero point of $\frac{x\log x-a}{x\log x}$ in the interval $(e,b+a\log\log b)$. Thus,
$\max_{e\leq x\leq b+a\log\log b}x-a\log\log x-b=\max\{x-a\log\log x-b|_{x=e},x-a\log\log x-b|_{x=b+a\log\log b}\}.$
Easy to see
$e-a\log\log e-b=e-b<0,$
and
$b+a\log\log b-a\log\log(b+a\log\log b)-b$
$=$ $a\log\log b-a\log\log(b+a\log\log b)$
$<$ $a\log\log b-a\log\log(b)$
$=$ $0.$
That means $\max\limits_{e\leq x\leq b+a\log\log b}x-a\log\log x-b<0$. The second conclusion is done. ∎

### v2

@@
 Abstract
-1-identification is a fundamental multi-armed bandit formulation on pure exploration.
-An agent aims to determine whether there exists a qualified arm whose mean reward is not less than a known threshold $\mu_{0}$, or to output None if it believes such an arm does not exist.
-The agent needs to guarantee its output is correct with probability at least $1-\delta$, while making expected total pulling times $\mathbb{E}\tau$ as small as possible.
-We work on 1-identification with two main contributions.
-(1) We utilize an optimization formulation to derive a new lower bound of $\mathbb{E}\tau$, when there is at least one qualified arm.
-(2) We design a new algorithm, deriving tight upper bounds whose gap to lower bounds are up to a polynomial of logarithm factor across all problem instance.
-Our result complements the analysis of $\mathbb{E}\tau$ when there are multiple qualified arms, which is an open problem left by history literature.
-Keywords 1-identification, Fixed Confidence, Pure Exploration
+The 1-identification problem is a fundamental pure-exploration problem in multi-armed bandits.
+An agent aims to determine whether there exists an arm whose mean reward exceeds a known threshold $\mu_{0}$, or to output None otherwise.
+The agent must guarantee correctness with probability at least $1-\delta$, while minimizing the expected number of arm pulls $\mathbb{E}[\tau]$.
+We study the 1-identification problem and make two main contributions.
+First, for instances with at least one qualified arm, we derive a new lower bound on $\mathbb{E}[\tau]$ via a novel optimization formulation.
+Second, we propose a new algorithm and establish upper bounds that match the lower bounds up to polynomial logarithmic factors uniformly over all instances.
+Our result complements the analysis of $\mathbb{E}\tau$ when there are multiple qualified arms, which is an open problem in the literature.
 1 Introduction
 1.1 Model & Problem Formulation
-We study the 1-identification problem [^1] This name origins from Katz-Samuels and Jamieson (2020), Li and Cheung (2025).
-In Degenne and Koolen (2019), it is also called Any Low Problem., a variant of the best arm identification problem.
-An instance of 1-identification consists of a tuple of four elements: $\nu=([K],\rho=\{\rho_{a}\}_{a\in[K]},\mu_{0},\delta)$.
-The set $[K]$ denotes the arm set $\{1,2,\cdots,K\}$, and we assume $K\geq 2$.
-For each $a\in[K]$, $\rho_{a}$ is the reward probability distribution associated with arm $a$.
-The reward distribution $\rho_{a}$ is not known to the agent for each $a$.
-By pulling arm $a$ once, the agent receives a random reward distributed as $\rho_{a}$.
-We denote $r_{a}^{\nu}:=\mathbb{E}_{R\sim\rho_{a}}[R]$ as the mean reward of arm $a$ under reward distribution $\rho_{a}$.
-We assume the random noise $R_{a}-r_{a}^{\nu}$ to be 1-sub-Gaussian.
-The real number $\mu_{0}$ is a threshold known to the agent.
-The agent’s main goal is to predict whether there is an arm $a\in[K]$ with $r_{a}^{\nu}\geq\mu_{0}$, and to identify such an arm in the positive case, while outputting “None” in the negative case.
-The quantity $\delta\in(0,1)$ is the tolerant parameter on the probability of a wrong prediction.
-The agent’s pulling strategy is specified by a sampling rule $\pi=\{\pi_{t}\}_{t=1}^{\infty}$, a stopping time $\tau$ and a recommendation rule $\hat{a}\in[K]\cup\{\textsf{None}\}$.
-When a pulling strategy is applied on a 1-identification problem instance with reward distribution $\rho$, the strategy generates a history $\{U_{0}\}\cup\{A_{t},X_{t},U_{t}\}_{t=1}^{\tau}$, where $A_{t}\in[K]$ denotes the arm pulled at $t$, $X_{t}\sim\rho_{A_{t}}$ denotes the reward received, and $U_{t}\sim\text{Unif}(0,1)$ is an internal random seed.
-Arm $A_{t}\in[K]$ is chosen contingent upon the history $H(t-1)=\{U_{0}\}\cup\{(A_{s},X_{s},U_{s})\}^{t-1}_{s=1}$, via the function $\pi_{t}(H(t-1))$.
-The agent stops pulling at the end of time step $\tau$, where $\tau$ is a stopping time with respect to the filtration $\{\sigma(H(t))\}^{\infty}_{t=1}$.
-After time step $\tau$, the agent outputs the prediction $\hat{a}\in[K]\cup\{\textsf{None}\}$.
-We require that $\hat{a}$ is measurable with respect to $\sigma(H(\tau))$, i.e.
-determined by the information $H(\tau)$.
-Outputting $\hat{a}\in[K]$ means that the agent concludes with arm $\hat{a}$ satisfying $r_{\hat{a}}^{\nu}\geq\mu_{0}$, while outputting $\hat{a}=\textsf{None}$ means that the agent concludes with no arm having mean reward $\geq\mu_{0}$.
-We say that $\nu$ is a positive instance if $\max_{a}r_{a}^{\nu}>\mu_{0}$, and say that $\nu$ is a negative instance if $\max_{a}r_{a}^{\nu}<\mu_{0}$.
-For $\Delta>0$, we define $\mathcal{S}^{\text{pos}}_{\Delta}=\{\nu:\max_{a}r_{a}^{\nu}-\mu_{0}\geq\Delta\}$, $\mathcal{S}^{\text{neg}}_{\Delta}=\{\nu:\mu_{0}-\max_{a}r_{a}^{\nu}\geq\Delta\}$, and take $\mathcal{S}^{\text{pos}}=\cup_{\Delta>0}\mathcal{S}^{\text{pos}}_{\Delta},\mathcal{S}^{\text{neg}}=\cup_{\Delta>0}\mathcal{S}^{\text{neg}}_{\Delta}$.
-We define $i^{*}(\nu)$ as the set of correct answers.
-For $\nu\in\mathcal{S}^{\text{pos}}_{\Delta}$, we have $i^{*}(\nu)=\{a:r_{a}^{\nu}\geq\mu_{0}\}$, while for instance $\nu\in\mathcal{S}^{\text{neg}}_{\Delta}$, we have $i^{*}(\nu)=\{\textsf{None}\}$.
-We don’t consider instance whose $\max r_{a}^{\nu}=\mu_{0}$ and mainly focus on analyzing PAC algorithms, with the following definitions.
-Definition 1.
-A pulling strategy is $\delta$-PAC, if for any $\delta\in(0,1),\nu\in\mathcal{S}^{\text{pos}}\cup\mathcal{S}^{\text{neg}}$, it satisfies $\Pr_{\nu}(\tau<+\infty,\hat{a}\in i^{*}(\nu))>1-\delta$.
-Appendix A.3 in Li and Cheung (2025) shows that a $\delta$-PAC algorithm can still suffer from $\mathbb{E}_{\nu}\tau=+\infty$ for some $\nu\in\mathcal{S}^{\text{pos}}$.
-Thus, we further introduce the following definition, which can be considered “Stronger $\delta$-PAC”.
-Definition 2.
-A pulling strategy is $(\Delta,\delta)$-PAC, if it is $\delta$-PAC, and for any $\Delta,\delta>0$, we have $\sup_{\nu\in\mathcal{S}^{\text{pos}}_{\Delta}\cup\mathcal{S}^{\text{neg}}_{\Delta}}\mathbb{E}_{\nu}\tau<+\infty$.
+We study the 1-identification problem [^1] The terminology originates from Katz-Samuels and Jamieson (2020); Li and Cheung (2025).
+Degenne and Koolen (2019) call it Any Low Problem., a variant of the best arm identification problem.
+An instance of 1-identification is specified by the tuple $\nu=([K],\rho=\{\rho_{a}\}_{a\in[K]},\mu_{0},\delta)$, where $K]\coloneqq\{1,2,\ldots,K\}$ denotes the set of arms with $K\geq 2$, and $\rho_{a}$ is the unknown reward distribution associated with arm $a$.
+Upon pulling arm $a$, the agent observes a random reward $R\sim\rho_{a}$.
+We denote by $r_{a}^{\nu}\coloneqq\mathbb{E}_{R\sim\rho_{a}}[R]$ the mean reward of arm $a$ under instance $\nu$, and assume that the noise $R-r_{a}^{\nu}$ is $1$-sub-Gaussian.
+The threshold $\mu_{0}\in\mathbb{R}$ is known to the agent.
+The confidence parameter $\delta\in(0,1)$ controls the allowable probability of an incorrect decision.
+The objective of the agent is to determine whether there exists an arm $a\in[K]$ such that $r_{a}^{\nu}\geq\mu_{0}$.
+In the affirmative case, the agent must identify (at least) one such arm; otherwise, it must output the symbol None.
+The 1-identification problem serves as a framework for many decision-making scenarios.
+For instance, a clinical trial screening phase where researchers must decide whether any of several new drug candidates meet a specific efficacy threshold to justify the cost of further development.
+Similarly, considering anomaly detection in power grids, an agent may need to determine if if any substation is experiencing a voltage fluctuation that exceeds safety limits.
+In both cases, the agent does not necessarily need to rank every option; rather, it must quickly identify at least one "qualified" arm or confidently conclude that none exist.
 […]
-1.2 Lit Review on 1-identification
-Before illustrating our main contribution, we first review the weakness in the history literature and we start the review by unifying the notations.
-In the following paper, we consider a reward vector $\{\mu_{a}\}_{a=1}^{K}\subset[0,1]^{K}$ satisfies that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$.
-Also, we consider $\mu_{0}\in[0,1]$ and define $\Delta_{i,j}=|\mu_{i}-\mu_{j}|$ for $\forall i,j\in[K]\cup\{0\}$.
-For an instance $\nu$, we assume $r_{a}^{\nu}=\mu_{a}$ without further explanation.
-In addition, we introduce different complexity terms occurred in the history result.
+Instance classes.
+Following Li and Cheung (2025), an instance $\nu$ is called a positive instance if $\max_{a\in[K]}r_{a}^{\nu}>\mu_{0}$, and a negative instance if $\max_{a\in[K]}r_{a}^{\nu}<\mu_{0}$.
+For $\Delta>0$, we define the margin-separated classes $\mathcal{S}^{\mathrm{pos}}_{\Delta}:=\bigl\{\nu:\max_{a}r_{a}^{\nu}-\mu_{0}\geq\Delta\bigr\}$, $\mathcal{S}^{\mathrm{neg}}_{\Delta}\coloneqq\bigl\{\nu:\mu_{0}-\max_{a}r_{a}^{\nu}\geq\Delta\bigr\}$.
+Let $\mathcal{S}^{\mathrm{pos}}\coloneqq\bigcup_{\Delta>0}\mathcal{S}^{\mathrm{pos}}_{\Delta}$, $\mathcal{S}^{\mathrm{neg}}\coloneqq\bigcup_{\Delta>0}\mathcal{S}^{\mathrm{neg}}_{\Delta}$.
+We define the set of correct answers $i^{*}(\nu)$ by $i^{*}(\nu)\coloneqq\begin{cases}\{a\in[K]:r_{a}^{\nu}\geq\mu_{0}\},&\text{if }\nu\in\mathcal{S}^{\mathrm{pos}},\\[4.0pt] \{\textsf{None}\},&\text{if }\nu\in\mathcal{S}^{\mathrm{neg}}.\end{cases}$.
+In this work, We ignore instances for which $\max_{a}r_{a}^{\nu}=\mu_{0}$.
+PAC guarantees. We adopt a PAC framework and focus on algorithms with correctness guarantees.
+Definition 1 ($\delta$-PAC).
+An algorithm is $\delta$-PAC, if for any $\nu\in\mathcal{S}^{\mathrm{pos}}\cup\mathcal{S}^{\mathrm{neg}}$, $\Pr_{\nu}\!\left(\tau<\infty,\;\hat{a}\in i^{*}(\nu)\right)\geq 1-\delta.$
+As shown in Appendix A.3 in Li and Cheung (2025), a $\delta$-PAC algorithm may still satisfy $\mathbb{E}_{\nu}[\tau]=\infty$ for some $\nu\in\mathcal{S}^{\mathrm{pos}}$.
+To rule out such pathological behavior, we introduce a stronger notion.
+Definition 2 (Strong $\delta$-PAC).
+An algorithm is Strong $\delta$-PAC if it is $\delta$-PAC and, for all $\Delta,\delta>0$, we have $\sup_{\nu\in\mathcal{S}^{\mathrm{pos}}_{\Delta}\cup\mathcal{S}^{\mathrm{neg}}_{\Delta}}\mathbb{E}_{\nu}[\tau]<\infty$.
+[…]
+1.2 Literature Review
+Before illustrating our main contribution, we review the existing literature on the 1-identification problem and highlight its current limitations.
+We begin by unifying notation across prior works in order to facilitate comparison.
+Throughout this paper, we consider a reward vector $(\mu_{a})_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{K}$, together with a known threshold $\mu_{0}\in[0,1]$.
+For any $i,j\in[K]\cup\{0\}$, we define the gap $\Delta_{i,j}\coloneqq|\mu_{i}-\mu_{j}|$.
+In the analysis of upper bound(section 3), we only consider instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$.
+[^2] The agent does not know the monotonic ordering of $\{\mu_{a}\}_{a=1}^{K}$, and this default setting is for convenience in discussions.
+In the analysis of lower bound(section 2), we further consider instances $\nu^{\prime}$ whose reward vector $r_{a}^{\nu^{\prime}}$ is not ordered.
+We next recall several instance-dependent complexity measures that appear in the literature:
 $\begin{split}H_{1}^{\text{neg}}=&\sum_{a=1}^{K}\frac{2}{\Delta^{2}_{0,a}},H^{\text{low}}_{1}=\sum_{a:\mu_{a}<\mu_{0}}\frac{2}{\Delta_{1,a}^{2}},H=\frac{2}{\Delta^{2}_{0,1}}\\ H_{1}=&\sum_{a=2}^{K}\frac{2}{\Delta^{2}_{1,a}},H_{0}=\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta^{2}_{0,a}}\\ H^{\text{u}}=&\frac{(\frac{K}{m}-1)}{\Delta^{2}},\Delta:=\min_{a:\mu_{a}>\mu_{0}}\Delta_{a,0},m:=\arg\max\{a:\mu_{a}>\mu_{0}\}\\ H(j)=&\frac{1}{j}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{j,a}^{2},\Delta_{a,0}^{2}\}}\end{split}$ (1)
-The discussion on the sample complexity focuses on two cases: instance $\nu$ is negative or positive.
-There has been abundant research achieving the tight approximation of sample complexity on a negative instance, which is similar to the Best Arm Identification.
-Regarding the lower bound of $\mathbb{E}\tau$, Theorem 5.4 in Li and Cheung (2025) asserts $\mathbb{E}\tau\geq\Omega(H_{1}^{\text{neg}}\log\frac{1}{\delta})$, if we apply any $\delta$-PAC algorithm on an instance $\nu$.
-Meanwhile, Algorithm HDoC in Kano et al.
-(2017), APGAI in Jourdan and Réda (2023), SEE in Li and Cheung (2025), are able to guarantee $\mathbb{E}\tau\leq O(H_{1}^{\text{neg}}(\log\frac{1}{\delta}+\log H_{1}^{\text{neg}}))$, ignoring some logarithm factor.
-In addition, Algorithm Sticky Track and Stop (S-TaS) in Degenne and Koolen (2019) is able to guarantee $\lim\sup\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\leq T^{*}(\nu)$, where $T^{*}(\nu)=H_{1}^{\text{neg}}$ if $\nu$ is negative and the reward follows Gaussian distribution with variance $1$.
-Contrast with the negative instance, the complexity of positive instance remains a mystery, as the proposed upper bounds for a positive instance suffer from different looseness.
-For simplicity, we consider a positive instance $\nu$ and denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and only consider reward following Gaussian distribution with variance $1$.
-We propose Table 1 to summarize the looseness in the current result, with imprecise formulation because of space limit.
-The actual upper bound existing in the history literature is not smaller than our presented formulation, and we can still observe the looseness in our presented formulation.
-Following the Appendix A.2 in Li and Cheung (2025), we put details of simplification on history upper bound in Appendix A.
+The study of sampling complexity naturally separates into two regimes, depending on the instance $\nu$ is negative or positive.
+For negative instances, the problem is closely related to the classical Best-Arm Identification problem, and the sample complexity is well studied.
+In particular, Theorem 5.4 of Li and Cheung (2025) establishes that, for any $\delta$-PAC algorithm and any negative instance $\nu$, $\mathbb{E}_{\nu}[\tau]\geq\Omega\!\left(H_{1}^{\mathrm{neg}}\log\frac{1}{\delta}\right)$.
+On the algorithmic side, HDoC Kano et al.
+(2017), APGAI Jourdan and Réda (2023), and SEE Li and Cheung (2025) guarantee $\mathbb{E}_{\nu}[\tau]\leq O(H_{1}^{\mathrm{neg}}(\log(1/\delta)+\log H_{1}^{\mathrm{neg}}))$, up to logarithmic factors.
+Moreover, for Gaussian rewards with unit variance, the Sticky Track-and-Stop (S-TaS) algorithm Degenne and Koolen (2019) achieves the asymptotically optimal guarantee $\limsup_{\delta\to 0}\frac{\mathbb{E}_{\nu}[\tau]}{\log(1/\delta)}\leq T^{*}(\nu)$, where $T^{*}(\nu)=H_{1}^{\mathrm{neg}}$ for negative instances.
+In contrast, the sample complexity of positive instances remains unclear.
+Existing upper bounds suffer from various looseness and fail to match known lower bounds except in restricted instances.
+To illustrate with simplicity, we consider a positive instance $\nu$ satisfying $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, and restrict attention to Gaussian rewards with variance 1.
+Table 1 summarizes representative non-asymptotic upper and lower bounds from the literature.
+Due to space limit, the bounds in the table are simplified; the original bounds are larger than those reported here.
+Even under these simplifications, the looseness remains apparent.
+Following Appendix A.2 of Li and Cheung (2025), we provide a detailed derivation of the simplified bounds in Appendix A.
 Table 1: Comparison of non-asymptotic upper bounds on $\mathbb{E}\tau$ for a positive instance $\nu$.
@@
 Algorithm Non-asymptotic Bound Optimality
-S-TaS, Degenne and Koolen (2019)
-Poiani et al. (2025)
+S-TaS,
+proposed by Degenne and Koolen (2019)
+analyzed by Poiani et al. (2025)
 $\mathbb{E}\tau\leq O(\frac{\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\Delta_{0,1}^{4}})$ $\times$
@@
 SEE, Li and Cheung (2025) $\mathbb{E}\tau\leq O(\frac{\log\frac{1}{\delta}}{\Delta_{0,1}^{2}}+H(1)\log\frac{K}{\Delta_{0,1}^{2}})$ $\surd_{\text{cond}}$
-Parallel-SEE, This work $\mathbb{E}\tau\leq O\Big(\log^{4}K\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)\log\frac{1}{\Delta_{j,0}^{2}}\Big)$ $\surd$
+PSEEB, This work $\mathbb{E}\tau\leq O\Big(\log K\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)\log^{4}K\log\frac{1}{\Delta_{j,0}^{2}}\right\}\Big)$ $\surd$
 LB, Katz-Samuels and Jamieson (2020) $\mathbb{E}\tau\geq\Omega(\frac{1}{m}H^{\text{low}}_{1}-\frac{1}{\Delta^{2}_{1,m+1}})$ $\times$
 LB, Li and Cheung (2025) $\mathbb{E}\tau\geq\Omega(H\log\frac{1}{\delta}+\frac{1}{m}H^{\text{low}}_{1}-\frac{1}{\Delta^{2}_{1,m+1}})$ $\surd_{\text{cond}}$
-LB, This work $\mathbb{E}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$ $\surd$
-Degenne and Koolen (2019) propose algorithm S-TaS, which provides asymptotic guarantee for both negative and positive instances.
-S-TaS is able to guarantee $\lim\sup\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\leq\frac{2}{\Delta_{0,1}^{2}}$.
-Together with their proposed lower bound $\lim\limits_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log\frac{1}{\delta}}\geq\frac{2}{\Delta_{0,1}^{2}}$, we know S-TaS is asymptotically optimal.
-Degenne and Koolen (2019) leaves non-asymptotic analysis of S-TaS as a future direction, which is complemented by Poiani et al.
+LB, This work $\mathbb{E}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right\}\right)$ $\surd$
+Degenne and Koolen (2019) propose the S-TaS algorithm, which attains asymptotically optimal guarantees for both negative and positive instances.
+In particular, they show $\limsup_{\delta\to 0}\frac{\mathbb{E}_{\nu}[\tau]}{\log(1/\delta)}\leq\frac{2}{\Delta_{0,1}^{2}}$, together with a matching lower bound $\lim\inf_{\delta\rightarrow 0}\frac{\mathbb{E}\tau}{\log(1/\delta)}\geq\frac{2}{\Delta_{0,1}^{2}}$, establishing asymptotic optimality.
+A non-asymptotic analysis is left as a future direction in Degenne and Koolen (2019) and later addressed by Poiani et al.
 (2025).
-Poiani et al.
-(2025) present the non-asymptotic upper bound in an implicit way, but we are able to derive a smaller but explicit formulation.
-The weakness of Poiani et al.
-(2025) comes from the $\delta$-independent part $\frac{K^{2}}{\Delta_{0,1}^{4}}$, which is inconsistent with any lower bound.
-Algorithm HDoC in Kano et al.
-(2017) proves an upper bound not smaller than $C^{\text{HDoC}}_{\text{pos}}:=\frac{\log(K/\delta)}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}}$, where $\epsilon=\min\Big\{\min\limits_{a\in[K]}\Delta_{0,a},\min\limits_{a\in[K-1]}\frac{\Delta_{a,a+1}}{2}\Big\}$.
-The suboptimality of $C^{\text{HDoC}}_{\text{pos}}$ consists of two parts.
-The first part is $H_{1}\log\log\frac{1}{\delta}$, with comments in Kano et al.
+The resulting non-asymptotic bound, however, contains a $\delta$-independent term $\frac{K^{2}}{\Delta_{0,1}^{4}}$, which has no counterpart in known lower bounds and leads to looseness.
+The HDoC algorithm of Kano et al.
+(2017) yields an upper bound $C^{\text{HDoC}}_{\text{pos}}:=\frac{\log(K/\delta)}{\Delta_{0,1}^{2}}+H_{1}\log\log\frac{1}{\delta}+\frac{K}{\epsilon^{2}}$, where $\epsilon=\min\{\min_{a}\Delta_{0,a},\,\min_{a}\Delta_{a,a+1}/2\}$.
+Its suboptimality arises from two sources.
+The first source is $H_{1}\log\log\frac{1}{\delta}$, which is explicitly acknowledged in Kano et al.
 (2017): “Therefore, in the case that $\ldots$ $K=\Omega(\frac{\log\frac{1}{\delta}}{\log\log\frac{1}{\delta}})$there still exists a gap between the lower bound in …and the upper bound in …”.
-The second part is $\frac{K}{\epsilon^{2}}$, which can be infinity if $\mu_{i}=\mu_{j}$ holds.
-for some $i,j\in[K]$.
-Based on HDoC, Tsai et al.
-(2024) propose Algorithm lilHDoC, which incorporate the LIL concentration event (Lemma 3 in Jamieson et al.
-(2014)) and add an extra warm-up stage.
-After adding these designs, the proposed algorithm lilHDoC still suffers from the same sub-optimality as HDoC.
-Kano et al.
-(2017) design Algorithm APGAI to address Good Arm Identification, aiming to output all the arms whose reward is above a given threshold $\mu_{0}$.
-Though it is not the main focus, they also propose an upper bound for 1-identification.
-The weakness of upper bound $O(H_{0}(\log\frac{K}{\delta}))$ mainly comes from $H_{0}$.
-Easy to see $H_{0}$ is proportional to $K$, suggesting that the coefficient of $\log\frac{1}{\delta}$ is suboptimal compared to the others.
-What’s more, if there exists an arm $a$ such that $\mu_{a}=\mu_{0}$, we have $H_{0}=\infty$, suggesting that the upper bound is vacuous.
-Katz-Samuels and Jamieson (2020) work on $\epsilon$-good identification and 1-identification simultaneously.
-They define a new concept Unverifiable Sample Complexity and prove upper bounds for it.
-Though Unverifiable Sample Complexity is different with Sample Complexity in this paper, we can still adapt their result, as discussed by Appendix A.1 in Li and Cheung (2025).
-The suboptimality of its upper bound consists of two points.
-First, its upper bound relies on $\log\frac{1}{\delta}\log\log\frac{1}{\delta}$ instead of $\log\frac{1}{\delta}$.
-Second, the complexity term $H^{u}$ relies on the minimum gap between a qualified mean reward and the threshold $\mu_{0}$, which can be much smaller than $\Delta_{0,1}$.
-In the case $m$ is much smaller than $K$ and $\Delta<\Delta_{0,1}$, this upper bound is much larger than their proposed lower bound.
-Li and Cheung (2025) design Algorithm SEE, splitting the upper bound of $\mathbb{E}\tau$ into two parts which are $\delta$-dependent and $\delta$-independent.
-They prove both upper and lower bounds of $\mathbb{E}\tau$ with explicit formulation.
-They show when $m=1$, i.e.
-there is only one arm whose reward is greater than $\mu_{0}$, the gap between their proposed upper and lower bounds are up to a polynomial of $\log K,\log\frac{1}{\Delta_{0,1}^{2}}$, suggesting the optimality.
-However, they also comment $H(1)\log\frac{K}{\Delta_{0,1}^{2}}$, which is $\delta$-independent part, might be improved if $m>1$.
-They leave the possible improvement as an open direction.
+The second source is $\frac{K}{\epsilon^{2}}$, which is infinite when $\epsilon=0$.
+The lilHDoC algorithm Tsai et al.
+(2024), despite incorporating LIL-based confidence bounds and an additional warm-up phase, inherits the same fundamental limitations.
+The APGAI algorithm Jourdan and Réda (2023) is originally designed for Good Arm Identification, aiming to output all the arms whose reward is above a given threshold $\mu_{0}$.
+But the analysis in Jourdan and Réda (2023) also implies an upper bound $O(H_{0}\log(K/\delta))$ for 1-identification.
+This bound is loose primarily due to the dependence on $H_{0}$, which might scale linearly with $K$, leading to a suboptimal coefficient of $\log(1/\delta)$.
+Moreover, if there exists an arm with $\mu_{a}=\mu_{0}$, then $H_{0}=\infty$.
+In this case, the bound is vacuous.
+Katz-Samuels and Jamieson (2020) study $\epsilon$-good identification and 1-identification via the notion of Unverifiable Sample Complexity.
+Although this notion differs from the sampling complexity considered here, their results can be adapted as discussed in Appendix A.1 of Li and Cheung (2025).
+The resulting upper bound is suboptimal for two reasons: it scales as $\log(1/\delta)\log\log(1/\delta)$ rather than $\log(1/\delta)$, and it depends on the complexity term $H^{\mathrm{u}}$, which is governed by the smallest gap above $\mu_{0}$ and may be much smaller than $\Delta_{0,1}$.
+When $m\ll K$ and $\Delta\ll\Delta_{0,1}$, this leads to a significant gap compared to the lower bound.
+Li and Cheung (2025) propose the SEE algorithm and derive explicit upper and lower bounds, decomposing $\mathbb{E}_{\nu}[\tau]$ into $\delta$-dependent and $\delta$-independent components.
+They show that when $m=1$, i.e.
+there is only one arm whose reward is greater than $\mu_{0}$, the gap between upper and lower bounds is at most polynomial in $\log K$ and $\log(1/\Delta_{0,1}^{2})$, implying near-optimality.
+For $m>1$, however, they comment the $\delta$-independent term $H(1)\log\!\bigl(K/\Delta_{0,1}^{2}\bigr)$ may be improvable, which is explicitly left as an open problem.
 1.3 Main Contribution
-In this paper, we develop novel upper and lower bounds for the 1-identification problem. Regarding the lower bound, we conclude
+In this paper, we develop novel upper and lower sample complexity bounds for the 1-identification problem.
+Our results characterize the sample complexity in both positive and negative instances and close the gap for instances with multiple qualified arms.
+We first establish a new lower bound for positive instances.
 Theorem 1.
-For any $\delta$-PAC alg and any reward vector $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, any tolerance level $\delta$ such that $\delta<\min\{\frac{1}{10^{-8}},\frac{1}{64m^{2}}\}$, we can find a permutation $\sigma:[K]\rightarrow[K]$ and define a positive instance $\nu$ with $r_{\sigma(a)}^{\nu}=\mu_{a},a\in[K]$ with unit-variance Gaussian reward, such that $\mathbb{E}_{\text{alg},\nu}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$.
-Regarding the upper bound, we merge and adapt the idea in Katz-Samuels and Jamieson (2020) and Li and Cheung (2025) and design a new algorithm Parallel Sequential Exploration Exploitation on Brackets(Algorithm 1, PSEEB).
-We can prove the following two theorems.
+For any $\delta$-PAC alg and any reward vector $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, any tolerance level $\delta$ such that $\delta<\min\{10^{-8},\frac{1}{64m^{2}}\}$, we can find a permutation $\sigma:[K]\rightarrow[K]$ and define a positive instance $\nu$ with $r_{\sigma(a)}^{\nu}=\mu_{a},a\in[K]$ with unit-variance Gaussian reward, such that $\mathbb{E}_{\text{alg},\nu}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right\}\right)$.
+Theorem 1 is our main contribution, achieving nearly joint optimality on both $\delta$-dependent and $\delta$-independent term.
+To the best of our knowledge, the “Simulator” framework Simchowitz et al.
+(2017); Katz-Samuels and Jamieson (2020) was the only known method for deriving $\delta$-independent lower bounds.
+However, results in Katz-Samuels and Jamieson (2020) are only tight under restricted settings ($\mu_{1}=\dots=\mu_{m}>\mu_{0}$), without joint optimality on both parts.
+The bottleneck is constructing a “Simulator” that depends on the sampling path and the reward vector.
+Our proposed method reduces this complex construction to finding a feasible solution for a convex optimization problem, which not only significantly simplifies the proof burden but also yields a tighter, jointly optimal result.
+We provide sketch proof in Section 2 and full proof in Appendix B.
+On the algorithmic side, we combine and refine ideas from Katz-Samuels and Jamieson (2020); Li and Cheung (2025) to propose a new algorithm, Parallel Sequential Exploration–Exploitation on Brackets (Algorithm 1, abbreviated as PSEEB).
+We establish the following guarantees.
 Theorem 2.
@@
 Apply Algorithm 1 to a 1-identification instance $\nu$, we have
-$\mathbb{E}_{\nu}\tau\leq\begin{cases}O\left(\log K\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\log^{3}K(\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}})H(j)\right)&\nu\text{ is positive}\\ O\left((\log K)H_{1}^{\text{neg}}\log\frac{H_{1}^{\text{neg}}}{\delta}\right)&\nu\text{ is negative}\end{cases}$
-combining Theorem 2 and 3, we know algorithm 1 is $(\Delta,\delta)$-PAC.
-What’s more, the gap between Theorem 1 and 3 is up to a polynomial of $(\log K,\log\{\frac{1}{\Delta_{0,a}^{2}}\}_{a=1}^{m})$, suggesting that our proposed algorithm 1 is nearly optimal for all 1-identification instances, and $\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+H(j)$ is an appropriate complexity measure.
-Compared to history literature, our proposed Algorithm 1 complement one of the open direction left in Li and Cheung (2025), by deriving a new upper bound for $\mathbb{E}\tau$ when there are multiple qualified arms.
+$\mathbb{E}_{\nu}\tau\leq\begin{cases}O\left(\log K\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\log^{3}K(\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}})H(j)\right\}\right)&\nu\text{ is positive}\\ O\left((\log K)H_{1}^{\text{neg}}\log\frac{H_{1}^{\text{neg}}}{\delta}\right)&\nu\text{ is negative}\end{cases}$
+Combining Theorems 2 and 3, we conclude that Algorithm 1 is Strong $\delta$-PAC.
+Moreover, the gap between the lower bound in Theorem 1 and the upper bound in Theorem 3 is at most polynomial in $\log K$ and $\{\log(1/\Delta_{0,a}^{2})\}_{a=1}^{m}$.
+This shows that PSEEB is nearly optimal for all 1-identification instances and identifies $\min_{1\leq j\leq m}\left\{\frac{\log(1/\delta)}{\Delta_{j,0}^{2}}+H(j)\right\}$ as the appropriate instance-dependent complexity measure on positive instances.
+Compared to prior work, our results resolve an open problem posed in Li and Cheung (2025) by providing the first nearly optimal upper bound for instances with multiple qualified arms.
 1.4 Literature Review on Other Related Works
-If we consider the given threshold $\mu_{0}$ as an arm with deterministic reward, we can address 1-identification by solving a Best Arm Identification(BAI) problem Even-Dar et al.
-(2002), Gabillon et al.
-(2012), Jamieson et al.
-(2013), Kalyanakrishnan et al.
-(2012), Karnin et al.
-(2013), Jamieson and Nowak (2014), Chen and Li (2015).
-However, this is an inefficient solution, suggested by the asymptotic result in Degenne and Koolen (2019), Garivier and Kaufmann (2016) and non-asymptotic result in Li and Cheung (2025), Kano et al.
-(2017).
-The reason is 1-identification doesn’t care about an exact comparison between arms with high confidence level, and we should not pull arms as frequently as BAI to distinguish two arms.
-Besides 1-identification, there are other variants of pure exploration with multiple answers, as mentioned by Degenne and Koolen (2019).
-The agent needs to find all the arms Mason et al.
-(2020) in $\{a:\mu_{a}\geq\mu_{1}-\epsilon\}$ or only one arm Even-Dar et al.
-(2002), Gabillon et al.
-(2012), Kalyanakrishnan et al.
-(2012), Kaufmann and Kalyanakrishnan (2013), Katz-Samuels and Jamieson (2020).
-As discussed in Katz-Samuels and Jamieson (2020), if the gap $\mu_{1},\mu_{0}$ are known and $\mu_{1}>\mu_{0}$, by taking $\epsilon=\Delta_{0,1}$, the answer set of $\epsilon$-Good Arm Identification and 1-identification are the same.
-We can adopt same pulling strategy and recommendation rule.
-But it is unclear about the relationship if there are unknown.
-What’s more, Locatelli et al.
-(2016), Mukherjee et al.
-(2017) work on Thresholding Bandits, which aims to output a set of arms whose reward are above a given $\mu_{0}$, within fixed budget.
-The optimal set might not be unique under their proposed problem setting.
-In this paper, we adopt the idea of parallel algorithm copies, which also exists in the history literature.
-Chen and Li (2015), Chen et al.
-(2017), Katz-Samuels and Jamieson (2020) all require a main algorithm to manage different algorithm copies.
-Katz-Samuels and Jamieson (2020) allocate $2^{\ell}$ time rounds to a copy each time with increasing index $\ell$.
-Chen and Li (2015), Chen et al.
-(2017) maintain an overall round index $t\in\mathbb{N}$ and further assign algorithm index to each copy.
-They execute an algorithm copy only when its algorithm index divide $t$.
-In our work, we apply round-robin procedure to execute each algorithm copy.
-2 Lower Bound of Sample Complexity, for a Positive Instance
-In this section, we continue to work on $\{\mu_{a}\}_{a=1}^{K}$ such that $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$.
-We assume each reward distribution follows a unit-variance Gaussian distribution in our lower bound analysis.
-For an instance $\nu^{\prime}$ and a permutation $\sigma:[K]\rightarrow[K]$, represented as a bijection from $[K]$ to $[K]$, we define the instance $\sigma(\nu^{\prime})$ as $r^{\sigma(\nu^{\prime})}_{\sigma(a)}=r_{a}^{\nu^{\prime}}$.
-We first introduce the definition of a symmetric algorithm.
-Definition 3.
-An algorithm $\widetilde{\text{alg}}$ is said to be symmetric if the following statement holds.
-For any instance $\nu$ with reward vector $\{\mu_{a}\}_{a=1}^{K}$, any sequence $\{a_{t}\}_{t=1}^{T}\in[K]^{T}$, natural number $T\geq 1$, and two different permutations $\sigma^{\prime},\sigma^{\prime\prime}$, the algorithm $\widetilde{\text{alg}}$ satisfies
-$\Pr_{\sigma^{\prime}(\nu),\tilde{\text{alg}}}(\forall t\in[T],A_{t}=\sigma^{\prime}(a_{t}),\tau>T)=\Pr_{\sigma^{\prime\prime}(\nu),\tilde{\text{alg}}}(\forall t\in[T],A_{t}=\sigma^{\prime\prime}(a_{t}),\tau>T).$
+If the threshold $\mu_{0}$ is viewed as an additional arm with deterministic reward, then 1-identification can be reduced to a standard Best-Arm Identification (BAI) problem Even-Dar et al.
+(2002); Gabillon et al.
+(2012); Jamieson et al.
+(2013); Kalyanakrishnan et al.
+(2012); Karnin et al.
+(2013); Jamieson and Nowak (2014); Chen and Li (2015).
+However, this reduction is known to be inefficient.
+Both asymptotic analyses Degenne and Koolen (2019); Garivier and Kaufmann (2016) and non-asymptotic results Li and Cheung (2025); Kano et al.
+(2017) show that such a formulation incurs unnecessary sampling.
+Intuitively, 1-identification does not require high-confidence pairwise comparisons among all arms, and therefore does not necessitate the intensive sampling in BAI Algorithms.
+Beyond 1-identification, Degenne and Koolen (2019) research on several pure-exploration variants with multiple correct answers.
+These include identifying all arms in $\{a:\mu_{a}\geq\mu_{1}-\epsilon\}$ Mason et al.
+(2020), or identifying a single arm from this set Even-Dar et al.
+(2002); Gabillon et al.
+(2012); Kalyanakrishnan et al.
+(2012); Kaufmann and Kalyanakrishnan (2013); Katz-Samuels and Jamieson (2020).
+As observed in Katz-Samuels and Jamieson (2020), when the gap $\Delta_{0,1}$ is known and $\mu_{1}>\mu_{0}$, the answer sets of $\epsilon$-good arm identification and 1-identification coincide by setting $\epsilon=\Delta_{0,1}$, allowing the same sampling and recommendation rules.
+When the gaps are unknown, however, the precise relationship between these problems remains unclear.
+In addition, thresholding bandits Locatelli et al.
+(2016); Mukherjee et al.
+(2017) aim to identify arms with mean value above $\mu_{0}$ under a fixed budget.
+The correct output might not be unique under their proposed problem setting.
+Kano et al.
+(2017) also target at identifying all the arms with mean value above $\mu_{0}$, but they focus on the fixed confidence setting.
+Since Kano et al.
+(2017) provide an upper bound for each outputting time index, their result can be generalized to 1-identification.
+Finally, our algorithmic design builds on the idea of running multiple algorithm copies in parallel, which has appeared in prior work.
+In particular, Chen and Li (2015); Chen et al.
+(2017); Katz-Samuels and Jamieson (2020) employ a master procedure to manage multiple executions of a base algorithm.
+The approach of Katz-Samuels and Jamieson (2020) allocates $2^{\ell}$ rounds to the execution indexed by $\ell$, while Chen and Li (2015); Chen et al.
+(2017) schedule parallel executions according to a global round index $t\in\mathbb{N}$, activating an execution when its index divides $t$.
+In contrast, our method adopts a simple round-robin scheduling scheme, which enables a balanced execution of parallel algorithm instances.
+2 Lower Bound of Sample Complexity, for Positive Instances
+In this section, we continue to work with a reward vector $\{\mu_{a}\}_{a=1}^{K}$ satisfying $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and consider the instance $\nu$ defined by $r_{a}^{\nu}=\mu_{a}$.
+Throughout the lower-bound analysis, we assume unit-variance Gaussian rewards.
+For any instance $\nu^{\prime}$ and any permutation $\sigma:[K]\to[K]$ (a bijection), we define the permuted instance $\sigma(\nu^{\prime})$ by $r^{\sigma(\nu^{\prime})}_{\sigma(a)}\coloneqq r_{a}^{\nu^{\prime}},\forall a\in[K]$.
+We begin by introducing a symmetry notion that allows us to restrict attention to algorithms that do not prefer specific arm labels.
+Definition 3 (Symmetric algorithm).
+An algorithm $\widetilde{\textnormal{alg}}$ is symmetric if for any instance $\nu$ with reward vector $\{\mu_{a}\}_{a=1}^{K}$, any sequence $\{a_{t}\}_{t=1}^{T}\in[K]^{T}$, any integer $T\geq 1$, and any two permutations $\sigma^{\prime},\sigma^{\prime\prime}:[K]\to[K]$, it holds that
+$\Pr_{\sigma^{\prime}(\nu),\,\widetilde{\textnormal{alg}}}\!\bigl(\forall t\in[T],\,A_{t}=\sigma^{\prime}(a_{t}),\,\tau>T\bigr)=\Pr_{\sigma^{\prime\prime}(\nu),\,\widetilde{\textnormal{alg}}}\!\bigl(\forall t\in[T],\,A_{t}=\sigma^{\prime\prime}(a_{t}),\,\tau>T\bigr).$
 Intuitively, under a symmetric algorithm $\widetilde{\text{alg}}$, the reward distribution of the pulled arm $A_{t}$ at time $t$ is invariant under any permutation on the arms’ indexes.
-In particular, for a symmetric algorithm $\widetilde{\text{alg}}$, we have $\mathbb{E}_{\sigma^{\prime}(\nu),\tilde{\text{alg}}}\tau=\mathbb{E}_{\sigma^{\prime\prime}(\nu),\tilde{\text{alg}}}\tau$ for any two permutations $\sigma^{\prime},\sigma^{\prime\prime}$.
-A similar definition also appears in Definition 12 of Simchowitz et al.
-(2017), which further requires the distribution of output $\hat{a}$ is also invariant regarding the distribution.
-This extra requirement is unnecessary in our analysis.
-By the following lemma, it suffices to consider only symmetric algorithms for establishing our lower bound.
+In particular, we have $\mathbb{E}_{\sigma^{\prime}(\nu),\,\widetilde{\textnormal{alg}}}[\tau]=\mathbb{E}_{\sigma^{\prime\prime}(\nu),\,\widetilde{\textnormal{alg}}}[\tau],\text{for permutations }\sigma^{\prime},\sigma^{\prime\prime}$.
+A similar definition also appears in (Simchowitz et al., 2017, Definition 12), which further requires the invariance of output distribution; this extra requirement is not needed here.
+The following lemma shows that it suffices to consider symmetric algorithms for lower bounds.
 Lemma 4.
-Denote $S$ as the set of all permutations of $[K]$.
-Given any $\delta$-PAC algorithm alg, we can always find a $\delta$-PAC symmetric algorithm $\widetilde{\text{alg}}$, such that for any instance $\nu^{\prime}$, we have $\max\limits_{\sigma\in S}\mathbb{E}_{\text{alg},\sigma(\nu^{\prime})}\tau\geq\mathbb{E}_{\tilde{\text{alg}},\nu^{\prime}}\tau$.
-Full proof of Lemma 4 is in Appendix B.1.
-We can shuffle all arms before applying any alg to create a symmetric algorithm $\widetilde{\text{alg}}$.
-Given Lemma 4, to prove Theorem 1, we suffice to prove the following Theorem.
+Let $S$ denote the set of all permutations of $[K]$.
+For any $\delta$-PAC algorithm alg, there exists a $\delta$-PAC symmetric algorithm $\widetilde{\textnormal{alg}}$ such that for every instance $\nu^{\prime}$, $\max\limits_{\sigma\in S}\mathbb{E}_{\textnormal{alg},\,\sigma(\nu^{\prime})}[\tau]\geq\mathbb{E}_{\widetilde{\textnormal{alg}},\,\nu^{\prime}}[\tau]$.
+The proof of Lemma 4 is in Appendix B.1.
+It is obtained by randomizing a permutation of arm labels prior to executing alg.
+Given Lemma 4, to prove Theorem 1, it suffices to establish the following lower bound for symmetric algorithms.
 Theorem 5.
-Suppose $\delta<\min\{10^{-8},1/(64m^{2})\}$. For any $\delta$-PAC symmetric algorithm, and for any positive instance $\nu$ with $m\geq 1$ eligible arms, we have
-$\mathbb{E}\tau\geq\Omega\left(\min_{1\leq j\leq m}\frac{\log(1/\delta)}{\Delta_{j,0}^{2}}+\frac{1/j}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$
-The assumption $\delta<\min\{10^{-8},1/(64m^{2})\}$ is to ensure $\delta\cdot m^{2}$ is upper bounded by a non-trivial absolute constant.
+For any symmetric and $\delta$-PAC alg and any reward vector $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ satisfying that $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, any tolerance level $\delta$ such that $\delta<\min\{10^{-8},\frac{1}{64m^{2}}\}$, define a positive instance $\nu$ with $r_{a}^{\nu}=\mu_{a},a\in[K]$ with unit-variance Gaussian reward.
+We have $\mathbb{E}_{\text{alg},\nu}\tau\geq\Omega\left(\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right\}\right)$.
+The assumption $\delta<\min\{10^{-8},1/(64m^{2})\}$ is to ensure $\delta\cdot m^{2}$ is upper bounded by an absolute constant.
 The upper bound eliminates the possibility of being $\delta$-PAC by outputting a uniformly chosen arm without any arm pulls.
 Sketch Proof of Theorem 5.
-In what follows, we present the main steps to prove Theorem 5.
-The full proof is in Appendix B.2, which consists of the statements and proofs of auxiliary lemmas.
-The main idea is to establish Lemma 6,
+We outline the main steps; the full proof is in Appendix B.2. The main idea is to establish Lemma 6,
 Lemma 6.
-Assume $\delta<10^{-8}$, $\frac{1}{8m}\geq\sqrt{\delta}$, we can assert $\mathbb{E}_{\nu}\tau$ is lower bounded by the $\frac{1}{3200}$ times optimal value of the following optimization problem
+Assume $\delta<\min\{10^{-8},1/(64m^{2})\}$, for any symmetric and $\delta$-PAC alg, $\mathbb{E}_{\nu,\text{alg}}\tau$ is lower bounded by the $\frac{1}{3200}$ times optimal value of the following optimization problem
 $\min_{v,\{p_{j}\}_{j=1}^{m}}\quad$ $v$ (2)
@@
 $v\geq\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}\quad\forall j\in[m].$ (C5)
-If Lemma 6 holds, we first prove the above optimization problem (2) is convex.
-Then, we turn to construct a feasible solution to the dual problem, and prove the objective value of this feasible solution is lower bounded by $\Omega\left(\min\limits_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right)$.
+Given Lemma 6, we proceed to prove the Theorem 1 as follows.
+We first prove the above optimization problem (2) is convex.
+Then, we turn to construct an explicit feasible dual solution, and prove the objective value of this feasible solution is lower bounded by $\Omega\left(\min\limits_{1\leq j\leq m}\left\{\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{H(j)}{\log^{2}(m+1)}\right\}\right)$.
 By the strong duality, we complete the proof of Theorem 5.
 Denote event $\mathcal{E}_{j}=\left\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right\}$.
-To prove Lemma 6, we are going to show taking $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ for $j\in[m]$ is a feasible solution to (2).
-The overall idea is to apply the Transportation Equality(Lemma 18 in Kaufmann et al.
-(2016)) to prove the relationship between $\Pr_{\nu}(\mathcal{E}_{j})$ and $\Pr_{\nu^{\prime}}(\mathcal{E}_{j})$ for some alternative instances $\nu^{\prime}$, further the relationship between $3200\mathbb{E}_{\nu}\tau$ and $\{\Pr_{\nu}(\mathcal{E}_{j})\}_{j=1}^{m}$.
+To prove Lemma 6, we show taking $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ for $j\in[m]$ is a feasible solution to (2).
+The argument repeatedly applies the transportation (change-of-measure) inequality (Kaufmann et al., 2016, Lemma 18) to relate $\Pr_{\nu}(\mathcal{E}_{j})$ to probabilities under suitable alternative instances.
 We prove the feasibility of $v=3200\mathbb{E}_{\nu}\tau$, $p_{j}=\Pr_{\nu}(\mathcal{E}_{j})$ by going through the constraint (C1), (C2), (C3), (C4) (C5).
-For (C1), (C2), they are trivial. As the expected pulling times must be non-negative and the probability must be in the interval $[0,1]$.
+For (C1), (C2), they are evident. As the expected pulling times must be non-negative and the probability must be in the interval $[0,1]$.
 For (C3), notice that $\mathds{1}(\tau=+\infty)+\mathds{1}(\tau<+\infty,\hat{a}\in[K]\cup\{\textsf{None}\})=1$, we have $\sum_{j=1}^{m}\mathds{1}(\hat{a}=j)\leq 1$, further $\sum_{j=1}^{m}\mathds{1}(\mathcal{E}_{j})\leq 1$.
 This observation concludes
-$3200\mathbb{E}\tau\geq$ $\sum_{j=1}^{m}\mathbb{E}\tau\mathds{1}(\mathcal{E}_{j})\geq\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr_{\nu}(\mathcal{E}_{j}).$
-For (C4), we start by proving the inequality (3).
-$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])\geq 1-5\delta.$ (3)
-To prove (3), we consider two cases, $\{a:\mu_{a}=\mu_{0}\}$ is empty or not. The $\delta$-PAC requirement suggests that (3) must hold for both cases.
+$3200\mathbb{E}\tau\geq\mathbb{E}\tau\geq\sum_{j=1}^{m}\mathbb{E}\tau\mathds{1}(\mathcal{E}_{j})\geq\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr_{\nu}(\mathcal{E}_{j}).$
+For (C4), we start by proving
+$\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])\geq 1-5\delta,$ (3)
+which must hold regardless of whether $\{a:\mu_{a}=\mu_{0}\}$ is empty. The main idea is to utilize the assumption that the algorithm is $\delta$-PAC.
 Then, we are ready to prove (C4) by contradiction.
-Denote $\bar{\mathcal{E}_{j}}=\{N_{j_{0}}(\tau)<\frac{\log\frac{1}{\delta}}{100\Delta_{j_{0},0}^{2}},\hat{a}=j_{0}\}$.
+Denote $\bar{\mathcal{E}_{j}}=\{N_{j}(\tau)<\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\}$.
 If $\sum_{j=1}^{m}\Pr_{\nu}(\mathcal{E}_{j})<\frac{1}{2}$, together with (3), we can find a $j_{0}\in[m]$ such that
@@
 We consider an alternative instance $\nu^{\prime}_{j_{0}}(\epsilon)$ with $r_{a^{\prime}}^{\nu^{\prime}_{j_{0}}}=\begin{cases}\mu_{0}-\epsilon&a^{\prime}=j_{0}\\ r_{a^{\prime}}^{\nu}&else\end{cases}$.
-The $\delta$-PAC requirement of algorithm conclude
+The $\delta$-PAC requirement of algorithm means that
 $\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})<\delta$ (5)
 Combining (5),(4), we have $\log\frac{\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)}{\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})}\geq\frac{1}{2}\log\frac{1}{\delta}$ holds for all $\epsilon>0$.
-Meanwhile, Lemma 8, which utilizes the Transportation Equality(Lemma 18 in Kaufmann et al.
-(2016)), suggests that
+Meanwhile, Lemma 9, which utilizes the Transportation Equality (Kaufmann et al., 2016, Lemma 18), suggests that
 $\log\frac{\Pr_{\nu}\left(\bar{\mathcal{E}}_{j_{0}}\right)}{\Pr_{\nu^{\prime}_{j_{0}}(\epsilon)}(\bar{\mathcal{E}}_{j_{0}})}$
@@
 $\mathbb{E}_{\nu}N_{a}(\tau)\geq\frac{1}{4}T_{j}(a)\Pr_{\nu}\left(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\right).$ (7)
-for each $j\in[m],a\in[K]$, where $T_{j}(a)=\lceil\frac{1}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\Big(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}\Big)}\rceil$.
-If Lemma 11 holds, we can sum up (7) for each $a\in[K]$ and use the fact that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$, concluding (C5) holds.
+for each $j\in[m],a\in[K],j\neq a$, where $T_{j}(a)=\lceil\frac{1}{200\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\Big(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}\Big)}\rceil$.
+Given Lemma 12 holds, sum up (7) for each $a\in[K]$ and use the fact that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$, we prove (C5).
 Then, the remaining work is to prove (7).
 In the following, we present the sketch idea, which highly relies on the definition of $T_{j}(a)$.
-For $n\in\mathbb{N}$, denote
-$S_{j,a}^{n}=$ $\{t:N_{j}(t)+N_{a}(t)=n,A_{t+1}\in\{j,a\}\}$
-$\tau_{n}^{j}(a)=$ $\min S_{j,a}^{n}$
-For any sample path $\{U_{s},A_{s},X_{s}\}_{s=1}^{\tau}$, $S_{j,a}^{n}$ contains at most one element, which is the round that an agent is going to pull arm $a$ or $j$.
-We suffice to focus on these rounds.
-In the case that $S_{j,a}^{n}=\emptyset$, we take $\tau_{n}^{j}(a)=+\infty$, which means $\tau_{n}^{j}(a)\geq\tau$.
-Since $A_{t+1}$ only depends history $H_{t}$, we know $\tau_{n}^{j}(a)$ is a stopping time.
-Regarding $\tau_{n}^{j}(a)$, we are going to prove two inequalities
-$\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)\geq\Pr_{\nu}(\mathcal{E}_{j}),\forall n\leq T_{j}(a).$ (8)
-$\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau)\geq\frac{1}{4}\Pr_{\nu}(\tau_{n}^{j}(a)<\tau),\forall n\leq T_{j}(a).$ (9)
-If (8) (9) both hold, from the fact that we only pull arm a at $\tau_{n}^{j}(a)+1$, we can conclude
-$\mathbb{E}_{\nu}N_{a}(\tau)\geq$ $\sum_{n=0}^{T_{j}(a)}\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<+\infty)$
-$\geq$ $\sum_{n=0}^{T_{j}(a)}\frac{1}{4}\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$
-$\geq$ $\frac{1}{4}\sum_{n=0}^{T_{j}(a)}\Pr_{\nu}(\mathcal{E}_{j})$
-$=$ $\frac{1}{4}T_{j}(a)\Pr_{\nu}(\mathcal{E}_{j}).$
-The second step is by (9) and the third step is by (8). Thus, to prove (7), we suffice to show (8) (9) are both correct.
-For (8), we notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,0}^{2}}$, further $T_{j}(a)<\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}$.
+For $n\in\mathbb{N}$, we define $\tau_{n}^{j}(a)$ by looking for an time index $t$, such that $N_{j}(t)+N_{a}(t)=n,A_{t+1}\in\{j,a\}$.
+Such a time index $t$ is unique and $t<\tau$, if it exists.
+We take $\tau_{n}^{j}(a)=t$ if $t$ exists, or $\tau_{n}^{j}(a)=+\infty$ if $t$ doesn’t exist.
+Since $A_{t+1}$ is fully determined by the history up to time index $t$, for $t\in\mathbb{N}$.
+$\tau_{n}^{j}(a)$ is a stopping time regrading the filtration $\{\sigma(H(t))\}_{t\geq 1}$.
+Denote $\tilde{\mathcal{E}}^{aj}(n)=\{\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{j}(n)=\{A_{\tau_{n}^{j}(a)+1}=j,\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{a}(n)=\{A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau\}$.
+We prove two inequalities.
+$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n))\geq\Pr_{\nu}(\mathcal{E}_{j}),\forall n\leq T_{j}(a)-1.$ (8)
+$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\geq\frac{1}{4}\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)-1.$ (9)
+Given (8) and (9), from the fact that we only pull arm a at $\tau_{n}^{j}(a)+1$, we can conclude
+[…]
+[…]
+Thus, to prove (7), we suffice to show (8) (9) are both correct.
+For (8), we notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,0}^{2}}$, and that $T_{j}(a)<\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}$.
 That means for $n\leq T_{j}(a)$, we have $\mathcal{E}_{j}\Rightarrow\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\}\Rightarrow\{\tau_{n}^{j}(a)<\tau\}$, which suggests (8).
 For (9), we introduce an alternative instance $\nu_{j,a}$, with $r_{a^{\prime}}^{\nu_{j,a}}=\begin{cases}r_{a}^{\nu}&a^{\prime}=j\\ r_{j}^{\nu}&a^{\prime}=a\\ r_{a^{\prime}}^{\nu}&else\end{cases}$.
-$\nu_{j,a}$ swaps the position of arm $a$ and $j$ compared to the instance $\nu$.
-Denote $\tilde{\mathcal{E}}^{aj}(n)=\{\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{j}(n)=\{A_{\tau_{n}^{j}(a)+1}=j,\tau_{n}^{j}(a)<\tau\}$, $\tilde{\mathcal{E}}^{aj}_{a}(n)=\{A_{\tau_{n}^{j}(a)+1}=a,\tau_{n}^{j}(a)<\tau\}$.
-Notice that $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,a}^{2}}$, further $T_{j}(a)\leq\frac{O(1)}{\Delta_{j,a}^{2}(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})})}$, we can use Lemma 7 and 10, which apply the Transportation Equality(Lemma 18 in Kaufmann et al.
-(2016)), to conclude
-$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)$ (10)
+The instance distribution $\nu_{j,a}$ swaps the position of arm $a$ and $j$ compared to the instance $\nu$.
+Notice that
+$\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq\frac{1}{\Delta_{j,a}^{2}},\quad T_{j}(a)\leq\frac{O(1)}{\Delta_{j,a}^{2}(1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})})}$ (10)
+Lemma 8 and 11 combine (10) and the Transportation Equality(Lemma 18 in Kaufmann et al. (2016)), to conclude
+$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)-1$ (11)
 where $\text{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y}$.
-To complete this step, we have to put $1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}$ in the denominator of $T_{j}(a)$.
+$1+\log\frac{1}{\Pr_{\nu}(\mathcal{E}_{j})}$ in the denominator of $T_{j}(a)$ is crucial to show (11).
 (11) suggests that the algorithm is unable to distinguish $\nu$ and $\nu_{ja}$ at the round $t$ such that $N_{a}(t)+N_{j}(t)\leq T_{j}(a)$.
 For a symmetric algorithm, we have $\Pr_{\nu_{ja}}(\tilde{\mathcal{E}}^{aj}_{j}(n))=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))$, which means
-$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)$ (11)
-By the Pinsker’s Inequality $kl(x,y)\geq 2(x-y)^{2}$ and $\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j})+\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a})=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj})$, we can derive
-$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\geq\frac{1}{4}\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a).$
-The above inequality suggests that we we have proved (9), further the correctness of (C5). ∎
-3 Upper Bound on the Complexity of 1-identification
-In this section, we present our proposed Algorithm PSEEB with rough analysis.
-3.1 Algorithm
+$\text{kl}\Big(\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n)),\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\Big)\leq 0.4\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)-1$ (12)
+By the Pinsker’s Inequality $\text{kl}(x,y)\geq 2(x-y)^{2}$ and $\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{j}(n))+\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))=\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n))$, we can derive
+$\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}_{a}(n))\geq\frac{1}{4}\Pr_{\nu}(\tilde{\mathcal{E}}^{aj}(n)),\forall n\leq T_{j}(a)-1.$
+The above inequality suggests that we have proved (9), hence (C5) is justified. ∎
+3 Algorithm Design for the Upper Bound of the Sample Complexity
+In this section, we present our proposed algorithm, Parallel Sequential Exploration Exploitation on Brackets (PSEEB), together with a high-level discussion of its design and analysis.
 1Input: Threshold $\mu_{0}$, Error tolerance level $\delta$, Arm number $K$;
 2
-Randomly generate a permutation of $\sigma:[K]\rightarrow[K]$, take $B_{b}=\{\sigma(i)\}_{i=1}^{\min\{2^{b-1},K\}}$ for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$;
-// $\sigma(i)$ denotes the arm index in position $i$
+Randomly generate a permutation of $\sigma:[K]\rightarrow[K]$, take $B_{b}=\{\sigma^{-1}(i)\}_{i=1}^{\min\{2^{b-1},K\}}$ for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$;
+// $\sigma^{-1}(i)$ denotes the arm index in position $i$
 3
 4Create algorithm copies $\{\text{alg}_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ by specifying $\text{alg}_{b}$ through calling Algorithm 2 with $B\leftarrow B_{b}$, $\mu_{0}\leftarrow\mu_{0}$, $\delta\leftarrow\frac{\delta}{\lceil\log_{2}K\rceil+1}$, $C\leftarrow 1.01$, $K\leftarrow K$.
-Samples across bracket are not shared;
+Samples across brackets are not shared;
 5 for round index $t=1,2,\cdots$ do
@@
 Algorithm 1 Parallel Sequential Exploration Exploitation on Brackets(PSEEB)
-We propose Algorithm 1 to solve the 1-identification problem, which merges and adapts the idea in Katz-Samuels and Jamieson (2020) and Li and Cheung (2025).
-Katz-Samuels and Jamieson (2020) create each bracket by a unique and independent random shuffle, and a corresponding algorithm copy to pull arms in this bracket.
-In contrast to them, we specify all the brackets from a single random permutation.
-Algorithm 1 generates a random permutation of $[K]$ before the start of the pulling process.
-Based on the random permutation, Algorithm 1 create $\lceil\log_{2}K\rceil+1$ brackets $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$, satisfying $B_{1}\subset B_{2}\subset\cdots\subset B_{\lceil\log_{2}K\rceil+1}=[K]$.
+Algorithm 1 merges and adapts the bracketing idea of Katz-Samuels and Jamieson (2020) with the Sequential Exploration–Exploitation (SEE) framework of Li and Cheung (2025).
+In Katz-Samuels and Jamieson (2020), each bracket is generated by an independent random shuffle.
+In contrast, PSEEB defines all brackets from a single random permutation of $[K]$, yielding a nested family $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ satisfying $B_{1}\subset B_{2}\subset\cdots\subset B_{\lceil\log_{2}K\rceil+1}=[K]$.
 For a bracket index $1\leq b\leq\lceil\log_{2}K\rceil$, we have $|B_{b}|=2^{b}$, and $|B_{\lceil\log_{2}K\rceil+1}|=\min\{2^{\lceil\log_{2}K\rceil},K\}$.
-Our approach is to guarantee $\{B_{b}\}_{b=1}^{\lceil\log_{2}K\rceil+1}$ follow an increasing order, which is crucial to our Lemma 31.
-Lemma 31 asserts that with at least some probability, the first several brackets contain at least 1 qualified arm, if there are multiple qualified arms.
-Then, Algorithm 1 calls Algorithm 2, defined in Appendix C, as an oracle to set up $\lceil\log_{2}K\rceil+1$ algorithm copies.
-Algorithm 2 is adapted from the Algorithm 2,3,4,5 in Li and Cheung (2025), named as Sequential Exploration Exploitation(SEE).
+This monotone structure is crucial for our analysis, in particular for Lemma 7, which guarantees that with nontrivial probability the first several brackets contain at least one qualified arm when multiple qualified arms exist.
+Lemma 7.
+Given a qualified arm index $j\in[m]$, denote bracket index $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$.
+We have $\Pr(b_{j}\geq\tilde{b})\leq\exp\left(-\frac{j|B_{\tilde{b}-1}|}{K}\right),\forall\tilde{b}\in[\lceil\log_{2}K\rceil+1]$.
+Here we denote $B_{0}=\emptyset$ and take $|B_{0}|=0$.
+Then, PSEEB instantiates $\lceil\log_{2}K\rceil+1$ independent algorithm copies via Algorithm 2 (given in Appendix C).
+This procedure is adapted from Algorithms 2–5 of Li and Cheung (2025), named as Sequential Exploration Exploitation(SEE).
+We adapt SEE with modified budget allocation and stopping rules to accommodate bracketing.
 Algorithm copy $b\in[\lceil\log_{2}K\rceil+1]$ takes bracket $B_{b}$ as an input parameter and only pull arms in $B_{b}$.
-To fit in the bracketing idea, we modify its budget allocation and stopping rules.
 Detailed discussion of Algorithm 2 is in Appendix C.
-Algorithm 1 manages all the algorithm copies by a round-robin manner.
-In each round, Algorithm 1 execute one of the algorithm copies with one arm pull and check whether the copy terminates with an output $\hat{a}\in[K]\cup\{\textsf{None}\}$.
-By the design of Algorithm 2, $\text{alg}_{\lceil\log_{2}K\rceil+1}$ might terminate with taking output $\hat{a}\in[K]\cup\{\textsf{None}\}$.
-While for bracket $b$ whose size is smaller than $K$, $\text{alg}_{b}$ cannot output None but keep pulling arms, even if the $\text{alg}_{b}$ finds that $\max_{a\in B_{b}}\mu_{a}<\mu_{0}$.
-The output strategy differs with the algorithm SEE in Li and Cheung (2025).
-Once there exists an algorithm copy terminates with an output, Algorithm 1 will adopt the output as the final output and terminate the overall sampling process.
-We require each algorithm copy to pull arms in a parallel manner without sharing samples.
-Denote $\tau$ as the total length of algorithm 1, and $\tau(B_{b})$ as the total length of copy $\text{alg}_{b}$.
-We allow $\tau,\tau(B_{b})$ to take infinities.
-Since any termination of algorithm copy results in the termination of Algorithm 1, we have $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$, further $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{b})$.
-Generally speaking, for a positive instance $\nu$ with $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\mu_{K}$, we wish a bracket $B$ with size smaller than $K$ contains a qualified arm $j\in[m]$.
-Then, we wish the algorithm copy on this bracket $B$ is able to spend less pulling times to output arm $j$.
-While for a negative instance $\nu$ with $\mu_{0}>\mu_{1}\geq\mu_{2}\geq\cdots\mu_{K}$, it is possible that all the algorithm copies whose index $b\leq\lceil\log_{2}K\rceil$ never terminates.
-We only wish the copy $\lceil\log_{2}K\rceil+1$ is able to terminate with finite sampling and output $\hat{a}=\textsf{None}$ when encountering a negative instance.
-3.2 Performance Guarantee of Algorithm 1
-In the following of this section, we present the sketch proof of Theorem 2 and 3. We use Figure 1 in Appendix C.3 to summarize the overall steps.
-Sketch Proof of Theorem 2.
-The first step is prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)=0$, whose main idea is very similar to the proof of Theorem 5.2 in Li and Cheung (2025).
-Same as Li and Cheung (2025), denote $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ as the minimum phase index such that the concentration events of copy $b$ hold, we can prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)\leq\Pr(\max\{\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1},\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}\}=+\infty)=0$.
-Given the first step, we are able to conclude $\Pr\Big(\tau<+\infty\Big)=1$, by the fact that $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$.
-Then, to prove Algorithm 1 is $\delta$-PAC, we suffice to show $\Pr_{\nu}\left(\hat{a}=\textsf{None}\text{ or }\{\hat{a}\in[K],\mu_{\hat{a}}<\mu_{0}\}\right)<\delta$ for a positive instance $\nu$, and $\Pr_{\nu}\left(\hat{a}\in[K]\right)<\delta$ for a negative instance $\nu$.
-The proof for both targets is similar.
-Recall we deliver $\frac{\delta}{\lceil\log_{2}K\rceil+1}$ to each algorithm copy in the Line 1 o Algorithm 1.
-We are able to prove $\Pr(\text{alg}_{b}\text{ outputs incorrect }\hat{a})<\frac{O(1)\delta}{\lceil\log_{2}K\rceil+1}$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$.
-Taking the union bound, we complete the proof.
-∎
-Sketch Proof of Theorem 3.
-We split the proof into two cases, which are positive and negative $\nu$.
-If the instance $\nu$ is negative, we focus on analyzing $\tau(B_{\lceil\log_{2}K\rceil+1})$, which is very similar to the negative case of Theorem 5.3 in Li and Cheung (2025).
-By the fact that $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{\lceil\log_{2}K\rceil+1})$, we complete the proof.
-If the instance $\nu$ is positive, denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$.
-Denote $\tilde{T}_{j}(a)=\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, $\tilde{T}_{j}(a;k)=\frac{\log\frac{4K}{\delta_{k}}+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, we are going to show $\mathbb{E}\tau\leq O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\tilde{T}_{j}(a)\right)$ holds for all $j\in[m]$.
-What’s more, the constant hidden by the Big O notation is independent of $j$.
-For any given $j\in[m]$, we introduce $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$, which is the minimum bracket index that contains at least one arm not worse than $\mu_{j}$.
-As we adopt a random permutation to create brackets, $b_{j}$ is a random variable, and we have $\mathbb{E}\tau=\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\mathbb{E}\tau\mathds{1}(b_{j}=b)$.
-Then, we turn to analyze $O(\log K)\mathbb{E}\tau(B_{b})\mathds{1}(b_{j}=b)$, which is the upper bound of $\mathbb{E}\tau\mathds{1}(b_{j}=b)$.
-Denote $\tau^{\text{ee}}(B_{b}),\tau^{\text{et}}(B_{b})$ as the total length of exploration and exploitation periods of algorithm copy $\text{alg}_{b}$ separately.
-For any arm $a\in B_{b}$, denote $N_{a}^{\text{ee},B_{b}}$ as the pulling times of arm $a$ in the exploration period.
-Not hard to see $\tau(B_{b})=\tau^{\text{ee}}(B_{b})+\tau^{\text{et}}(B_{b})$.
-We are going to prove
-$\mathbb{E}\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)\Pr(b_{j}=b),$ (12)
-$\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)\leq\sum_{a=1}^{K}O(1)\tilde{T}_{j}(a)\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}+O(\log K)|B_{b}|\Pr(b_{j}=b).$ (13)
-If (12)(13) are both correct, we can sum up (12)(13) for $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$.
-Lemma 31, 33, 34 prove upper bounds for $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}|B_{b}|\Pr(b_{j}=b)$ and $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}$, which complete the proof of Theorem 3.
-To prove (12), we first show the algorithm must terminate no later than the end of phase $L_{b}^{\text{pos}}=\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{O(1)(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$, if $b_{j}=b$.
-Notice that $\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}=O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$, and Lemma 17 shows $\Pr(\kappa^{\text{ee}}_{b}\geq k)\leq\frac{\pi^{2}\delta_{k-1}}{6}$, $\Pr(\kappa^{\text{et}}_{b}\geq k)\leq\frac{\pi^{2}\delta}{6\alpha_{k-1}}$, we can take expectation on both sides of $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}\mathds{1}(b_{j}=b)$ to complete the proof of (12).
-To prove (13), we first show for each $a\in B_{b}$, we have
-$N_{a}^{\text{ee},B_{b}}\mathds{1}(b_{j}=b)\leq O(\tilde{T}_{j}(a;L_{b}^{\text{pos}}))(C+1)^{2}\mathds{1}(b_{j}=b).$ (14)
-The intuition is as follows.
-Given the fact that $j\in B_{b}$, once the pulling times of arm $a$ is higher than $O(\tilde{T}_{j}(a;L_{b}^{\text{pos}}))(C+1)^{2}$, copy $\text{alg}_{b}$ can assert either $\text{UCB}_{a}<\mu_{j}$ or $\text{LCB}_{a}>\mu_{0}$, with a constant confidence level.
-In both cases, $\text{alg}_{b}$ stops pulling arm $a$ in the exploration period.
-By the facts that $\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)=(|Q|+\sum_{a\in B_{b}}N_{a}^{\text{ee},B_{b}})\mathds{1}(b_{j}=b)$ and elements in $B_{b_{j}}$ are independent of $\kappa^{\text{ee}}_{b_{j}},\kappa^{\text{et}}_{b_{j}}$, we can sum up (14) for $a\in B_{b}$ and derive
-$\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)\leq\sum_{a=1}^{K}O(1)\tilde{T}_{j}(a)\Pr(a\in B_{b},b_{j}=b)+O(\log K)|B_{b}|\Pr(b_{j}=b).$
-Since $\Pr(a\in B_{b},b_{j}=b)\leq\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}$, and $\Pr(a\in B_{b})=\frac{|B_{b}|}{K}$, we can conclude (13). ∎
+Algorithm 1 manages all the algorithm copies in a round-robin manner.
+In each global round (Line 1), Algorithm 1 execute one of the algorithm copies with one arm pull, and the overall algorithm terminates as soon as some copy outputs $\hat{a}\in[K]\cup\{\textsf{None}\}$.
+By construction, only the copy $\textnormal{alg}_{\lceil\log_{2}K\rceil+1}$ (which operates on $B_{\lceil\log_{2}K\rceil+1}=[K]$) is permitted to output None.
+For any other bracket $B_{b}$ with $|B_{b}|<K$, the corresponding copy never outputs None; even if it becomes convinced that $\max_{a\in B_{b}}\mu_{a}<\mu_{0}$, it has to continue sampling.
+This differs from SEE Li and Cheung (2025) and is necessary for the analysis of $\mathbb{E}\tau$ under bracketing.
+Let $\tau$ denote the stopping time of PSEEB, and let $\tau(B_{b})$ denote the number of rounds executed by copy $\textnormal{alg}_{b}$ up to its own termination (possibly $\infty$).
+Since the Algorithm 1 terminates upon the first terminating copy and adopts the output, we have $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$, and hence $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{b})$.
+For a positive instance $\nu$, the intuition of sample complexity $O\Big(\log K\min_{1\leq j\leq m}\big\{\log(1/\delta)/\Delta_{j,0}^{2}+H(j)\log^{4}K\log 1/\Delta_{j,0}^{2}\big\}\Big)$ in Theorem 3 is as follows.
+Consider a “lucky” agent with prior knowledge $j^{*}=\arg\min_{1\leq j\leq m}\{(\log(1/\delta))/\Delta_{j,0}^{2}+H(j)\}$.
+This agent samples $K/j^{*}$ arms in $[K]$ uniformly at random, successfully capturing an arm $j^{*}$ with reward $\mu_{j^{*}}$.
+However, the arm’s unknown position requires the agent $H(j^{*})$ pulls to identify $j^{*}$ at a low confidence level.
+Subsequently, it must pull arm $j^{*}$ another $O(\log(1/\delta)/\Delta_{j^{*},0}^{2})$ times to confirm $j^{*}$ is qualified with confidence $1-\delta$.
+Algorithm 1 maintains multiple parallel copies, wishing $j^{*}$ is captured in an early bracket.
+Since the agent cannot pre-identify the correct bracket and low-confidence identification may be erroneous, the expected sample complexity includes an additional factor of $\log^{4}K\log\frac{1}{\Delta_{j,0}^{2}}$ for the rough identification phase.
+Theorems 2 and 3 are derived from several technical lemmas, leveraging Lemma 7. A proof sketch is provided in Appendix C.3, with the full proof in Appendix C.4.
+4 Conclusion
+This paper works on the open problem 1-identification, providing a novel lower bound for positive instances derived by a specialized optimization formulation.
+We further design an algorithm that is nearly optimal across all instances.
+Further refining logarithmic factors remains a future research.
 Appendix A Detailed Literature Review
@@
 We simplify the result in the history literature with smaller but explicit formulations, for the comparison in Table 1.
-A.1.1 Theorem 2 in Poiani et al. (2025)
+A.1.1 Theorem 2 in Poiani et al. [2025]
 Theorem 2 in Poiani et al.
-(2025) presents the Non-asymptotic analysis for algorithm S-TaS.
+[2025] presents the Non-asymptotic analysis for algorithm S-TaS.
 For an instance $\nu$ with $r_{a}^{\nu}=\mu_{a}$, $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$ and , they define $\neg a=\{\nu^{\prime}:r_{a}^{\nu^{\prime}}<\mu_{0}\}$ for $a\in[m]$, and
 $i_{F}(\nu)=\arg\max_{a\in[K]}\sup_{w\in\Delta_{K}}\inf_{\nu^{\prime}\in\neg a}\sum_{a^{\prime}\in K}\frac{w_{a^{\prime}}(r_{a^{\prime}}^{\nu^{\prime}}-r_{a^{\prime}}^{\nu})^{2}}{2}.$
-Then, Theorem 2 in Poiani et al. (2025) considers $\epsilon_{\nu}$ such that for all instance $\nu^{\prime}$ such that
+Then, Theorem 2 in Poiani et al. [2025] considers $\epsilon_{\nu}$ such that for all instance $\nu^{\prime}$ such that
 $\|\{r_{a}^{\nu^{\prime}}\}_{a=1}^{K}-\{r_{a}^{\nu}\}_{a=1}^{K}\|_{\infty}\leq\epsilon_{\nu}\implies i_{F}(\nu^{\prime})\subset i_{F}(\nu)\cup([K]\setminus i^{*}(\nu)).$
@@
 Thus, we can conclude $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\epsilon_{\nu}^{4}}$ is smaller than the proposed upper bound in Poiani et al.
-(2025).
+[2025].
 The remaining work is to show $\epsilon_{\nu}\leq\Delta_{0,1}$.
@@
 Thus, we have shown $\frac{2\log(1/\delta)}{\Delta_{0,1}^{2}}+\frac{K^{2}}{\Delta_{0,1}^{4}}$ is a lower bound of $10K^{4}+T_{0}(\delta)+\frac{\pi^{2}}{24}$.
-A.1.2 Theorem 3 in Kano et al. (2017)
-Following Appendix A.2 in Li and Cheung (2025), we can also simplify the Theorem 3 in Kano et al. (2017). By applying the algorithm HDoC, for $\nu\in\mathcal{S}^{\text{pos}}$, we have
+A.1.2 Theorem 3 in Kano et al. [2017]
+Following Appendix A.2 in Li and Cheung [2025], we can also simplify the Theorem 3 in Kano et al. [2017]. By applying the algorithm HDoC, for $\nu\in\mathcal{S}^{\text{pos}}$, we have
 $\mathbb{E}_{\nu}\tau\leq n_{1}+\sum_{a=2}^{K}\left(\frac{\log(K\max_{j\in[K]}n_{j})}{2(\Delta_{1,a}-2\epsilon)^{2}}+\delta n_{a}\right)+\frac{K^{2-\frac{\epsilon^{2}}{(\min_{a\in[K]}\Delta_{0,a}-\epsilon)^{2}}}}{2\epsilon^{2}}+\frac{K(5+\log\frac{1}{2\epsilon^{2}})}{4\epsilon^{2}},$
@@
 $\geq$ $\frac{\log\frac{K}{\delta}}{2\Delta_{0,1}^{2}}+\frac{H_{1}}{2}\log\log\frac{1}{\delta}+\frac{K}{4\epsilon^{2}}.$
-A.1.3 Thereom 2 in Jourdan and Réda (2023)
-Jourdan and Réda (2023) do not explicitly provide upper bound for $\mathbb{E}\tau$. Therorem 2 in Jourdan and Réda (2023) suggests that for a positive instance $\nu$.
+A.1.3 Thereom 2 in Jourdan and Réda [2023]
+Jourdan and Réda [2023] do not explicitly provide upper bound for $\mathbb{E}\tau$. Therorem 2 in Jourdan and Réda [2023] suggests that for a positive instance $\nu$.
 $\mathbb{E}_{\nu}\tau$ $\leq C^{\text{pos}}(\delta)+\frac{K\pi^{2}}{6}+1$
@@
 $\bar{W}_{-1}(x)=$ $-W_{-1}(-e^{-x})\approx x+\log x.$
-$W_{-1}$ is the negative branch of the Lambert W function, and $D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})$ is further defined in the Appendix F in Jourdan and Réda (2023), as
+$W_{-1}$ is the negative branch of the Lambert W function, and $D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})$ is further defined in the Appendix F in Jourdan and Réda [2023], as
 $D^{\text{pos}}(\{\mu_{a}\}_{a=1}^{K})=$ $S_{\mu}+1-\frac{6\log S_{\mu}}{\Delta_{0,1}^{2}}$
@@
 Thus, we can conclude $\left(\sum_{a:\mu_{a}\geq\mu_{0}}\frac{2}{\Delta_{0,a}^{2}}\right)\log\frac{K}{\delta}$ is a lower bound of $C^{\text{pos}}(\delta)+\frac{K\pi^{2}}{6}+1$.
-A.1.4 Theorem 8 in Katz-Samuels and Jamieson (2020)
-As discussed in the Appendix A.1 in Li and Cheung (2025), proposed Algorithm FWER-FWPD only work on the positive instance with a different definition of $\mathbb{E}\tau$.
+A.1.4 Theorem 8 in Katz-Samuels and Jamieson [2020]
+As discussed in the Appendix A.1 in Li and Cheung [2025], proposed Algorithm FWER-FWPD only work on the positive instance with a different definition of $\mathbb{E}\tau$.
 But we can adapt their result.
-Apply Algorithm FWER-FWPD on a positive instance $\nu$, Theorem 8 in Katz-Samuels and Jamieson (2020) suggests that
+Apply Algorithm FWER-FWPD on a positive instance $\nu$, Theorem 8 in Katz-Samuels and Jamieson [2020] suggests that
 $\mathbb{E}\tau\leq O\left(\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)\log\left((\frac{K}{m}-1)\cdot\frac{\log\frac{1}{\delta}}{\Delta^{2}}+\log\frac{\frac{K}{m}}{\delta}\right)\right),$
@@
 B.1 Proof of Preparation
-The first step is to prove Lemma 4, showing that we suffice to only consider symmetric algorithm
+The first step is to prove Lemma 4, showing that we suffice restrict attention to symmetric algorithms.
 Proof of Lemma 4.
+Let $S$ be the set of all possible permutations of $[K]$.
 Given the alg, we define $\widetilde{\text{alg}}$ as the following.
@@
 For any pulling sequence $\{a_{t}\}_{t=1}^{T}\in[K]^{T}$ and any permutation $\sigma$, we have
-$\Pr_{\widetilde{\text{alg}},\sigma(\nu^{\prime})}(\forall t\in[T],A_{t}=a,\tau>t)=\frac{\sum_{\tilde{\sigma}\in S}\Pr_{\text{alg},\tilde{\sigma}(\nu^{\prime})}(\forall t\in[T],A_{t}=a,\tau>t)}{K!},$
-which means $\tilde{\text{alg}}$ fits in the definition. Then, we have
+$\Pr_{\widetilde{\text{alg}},\sigma(\nu^{\prime})}(\forall t\in[T],A_{t}=\sigma(a),\tau>t)=\frac{\sum_{\tilde{\sigma}\in S}\Pr_{\text{alg},\tilde{\sigma}(\nu^{\prime})}(\forall t\in[T],A_{t}=\tilde{\sigma}(a),\tau>t)}{K!},$
+which means $\widetilde{\textnormal{alg}}$ fits in the definition. Then, we have
 $\max_{\sigma\in S}\mathbb{E}_{\text{alg},\sigma(\nu^{\prime})}\tau$
@@
 We can conclude $\widetilde{\text{alg}}$ is also $\delta$-PAC. ∎
-The next step, which is also the main trick, is to the prove Lemma 7.
+The next step establishes a technical change-of-measure bound that controls a conditional KL divergence.
+This lemma is the main device used to argue that a symmetric algorithm cannot reliably distinguish two arms when the total number of samples from the pair is small.
 Lemma 8.
-Given a pair $i,j\in[K],i\neq j$ and an arbitrary instance $\nu^{\prime}$, denote $\Delta=|r^{\nu^{\prime}}_{i}-r^{\nu^{\prime}}_{j}|$. Define
+Fix a pair $i,j\in[K],i\neq j$ and an arbitrary instance $\nu^{\prime}$, denote $\Delta=|r^{\nu^{\prime}}_{i}-r^{\nu^{\prime}}_{j}|$. Define
 $\tau_{n}=\min\{t:N_{i}(t)+N_{j}(t)=n,A_{t+1}\in\{i,j\}\}.$
@@
 $\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\frac{1}{2}n\Delta^{2}\Pr_{\nu^{\prime}}(\tau_{n}<\tau).$
-An observation is given the realized reward and a symmetric algorithm, the set $\{t:N_{i}(t)+N_{j}(t)=n,A_{t+1}\in\{i,j\}\}=:S_{ij}(n)$ contains at most one element.
-If $|S_{ij}(n)|=1$, $\tau_{n}$ will be the unique element, and we have $\tau_{n}<\tau$ holds in this case.
-If $|S_{ij}(n)|=0$, we have $\tau_{n}=\infty$, meaning that $\tau_{n}\geq\tau$.
-Another observation is
+An observation is given a fixed sample path and a symmetric algorithm, the set $\{t:N_{i}(t)+N_{j}(t)=n,A_{t+1}\in\{i,j\}\}=:S_{ij}(n)$ contains at most one element.
+If $|S_{ij}(n)|=1$, $\tau_{n}$ will be the unique element, and $\tau_{n}<\tau$ holds; if $|S_{ij}(n)|=0$, then $\tau_{n}=\infty$ and hence $\tau_{n}\geq\tau$.
+Moreover,
 $\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}+\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}$
-is the kl-divergence between two Bernoulli distribution with probability $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)$, $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)$, and not hard to see $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)+\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)=1$.
+is the kl-divergence between two Bernoulli distribution with probability $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)$, $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)$.
+From the definition of $\tau_{n}$, not hard to see $\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=i|\tau_{n}<\tau)+\Pr_{\nu^{\prime}}(A_{\tau_{n}+1}=j|\tau_{n}<\tau)=1$.
 Proof of Lemma 8.
 Denote $\nu^{\prime\prime}$ as the instance which switch the position of arm $i$ and arm $j$ in $\nu^{\prime}$, i.e.
-$r_{i}^{\nu^{\prime}}=r_{j}^{\nu^{\prime\prime}},r_{j}^{\nu^{\prime}}=r_{i}^{\nu^{\prime\prime}}$ and $r_{a}^{\nu^{\prime}}=r_{a}^{\nu^{\prime\prime}},a\neq i,j$.
+$r_{a}^{\nu^{\prime\prime}}=\begin{cases}r_{j}^{\nu^{\prime}}&a=i\\ r_{i}^{\nu^{\prime}}&a=j\\ r_{a}^{\nu^{\prime}}&\text{else}\end{cases}$.
 We first derive
-$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$ (15)
-$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$ (16)
+$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right)$ (13)
+$\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}\leq\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right)$ (14)
 where $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta^{2}}{2},\Delta^{2})$.
-Since these two inequalities are symmetric, we suffice to prove the first one.
-Take $\mathcal{E}=\{\tau_{n}<\tau,A_{\tau_{n}+1}=i\}$, from the Lemma 18 in Kaufmann et al.
-(2016), we have
-$\Pr_{\nu^{\prime\prime}}(\mathcal{E})=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)$
-$=$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})+$
-$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\neg\mathcal{E}\right]\Pr_{\nu^{\prime}}(\neg\mathcal{E})$
-$=$ $\mathbb{E}_{\nu^{\prime}}\left[\exp\left(-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})$
-$\geq$ $\exp\left(\mathbb{E}_{\nu^{\prime}}\left[-\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}|\mathcal{E}\right]\right)\Pr_{\nu^{\prime}}(\mathcal{E}),$
+Since these two inequalities are symmetric, we suffice to prove the first (13).
+Take $\mathcal{E}=\{\tau_{n}<\tau,A_{\tau_{n}+1}=i\}$.
+Notice that $\mathcal{E}$ is $\tau_{n}$-measurable, as both $\{\tau_{n}<\tau\},\{A_{\tau_{n}+1}=i\}$ are $\tau_{n}$-measurable.
+From the Lemma 18 in Kaufmann et al.
+[2016],
+$\Pr_{\nu^{\prime\prime}}(\mathcal{E})=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)$
+$=$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})+$
+$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\exp\left(-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)|\neg\mathcal{E}\right]\Pr_{\nu^{\prime}}(\neg\mathcal{E})$
+$=$ $\mathbb{E}_{\nu^{\prime}}\left[\exp\left(-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)|\mathcal{E}\right]\Pr_{\nu^{\prime}}(\mathcal{E})$
+$\geq$ $\exp\left(\mathbb{E}_{\nu^{\prime}}\left[-\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}-\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}|\mathcal{E}\right]\right)\Pr_{\nu^{\prime}}(\mathcal{E}),$
 where $\hat{KL}_{i,s}:=\log\frac{\exp(-\frac{(X_{i,s}-r_{i}^{\nu^{\prime}})^{2}}{2})}{\exp(-\frac{(X_{i,s}-r_{i}^{\nu^{\prime\prime}})^{2}}{2})}=\frac{(2X_{i,s}-r_{i}^{\nu^{\prime}}-r_{j}^{\nu^{\prime}})(r_{i}^{\nu^{\prime}}-r_{j}^{\nu^{\prime}})}{2}$, $X_{i,s}\sim N(r_{i}^{\nu^{\prime}},1)$, $\hat{KL}_{j,s}:=\log\frac{\exp(-\frac{(X_{j,s}-r_{j}^{\nu^{\prime}})^{2}}{2})}{\exp(-\frac{(X_{j,s}-r_{j}^{\nu^{\prime\prime}})^{2}}{2})}=\frac{(2X_{j,s}-r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})}{2}$, $X_{j,s}\sim N(r_{j}^{\nu^{\prime}},1)$.
@@
 Following the last line, we have
-$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}|\mathcal{E}\right]$
-$=$ $\frac{\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)\right]}{\Pr_{\nu^{\prime}}(\mathcal{E})},$
-and we can conclude $\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\geq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau)}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau)}\hat{KL}_{j,s}\right)\right]$.
+$\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}|\mathcal{E}\right]$
+$=$ $\frac{\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)\right]}{\Pr_{\nu^{\prime}}(\mathcal{E})},$
+and we can conclude $\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\geq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\left(\sum_{s=1}^{N_{i}(\tau_{n})}\hat{KL}_{i,s}+\sum_{s=1}^{N_{j}(\tau_{n})}\hat{KL}_{j,s}\right)\right]$.
 From the above calculation, we know $\hat{KL}_{i,s},\hat{KL}_{j,s}$ both follow $N(\frac{(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}})^{2}}{2},(r_{j}^{\nu^{\prime}}-r_{i}^{\nu^{\prime}}))$, and we can rewrite the conclusion as
-$\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right]$
+$\Pr_{\nu^{\prime}}(\mathcal{E})\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right]$
 where $Z_{s}|Z_{1},Z_{2},\cdots,Z_{s-1}\sim N(\frac{\Delta^{2}}{2},\Delta^{2})$. The above inequality suggests that we have proved (13) and (14).
-By the fact that the algorithm is symmetric, we can conclude $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)$ and $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)$.
+By symmetry of the algorithm, we can conclude $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)$ and $\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)=\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=i)$.
 Sum up inequality (13) and (14), we get
@@
 $\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)\log\frac{\Pr_{\nu^{\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}{\Pr_{\nu^{\prime\prime}}(\tau_{n}<\tau,A_{\tau_{n}+1}=j)}$
-$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right).$
+$\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right).$
 The remaining work is to prove
-$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$
+$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right)$
 $\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\frac{1}{2}n\Delta^{2}\Pr_{\nu^{\prime}}(\tau_{n}<\tau).$
-Denote $D:=\max_{1\leq t\leq n}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}}{2}\right)$, we have
-$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau)+N_{j}(\tau)}Z_{s}\right)$
+Let $D:=\max_{1\leq t\leq n}\sum_{s=1}^{t}\left(Z_{s}-\frac{\Delta^{2}}{2}\right)$. Since $\{\tau_{n}<\tau\}$ implies $N_{i}(\tau_{n})+N_{j}(\tau_{n})\leq n$, we have
+$\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\sum_{s=1}^{N_{i}(\tau_{n})+N_{j}(\tau_{n})}Z_{s}\right)$
 $\leq$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)\left(\max_{1\leq t\leq n}\sum_{s=1}^{n}Z_{s}\right)$
@@
 The remaining work is to prove $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\tau_{n}<\tau)D\leq\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\tau_{n}<\tau)$.
-To simplify the notation, we take event $\mathcal{E}=\{\tau_{n}<\tau\}$ here and we can conduct the following calculation.
-$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D\right]$
-$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D\right]$
-$=$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathds{1}(\mathcal{E})\mathds{1}(D\geq 0)D>s\Big)ds$
-$\leq$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathcal{E},D>s\Big)ds$
-$\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\Pr_{\nu^{\prime}}\Big(D>s\Big)\right\}ds$
-From Lemma 32, we can conclude $\Pr_{\nu^{\prime}}\big(D\geq s\big)\leq\exp(-\frac{s^{2}}{2n\Delta^{2}})$. Plug in the inequality, we get
-$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E})D\right]\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}),\exp(-\frac{s^{2}}{2n\Delta^{2}})\right\}ds$
-$=$ $\int_{0}^{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}\Pr_{\nu^{\prime}}(\mathcal{E})ds+\int_{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2n\Delta^{2}})ds$
-$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\frac{\sqrt{2\pi}\sqrt{n\Delta^{2}}}{\sqrt{2\pi}}\int_{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}^{+\infty}\exp(-\frac{s^{2}}{2})ds$
-$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n\Delta^{2}}\exp(-\frac{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}}{2})$
-$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}\Pr_{\nu^{\prime}}(\mathcal{E})+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\mathcal{E}).$
+To simplify the notation, we take event $\mathcal{E}_{n}=\{\tau_{n}<\tau\}$ here and use the layer-cake representation,
+$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E}_{n})D\right]$
+$\leq$ $\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E}_{n})\mathds{1}(D\geq 0)D\right]$
+$=$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathds{1}(\mathcal{E}_{n})\mathds{1}(D\geq 0)D>s\Big)ds$
+$\leq$ $\int_{0}^{+\infty}\Pr_{\nu^{\prime}}\Big(\mathcal{E}_{n},D>s\Big)ds$
+$\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}_{n}),\Pr_{\nu^{\prime}}\Big(D>s\Big)\right\}ds$
+From Lemma 32, we can conclude $\Pr_{\nu^{\prime}}\big(D\geq s\big)\leq\exp(-\frac{s^{2}}{2n\Delta^{2}})$. Plugging this tail bound yields
+$\mathbb{E}_{\nu^{\prime}}\left[\mathds{1}(\mathcal{E}_{n})D\right]\leq$ $\int_{0}^{+\infty}\min\left\{\Pr_{\nu^{\prime}}(\mathcal{E}_{n}),\exp(-\frac{s^{2}}{2n\Delta^{2}})\right\}ds$
+$=$ $\int_{0}^{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n})ds+\int_{\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}}^{+\infty}\exp(-\frac{s^{2}}{2n\Delta^{2}})ds$
+$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n})+\frac{\sqrt{2\pi}\sqrt{n\Delta^{2}}}{\sqrt{2\pi}}\int_{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}}^{+\infty}\exp(-\frac{s^{2}}{2})ds$
+$\leq$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n})+\sqrt{2\pi n\Delta^{2}}\exp(-\frac{\sqrt{2\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}}{2})$
+$=$ $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E}_{n})}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n})+\sqrt{2\pi n\Delta^{2}}\Pr_{\nu^{\prime}}(\mathcal{E}_{n}).$
 The second last step is by the equality $\int_{t}^{+\infty}\frac{1}{\sqrt{2\pi}}e^{-\frac{s^{2}}{2}}ds\leq e^{-\frac{t^{2}}{2}}$ for $t>0$. The reason is
@@
 $\log\frac{\Pr_{\nu^{\prime}}(\mathcal{E})}{\Pr_{\nu^{\prime\prime}}(\mathcal{E})}\leq\sum_{i=1}^{M}\sqrt{2n_{i}\Delta_{i}^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\mathcal{E})}}+\sqrt{2\pi n_{i}\Delta_{i}^{2}}+\frac{1}{2}n_{i}\Delta_{i}^{2}$
-Proof.
-From the Lemma 18 in Kaufmann et al. (2016), we have
+Proof of Lemma 9.
+From the Lemma 18 in Kaufmann et al. [2016], we have
 $\Pr_{\nu^{\prime\prime}}(\mathcal{E})=$ $\mathbb{E}_{\nu^{\prime}}\mathds{1}(\mathcal{E})\exp\left(-\sum_{i=1}^{M}\sum_{s=1}^{N_{i}(\tilde{\tau})}\hat{KL}_{i,s}\right)$
@@
 B.2 Proof of Theorem 5
-Following the sketch proof in section 2, we prove the lemmas one by one. The first step is to build a “constant probability lower bound”.
+Following the sketch proof in section 2, we prove the following lemmas one by one.
+The first step is to build a “probabilistic constant lower bound”.
+Without further description, we only consider symmetric algorithms in this section.
 Lemma 10.
@@
 Proof of Lemma 10.
-Recall the notation $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$, and the $\delta$-PAC assumption requires
+Recall the notation $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$. The $\delta$-PAC assumption implies
 $\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m]\cup\{a:\mu_{a}=\mu_{0}\})>1-\delta$
+Before the proof of the Lemma 10, we first show a claim
 […]
-We prove the claim by contradiction, if $Pr_{\nu}(\tau<+\infty,\hat{a}\in\{a:\mu_{a}=\mu_{0}\})\geq 4\delta$ holds, we can find a $T_{0}$, such that
+This claim is to show $\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m])\geq 1-\Theta(\delta)$.
+If $\{a:\mu_{a}=\mu_{0}\}=\emptyset$, (15) is immediate.
+We therefore assume $\{a:\mu_{a}=\mu_{0}\}\neq\emptyset$ and argue by contradiction.
+If $Pr_{\nu}(\tau<+\infty,\hat{a}\in\{a:\mu_{a}=\mu_{0}\})\geq 4\delta$ holds, then there exists a (possibly $\delta$-dependent) constant $T_{0}(\delta)$, such that
 $\Pr_{\nu}(\tau<T_{0},\hat{a}\in\{a:\mu_{a}=\mu_{0}\})\geq 2\delta.$ (16)
 We define $\nu^{\prime}(\epsilon)$ with $r_{a}^{\nu^{\prime}(\epsilon)}=\begin{cases}\mu_{0}-\epsilon&\mu_{a}=\mu_{0}\\ r_{a}^{\nu}&\text{else}\end{cases}$, where $\epsilon$ is an arbitrary positive number.
-We know $i^{*}(\nu^{\prime}(\epsilon))=[m]$ and we have
+We know $i^{*}(\nu^{\prime}(\epsilon))=[m]$.
+By $\delta$-PAC assumption,
 $\Pr_{\nu^{\prime}(\epsilon)}(\tau<+\infty,\hat{a}\in[m])>1-\delta$
@@
 […]
-[…]
+With the claim (15) established, we deduce
 $\Pr_{\nu}(\tau<+\infty,\hat{a}\in[m]\cup\{a:\mu_{a}=\mu_{0}\})>1-\delta$
@@
 which leads to a contradiction, as $\delta<10^{-8}$. ∎
-The next step is to prove a lemma analyzing the pulling behavior.
+We next quantify the algorithm’s inability to distinguish two arms when the total arm pulls of the pair is small.
 Lemma 11.
@@
 $<$ $0.4\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$
+Dividing $\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$ on both sides, the conclusion is equivalent to
+$\text{kl}\Big(\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a|\tau_{n}^{j}(a)<\tau),1-\Pr_{\nu}(A_{\tau_{n}^{j}(a)+1}=a|\tau_{n}^{j}(a)<\tau)\Big)\leq 0.4$
+where $\text{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y}$.
+The reason is $\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=j)+\Pr_{\nu}(\tau_{n}^{j}(a)<\tau,A_{\tau_{n}^{j}(a)+1}=a)=\Pr_{\nu}(\tau_{n}^{j}(a)<\tau)$.
 Proof of Lemma 11.
-The main idea is to utilize Lemma 7. From the definition, we know $T_{j}(a)\geq 1$. For $n=1,2,\cdots,T_{j}(a)-1$, we have
+The main idea is to utilize Lemma 8. Take $\Delta=\Delta_{a,j}$. From the definition, we know $T_{j}(a)\geq 1$. For $n=1,2,\cdots,T_{j}(a)-1$, we have
 $\sqrt{2n\Delta^{2}\log\frac{1}{\Pr_{\nu^{\prime}}(\tau_{n}<\tau)}}$
@@
 $\leq$ $\sqrt{\frac{1}{50}}.$
-The third step is by the fact that $\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\}\subset\{N_{a}(\tau)+N_{j}(\tau)>n\}$ holds for all $n=0,1,\cdots,T_{j}(a)-1$.
-The last step is by the fact that $\frac{\Delta_{a,j}^{2}}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq 1$.
-For the same reason, we have
+The third line is by the fact that $\{N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\}\subset\{N_{a}(\tau)+N_{j}(\tau)>n\}$ holds for all $n=0,1,\cdots,T_{j}(a)-1$.
+The last step is by the fact that $\frac{\Delta_{a,j}^{2}}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\leq 1$ and
+$\frac{\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}}{1+\log\frac{1}{\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)}}\leq 1.$
+Similarly, we have
 $\sqrt{2\pi n\Delta^{2}}\leq$ $\sqrt{\frac{\pi}{50}}$
@@
 Not hard to validate that $\sqrt{\frac{1}{50}}+\sqrt{\frac{\pi}{50}}+\frac{1}{200}<0.4$. Applying Lemma 8, we prove the result. ∎
-Based on the Lemma 10, we can prove a lower bound connecting $\mathbb{E}N_{A}(\tau)$ and $\Pr_{\nu}(N_{j}(\tau)\geq\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j)$, which is the following Lemma.
+The next lemma converts Lemma 11 into a conclusion relating $\mathbb{E}_{\nu}N_{a}(\tau)$ to the probability of correctly recommending a given qualified arm with sufficiently many samples.
 Lemma 12.
@@
 Proof of Lemma 12.
-We firstly prove a claim: Assume $\tau_{n+1}^{j}(a)<\tau$, for all $t$ such that $\tau_{n}+1^{j}(a)<t<\tau_{n+1}^{j}(a)+1$, we have $A_{t}\neq a$.
+We first claim that if $\tau_{n+1}^{j}(a)<\tau$, for all $t$ such that $\tau_{n+1}^{j}(a)<t<\tau_{n+1}^{j}(a)+1$, we have $A_{t}\neq a$.
 We prove this claim by contradiction.
@@
 $=$ $n+2.$
-However, $t\leq\tau_{n+1}^{j}(a)$ suggests that $N_{j}(t)+N_{a}(t)\leq N_{1}(\tau_{n+1}^{j}(a))+N_{a}(\tau_{n+1}^{j}(a))=n+1$, which leads to a contradiction.
-Then we can rewrite $\mathbb{E}_{\nu}N_{a}(\tau)$ as the following.
+However, $t\leq\tau_{n+1}^{j}(a)$ suggests that $N_{j}(t)+N_{a}(t)\leq N_{1}(\tau_{n+1}^{j}(a))+N_{a}(\tau_{n+1}^{j}(a))=n+1$, which leads to a contradiction. The claim follows.
+Using the claim, we can rewrite
 $\mathbb{E}_{\nu}N_{a}(\tau)$
@@
 $=$ $\mathbb{E}\sum_{n=0}^{+\infty}\mathds{1}(A_{\tau_{n}^{j}(a)+1}=a)\mathds{1}(\tau>\tau_{n}^{j}(a))$
-$\geq$ $\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a,\tau>\tau_{n}^{j}(a)\Big).$
+$\geq$ $\sum_{n=0}^{T_{j}(a)-1}\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a,\tau>\tau_{n}^{j}(a)\Big).$ (17)
 From Lemma 11, we know
-$kl\Bigg(\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big),1-\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)\Bigg)<0.4$
-where $kl(x,y)=x\log\frac{x}{y}+(1-y)\log\frac{1-y}{1-x}$. From Pinsker’s Inequality, we can further conclude
+$\text{kl}\Bigg(\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big),1-\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)\Bigg)<0.4$
+where $\text{kl}(x,y)=x\log\frac{x}{y}+(1-y)\log\frac{1-y}{1-x}$. From Pinsker’s Inequality, we can further conclude
 $2\Bigg(1-2\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a|\tau>\tau_{n}^{j}(a)\Big)\Bigg)^{2}<0.4$
@@
 $\Rightarrow$ $\Pr_{\nu}\Big(A_{\tau_{n}^{j}(a)+1}=a,\tau>\tau_{n}^{j}(a)\Big)>\frac{1}{4}\Pr_{\nu}\Big(\tau>\tau_{n}^{j}(a)\Big).$
-The last line suggests that
+Combine the last line and (17), we get
 $\mathbb{E}_{\nu}N_{a}(\tau)$
@@
 which is the conclusion we wish to prove. ∎
-Given Lemma 11, we are able to take the optimal value of an optimization problem as the lower bound of $\mathbb{E}_{\nu}\tau$, which is Lemma 6.
+Finally, we use Lemma 12 to lower bound $\mathbb{E}_{\nu}\tau$ by the optimal value of an optimization problem, as stated in Lemma 6.
 Proof of Lemma 6.
@@
 $\geq$ $\sum_{a=1}^{K}\frac{1}{4}T_{j}(a)\Pr_{\nu}\Big(N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}},\hat{a}=j\Big).$
-Thus, we can conclude
+The third step is by the Lemma 12. Since the last line holds for any $j\in[m]$,
 $\mathbb{E}\tau$
@@
 $\geq$ $\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Pr_{\nu}(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}})$ (19)
-Merge (18) and (19), notice that Lemma 9 asserts
+Merge (18), (19), and notice that Lemma 10 asserts
 $\sum_{j=1}^{m}\Pr_{\nu}\Big(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Big)\geq\frac{1}{2},$
@@
 $\begin{array}[]{cl}\min&\frac{1}{800}\max\left\{\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j},\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}\right\}\\ s.t.&\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}\\ &0\leq p_{j}\leq 1\end{array}$
-Then, we turn to show $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$ holds for $j\in[m],a\in[K]$.
+where for any symmetric and $\delta$-PAC algorithm, $\Big\{p_{j}=\Pr_{\nu}\Big(\hat{a}=j,N_{j}(\tau)>\frac{\log\frac{1}{\delta}}{100\Delta_{j,0}^{2}}\Big)\Big\}_{j=1}^{m}$ is always a feasible solution.
+The next step is to show $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{4\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$ holds for $j\in[m],a\in[K]$.
 If $\mu_{a}\geq\mu_{j}$, we have $\frac{1}{\max\{\Delta_{j,0}^{2},\Delta_{a,j}^{2}\}}\geq\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}$.
@@
 $\begin{array}[]{cl}\min&\frac{1}{3200}\max\left\{\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j},\max_{1\leq j\leq m}\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)}\right\}\\ s.t.&\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}\\ &0\leq p_{j}\leq 1\end{array}$
-Reformulate the above optimization problem, we can turn to solve
+Reformulate the above optimization problem, we know the optimal value of (2)
 $\begin{array}[]{cl}\min&v\\ s.t.&v\geq\sum_{j=1}^{m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}p_{j}\\ &v\geq\sum_{a=1}^{K}\frac{p_{j}}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}\left(1+\log\frac{1}{p_{j}}\right)},\forall j\in[m]\\ &\sum_{j=1}^{m}p_{j}\geq\frac{1}{2}\\ &0\leq p_{j}\leq 1\\ &v\geq 0.\end{array}$
+is a lower bound of $\mathbb{E}\tau$. ∎
 Given Lemma 6, the last step is to prove a lower bound of the optimal value of (2).
@@
 $\text{opt}\geq\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right)$
-Proof.
-Proof of Lemma 12
+Proof of Lemma 13.
 We first show function $f(p)=\frac{p}{1+\log\frac{1}{p}}$ is convex for $p\in(0,1]$. Easy to calculate
@@
 $\min_{\{p_{j}\}\in[0,1]^{m},v\geq 0}L(\alpha^{0},\gamma^{0},\{\beta_{j}^{0}\}_{j=1}^{m};\{p_{j}\}_{j=1}^{m},v)$
-$\geq$ $\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$
-Before that, we prove a claim
+$\geq$ $\Omega\left(\min_{1\leq j\leq m}\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\frac{1}{j}}{\log^{2}(m+1)}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\right).$ (20)
+We start the proof of (20) by claiming
 $\min_{1\geq p_{j}\geq 0}\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p_{j}+\frac{p_{j}}{1+\log\frac{2}{p_{j}}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p_{j}\right)\geq-\frac{\gamma^{0}}{100m}$ (21)
+holds for all $j\in[m]$.
 Define $\bar{L}_{j}(p)=\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p+\frac{p}{1+\log\frac{2}{p}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p$ and $p^{*}_{j}=\arg\min\limits_{1\geq p\geq 0}\bar{L}_{j}(p)$.
-Not hard to validate $\bar{L}_{j}(p)$ is a convex function.
-We prove the claim by discussing two cases.
+Not hard to validate $\bar{L}_{j}(p)$ is a convex function for $p\in[0,1]$, and there are two cases.
 If $\frac{\partial\bar{L}_{j}(p)}{\partial p}\geq 0$ holds for all $0\leq p\leq 1$, we have
@@
 The last line is by the fact that $2\sum_{s=1}^{m}\frac{1}{2s}<1+\log(m+1)$ for $m\geq 1$.
-This inequality contradicts with the fact that $\bar{L}_{j}(p^{*}_{j})<0$.
-Thus, we can conclude
+This last line contradicts with the fact that $\bar{L}_{j}(p^{*}_{j})<0$.
+Thus, we can conclude $p^{*}_{j}<\frac{1}{100m}$, further
 $\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}\alpha^{0}\cdot p^{*}_{j}+\frac{p^{*}_{j}}{1+\log\frac{2}{p^{*}_{j}}}\sum_{a=1}^{K}\frac{1}{\max\{\Delta_{a,0}^{2},\Delta_{a,j}^{2}\}}\cdot\beta_{j}^{0}-\gamma^{0}p^{*}_{j}$
@@
 C.1 Rigorous Presentation of Algorithm
+Introduce
 $U(t,\delta)=\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}.$ (22)
-where $\lceil x\rceil:=\max\{x,1\}$
+where $\lceil x\rceil:=\max\{x,1\}$.
 Before presenting the full statement of the algorithm, we first introduce some notations
@@
 $\text{LCB}^{\text{et}}_{a}(\mathcal{H}^{et},\delta)=\hat{\mu}_{a}^{\text{et}}(\mathcal{H}^{et})-U(N_{a}^{\text{et}}(\mathcal{H}^{ee}),\frac{\delta}{K})$
-where we consider $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$ as stacks that can add index to the stored item, and denote $A_{s}(\mathcal{H}^{\text{ee}}),X_{s}(\mathcal{H}^{\text{ee}})$ as the action and reward stored with index $s$ in $\mathcal{H}^{\text{ee}}$.
-Same idea applies to $A_{s}(\mathcal{H}^{\text{et}}),X_{s}(\mathcal{H}^{\text{et}})$.
-If $N_{a}^{\text{ee}}(\mathcal{H}^{ee})=0$ holds for some $a\in[K]$, we take $\text{UCB}^{\text{ee}}_{a}=+\infty,\text{LCB}^{\text{ee}}_{a}=$-$\infty$.
-Same idea applies to $N_{a}^{\text{et}}(\mathcal{H}^{et}),\text{UCB}^{\text{et}}_{a},\text{LCB}^{\text{et}}_{a}$.
-Given the above notation, we introduce algorithm 2, which is adapted from Li and Cheung (2025).
-The following analysis is very similar to the Section 4 and Appendix B in Li and Cheung (2025), but we still remain the full description and proof for completeness.
-In the following subsection, we discuss each part of the algorithm line by line and highlight our adaptation of history literature in Appendix C.1.5.
+We treat $\mathcal{H}^{\mathrm{ee}}$ and $\mathcal{H}^{\mathrm{et}}$ as stacks with indexed entries: $A_{s}(\mathcal{H}^{\mathrm{ee}})$ and $X_{s}(\mathcal{H}^{\mathrm{ee}})$ denote, respectively, the action and reward stored at index $s$ in $\mathcal{H}^{\mathrm{ee}}$, and analogously for $\mathcal{H}^{\mathrm{et}}$.
+When $N_{a}^{\mathrm{ee}}(\mathcal{H}^{\mathrm{ee}})=0$ for some $a\in[K]$, we set $\mathrm{UCB}^{\mathrm{ee}}_{a}=+\infty$ and $\mathrm{LCB}^{\mathrm{ee}}_{a}=$-$\infty$; the same convention applies to $\mathrm{UCB}^{\mathrm{et}}_{a}$ and $\mathrm{LCB}^{\mathrm{et}}_{a}$ when $N_{a}^{\mathrm{et}}(\mathcal{H}^{\mathrm{et}})=0$.
+With this notation in place, we present Algorithm 2, adapted from Li and Cheung [2025].
+Our analysis follows closely Section 4 and Appendix B of Li and Cheung [2025]; nevertheless, we provide the full algorithmic description and proofs for completeness.
+In the subsequent subsections, we discuss the algorithm line by line and summarize our modifications relative to the prior literature in Appendix C.1.5.
 1Input: Bracket $B$, Threshold $\mu_{0}$, Tolerance level $\delta$, Tunable parameter $C>1$, Arm number $K$;
@@
 Algorithm 2 Sequential Exploration Exploitaion (SEE)
-C.1.1 From Line 2 to 2: Input and Initialized Parameter
-Algorithm 2 takes bracket $B$, threshold $\mu_{0}$, tolerance level $\delta$, arm number $K$, tunable parameter $C$ as input, and only pull arms in $B$.
-$\mu_{0},K$ are the same as what Algorithm 1 receives, while Algorithm 1 sets $\delta\leftarrow\frac{\delta}{\lceil\log_{2}K\rceil+1}$ when calling Algorithm 2.
-We take $C=1.01$ by default, but we can use any value greater than $1$.
-Besides input parameter specified by the problem nature and Algorithm 1, 2 also initialize some parameters to manage the pulling process.
+C.1.1 From Line 2 to 2: Input and Initialize Parameter
+Algorithm 2 takes a bracket $B$, a threshold $\mu_{0}$, a tolerance level $\delta$, the number of arms $K$, tunable parameter $C$ as input, and only pull arms in $B$.
+The parameters $\mu_{0}$ and $K$ coincide with those used by Algorithm 1.
+When Algorithm 1 calls Algorithm 2, it sets $\delta\leftarrow\delta/(\lceil\log_{2}K\rceil+1)$.
+We use $C=1.01$ by default, although any value strictly larger than $1$ is admissible.
+In addition to the problem parameters, 2 also initialize some parameters to manage the pulling process.
 Parameter $\{\delta_{k}=\frac{1}{3^{k}},\beta_{k}=2^{k},\alpha_{k}=5^{k}\}_{k=1}^{+\infty}$ are used to set up the tolerance level and budget during the pulling process.
-It is possible to specify other values for $\delta_{k},\beta_{k},\alpha_{k}$, see Appendix B.1 in Li and Cheung (2025).
-$\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}},Q$ are used to store reward collected from pulling arms.
-As we have discussed above, $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$ are two stacks that can maintain the arrival order of reward of each arm $a\in B$.
-This requirement is to guarantee we are able to transfer the latest collected reward out of $\mathcal{H}^{\text{et}},\mathcal{H}^{\text{ee}}$.
-$Q$ is used to guarantee $\text{LCB}_{a}^{\text{ee}}$ cannot be greater than $\mu_{0}$ at the start of each phase, which will be discussed in section C.1.4.
+It is possible to specify other values for $\delta_{k},\beta_{k},\alpha_{k}$, see Appendix B.1 in Li and Cheung [2025].
+The containers $\mathcal{H}^{\mathrm{ee}}$, $\mathcal{H}^{\mathrm{et}}$, and $Q$ store samples collected throughout the process.
+In particular, $\mathcal{H}^{\mathrm{ee}}$ and $\mathcal{H}^{\mathrm{et}}$ are ordered stacks that preserve the arrival order of samples for each arm in $B$, which allows us to remove and transfer the most recently collected sample when needed.
+The role of $Q$ is to ensure that $\mathrm{LCB}^{\mathrm{ee}}_{a}$ cannot be greater than $\mu_{0}$ at the beginning of each phase; see Section C.1.4.
 C.1.2 Line 2 to Line 2: Overall Idea of Algorithm 2
-At the start of Algorithm 2, it splits the arm distributions into two parallel and independent copies, one for exploration period, and another for exploitation period.
+At the start of Algorithm 2, it splits the arm distributions into two parallel and independent copies.
+One copy is for exploration period, and another copy is for exploitation period.
 Correspondingly, reward collected in the exploration periods are stored in $\mathcal{H}^{\text{ee}}$.
@@
 In each phase $k$, Algorithm 2 consists of two parts.
-The first part is the conditions from Line 2 to Line 2.
-The second part, which is called Exploration Period, ranges from Line 2 to 2.
-Exploration Period adopts UCB rule in Jamieson and Nowak (2014) to pull arms.
-These two parts are sequentially executed in the While Loop at Line 2.
-Though the condition of the While Loop is always True, but the conditions in the first part is able to guarantee that the Algorithm 2 never gets stuck in a infinity loop.
-See the discussion in section C.1.3.
-At a phase $k$, the exploration period utilizes $\delta_{k}$ as the tolerance level to apply UCB rule of pulling arms, while the exploitation period utilizes $\frac{\delta}{\alpha_{k}}$ as the tolerance level to calculate Upper and Lower Confidence Bounds and only pull a single arm.
-Once the tolerance level of both periods are below $\frac{\delta}{3}$ and $\hat{a}$ takes value in $K]\cup\{\textsf{None}\}$, Algorithm 2 will terminate.
-What’s more, Algorithm 2 also sets up a maximum pulling times for both periods.
-For the exploration period at phase $k$, the maximum pulling times for each arm $a\in B$ is $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$.
-For the exploitation period at phase $k$, the maximum pulling times of the single arm is $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$.
-Section C.1.3 and C.1.4 provide more details.
+Each phase consists of two components executed within the while-loop in Line 2:
+  • a sequence of stopping/transition checks (Lines 22), and
+  • an exploration period (Lines 22) that applies a UCB rule (as in Jamieson and Nowak [2014]) to select the next arm to sample.
+Although the loop condition is True, the stopping/transition checks ensure that the algorithm cannot run indefinitely; see Section C.1.3.
+Within phase $k$, exploration uses tolerance level $\delta_{k}$ to form UCB/LCB bounds, while exploitation uses tolerance level $\delta/\alpha_{k}$ and samples only a single candidate arm.
+Once $\hat{a}$ takes value in $K]\cup\{\textsf{None}\}$, Algorithm 2 will terminate.
+Moreover, each phase enforces explicit sampling budgets: during exploration, each arm $a\in B$ can be sampled at most $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ times, while during exploitation the candidate arm is sampled at most $\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}$ times.
+Further details appear in Sections C.1.3 and C.1.4.
 C.1.3 Line 2 to 2: Conditions on Phase Transition
@@
 Usage of $Q$ is to guarantee Lemma 23, requiring $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$.
-Similar to the reason described in the section 4.1 of Li and Cheung (2025), Lemma 22 is required to prove Lemma 23, further Theorem 3.
+Similar to the reason described in the section 4.1 of Li and Cheung [2025], Lemma 23 is required to prove Lemma 24, further Theorem 3.
 We put more discussion of $Q$ to section C.1.4.
@@
 The third condition is Line 2, which is correlated to the budget of the exploration period.
-For a phase $k$, Algorithm 2 sets up a maximum pulling times $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ for each arm $a\in B$.
-Recall we only consider picking $A^{\text{ee}}_{t}$ from those arms whose total pulling times from phase 1 to phase $k$ is below $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ in Line 2.
-The third condition is Line 2 is to enter the exploration period in phase $k+1$, if pulling times are no less than $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}-1$ holds for all $a\in B$.
-Line 2 and 2 also guarantee Algorithm 2 will not get stuck into a infinite loop of “While True” in Line 2.
+In phase $k$, each arm $a\in B$ is assigned an exploration budget of $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$.
+We only consider pull $A^{\text{ee}}_{t}$ from those arms whose total pulling times from phase 1 to phase $k$ is below $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}$ in Line 2.
+If all arms in $B$ have reached their budget (up to the 1-1 offset in the implementation), the algorithm sets $\hat{a}_{k}=\textsf{Not Completed}$ and transitions to phase $k+1$.
+Together, Lines 2 and 2 ensure that the while-loop in Line 2 cannot become an infinite loop.
 C.1.4 Line 2 to 2: Exploration
@@
 If not, Algorithm 2 samples a new random variable from $\nu_{A_{t}^{\text{ee}}}$.
-As we mentioned in section C.1.3, usage of $Q$ is to guarantee Lemma 22, requiring $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$.
-Someone may question whether we can directly abandon the latest collected sample of $A_{t}^{\text{ee}}$ in $\mathcal{H}$, which seems also able to guarantee $\text{LCB}_{a}^{\text{ee}}<\mu_{0}$ holds for all $a\in B$ at the start of phase $k$.
-We cannot to do so.
-By transferring samples into or out $Q$, we are able to derive
+As noted above, the purpose of $Q$ is to enforce Lemma 23, which requires $\mathrm{LCB}^{\mathrm{ee}}_{a}<\mu_{0}$ for all $a\in B$ at the start of each phase.
+One might ask whether it suffices to simply discard the most recent exploration sample instead of storing it in $Q$.
+This is not sufficient.
+The transfer mechanism preserves the per-arm arrival order of samples, which is needed for the following property.
 Lemma 14.
@@
 Condition on reward sequence in the exploration period $\{X_{a,s}^{\text{ee}}\}_{s=1}^{+\infty}$, for each $a\in B$, we can assert $\hat{\mu}_{a}(\mathcal{H}^{\text{ee}})\in\{\frac{\sum_{s=1}^{t}X_{a,s}^{\text{ee}}}{t}\}_{t=1}^{+\infty}$ throughout the pulling process.
-The reason is we do not change the arrival order of $\{X_{a,s}^{\text{ee}}\}_{s=1}^{+\infty}$ for each $a\in B$. To analyze Algorithm 2, we need to introduce
-$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\right\}$
-$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\right\}.$
-where $\lceil x\rceil^{+}=\max\{x,1\}$.
+The key point is that the algorithm never permutes the order of $\{X_{a,s}^{\mathrm{ee}}\}_{s\geq 1}$ for any arm $a\in B$; it only temporarily holds out the most recent sample and later reinserts it without changing order.
+For the analysis, we introduce the random phase indices
+$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<U(t,\frac{\delta_{k}}{K})\right\}$
+$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<U(t,\frac{\delta_{k}}{K})\right\}.$
 $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately.
@@
 Lemma 14 suggests that for phase $k\geq\kappa^{\text{ee}}_{b}$, we can guarantee $\text{LCB}_{a}^{\text{ee}}<\mu_{a}<\text{UCB}_{a}^{\text{ee}}$ holds for all $a\in B_{b}$, which further helps to prove upper bounds of the pulling times for all $a\in B_{b}$.
-Simply dropping the latest collected samples are unable to guarantee 13, making it much harder to conduct the analysis.
+Simply dropping the latest collected samples is unable to guarantee 14, making it much harder to conduct the analysis.
 Besides, we also care about the size of $Q$ during the pulling process, as we can observe the following Lemmas.
@@
 In this case, $A_{t}^{\text{ee}}$ just get pulled in the last round of exploration.
-Meanwhile, once $Q$ contains a sample for an arm $a\in B$, condition in 2 guarantees Algorithm 2 will never collect new samples of arm $a$ in the exploration period.
-This property implies following.
+Meanwhile, once $Q$ contains a sample for an arm $a\in B$, condition in Line 2 ensures that Algorithm 2 will not draw a fresh exploration sample for arm $a$ until that held-out sample is transferred back.
+This yields the following bound.
 Lemma 16.
 Consider an algorithm copy alg taking bracket $B$ as input, we have $|Q|\leq|B|$ throughout the pulling process
-At the end of this subsection, we present a Lemma similar to 13, but for the exploitation period.
+At the end of this subsection, Lemma 17 is similar to 14, but for the exploitation period.
 Lemma 17.
@@
 C.1.5 Adaptation of History Literature
-Compared to the Algorithm SEE in Li and Cheung (2025), our main adaptation consists of two points.
-The first adaptation is to put the stopping conditions (Line 2 to 2) before the exploration periods (Line 2 to 2).
-This design is to guarantee Lemma 29, guaranteeing Algorithm 2 will not pull arm $a$ whose $\text{UCB}_{a}^{\text{ee}}$ is below $\mu_{0}$.
-The second point is for each phase $k$, we set up a maximum pulling times of each arm in the exploration period, instead of maximum pulling times of the total pulling times (reference line index is to be added).
-The reason is the arm in a bracket $B$ is randomized, it is hard to analyze the required pulling times that a bracket need to output a $\hat{a}_{k}$.
+Compared to the Algorithm SEE in Li and Cheung [2025], our main adaptation consists of two points.
+The first adaptation is we place the stopping/transition conditions (Lines 22) before the exploration step (Lines 22).
+This ordering is required for Lemma 30, which ensures that the algorithm does not sample any arm $a$ whose exploration upper confidence bound satisfies $\mathrm{UCB}^{\mathrm{ee}}_{a}\leq\mu_{0}$.
+The second point is for each phase $k$, we set up a maximum pulling times of each arm in the exploration period, rather than a budget on the total number of exploration pulls across the bracket (Line 2 and 2).
+This modification is motivated by the randomization of arms within each bracket: with randomized arm assignments, it becomes difficult to track how a total budget is required for the bracket to produce a candidate $\hat{a}_{k}$.
 C.2 Properties of Algorithm 2
-This subsections follows the Appendix B in Li and Cheung (2025), which are nearly the same. Here we remain the full proof for completeness, as we have modified some implementation details and proof.
+This subsection closely follows Appendix B of Li and Cheung [2025], with only minor differences.
+Since we modify several implementation details (and correspondingly adjust parts of the argument), we include the full proofs for completeness.
 In Algorithm 1, we create $\lceil\log_{2}K\rceil+1$ independent copies of Algorithm 2.
@@
 We define
-$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K(\lceil\log_{2}t\rceil^{+})^{2}}{\delta_{k}}}}{t}\right\}$
-$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<\frac{\sqrt{2\cdot 2^{\lceil\log_{2}t\rceil^{+}}\log\frac{2K\alpha_{k}(\lceil\log_{2}t\rceil^{+})^{2}}{\delta}}}{t}\right\}.$
-where $\lceil x\rceil^{+}=\max\{x,1\}$.
+$\kappa^{\text{ee}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{ee}_{a,b}(t)-r^{\nu}_{a}|<U(t,\frac{\delta_{k}}{K})\right\}$
+$\kappa^{\text{et}}_{b}=\min\left\{k:\forall t,\forall a\in[K],|\hat{\mu}^{et}_{a,b}(t)-r^{\nu}_{a}|<U(t,\frac{\delta_{k}}{K})\right\}.$
 $\hat{\mu}^{ee}_{a,b}(t),\hat{\mu}^{et}_{a,b}(t)$ denote the empirical mean of arm $a$ of the first $t$ collected samples in bracket $B_{b}$, corresponding to the exploration and exploitation periods separately.
@@
 $=$ $(C+1)^{2}\beta_{k}\log\frac{4K}{\delta_{k}}.$
-The first step is still caused by the Line 2 in Algorithm 2. The second step is from the foundation. The last line shows that the induction holds. ∎
+The first step is still caused by the Line 2 in Algorithm 2. The second step is from the conclusion at phase index $k$. The last line shows that the induction holds. ∎
 Lemma 22.
@@
 $\tau_{k}^{\text{et}}(B)-\tau_{k-1}^{\text{et}}(B)$
-$=$ $\sum_{a=1}^{K}N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k}^{\text{et}}(B)\Big)-N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k-1}^{\text{et}}(B)\Big)\mathds{1}(\hat{a}_{k}=a)$
+$=$ $\sum_{a=1}^{K}\left(N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k}^{\text{et}}(B)\Big)-N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k-1}^{\text{et}}(B)\Big)\right)\mathds{1}(\hat{a}_{k}=a)$
 $\leq$ $\sum_{a=1}^{K}N_{\hat{a}_{k}}^{\text{et}}\Big(\tau_{k}^{\text{et}}(B)\Big)\mathds{1}(\hat{a}_{k}=a)$
@@
 Proof of Lemma 23.
-We apply induction on $k$. For $k=1$, this is obviously true, as we denote the $\text{LCB}_{a}^{\text{ee}}$ as -$\infty$ if the pulling times of arm $a$ is 0.
+We prove the Lemma by induction on $k$. For $k=1$, this is obviously true, as we denote the $\text{LCB}_{a}^{\text{ee}}$ as -$\infty$ if the pulling times of arm $a$ is 0.
 If the conclusion holds for the phase k1k-1, we prove the case of $k$ by discussing all the possible values of $\hat{a}_{k-1}$.
-Before the start of phase $k$, there are three possible scenarios for the value of $\hat{a}_{k-1}$.
-Phase k1k-1 may end with taking $\hat{a}_{k-1}$ takes value at the Line 2, 2 or 2.
+Since the pulling process enter phase $k$, we can conclude phase k1k-1 may end with taking $\hat{a}_{k-1}$ at the Line 2, 2 or 2.
 If $\hat{a}_{k-1}$ takes Not Completed at the Line 2, we have we $\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k})<\text{LCB}^{\text{ee}}_{a}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0}$, forall $a\in[K]$.
 If $\hat{a}_{k-1}$ takes Not Completed at the Line 2, We can also assert $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})<\mu_{0},\forall a\in[K]$.
-Prove by contradiction.
+We argue by contradiction.
 Assume there exists $a\in[K]$ such that $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k-1})\geq\mu_{0}$ at the end of phase k1k-1.
@@
 Once there exists an arm $a$ whose LCB is above $\mu_{0}$ after an arm pull, the algorithm will enter Line 2, suggesting that there is at most one arm whose LCB is above $\mu_{0}$.
+This observation guarantees that $\text{LCB}_{a}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$.
 Then, we turn to analyze $\text{LCB}_{\hat{a}_{k-1}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})$ at the start of phase $k$. Before the execution of Line 2, the following inequalities must hold
@@
 Since $U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k}}{K}\right)>U\left(N_{\hat{a}_{k-1}}^{\text{ee}}-1,\frac{\delta_{k-1}}{K}\right)$, we know before the start of phase $k$, $\text{LCB}_{\hat{a}_{k-1}}^{\text{ee}}(\mathcal{H}^{\text{ee}},\delta_{k})<\mu_{0}$ must hold.
+∎
 Following lemma suggests that the exploration period is able to output correct answer, given some concentration event.
@@
 The remaining work is to prove if $\hat{a}_{k}\in[K]$, we have $\mu_{\hat{a}_{k}}\geq\omega\mu_{j}+(1-\omega)\mu_{0}$, $\omega=\frac{C-1}{C+3}$.
-For simplicity, we denote $\tau_{k}^{\text{ee}}$ as $\tau_{k}^{\text{ee}}(B)$, and prove by contradiction.
-If $\hat{a}\in[K]$ and $\mu_{\hat{a}}<\omega\mu_{j}+(1-\omega)\mu_{0}<\omega\mu_{a^{*}(B)}+(1-\omega)\mu_{0}$, we have
-$A_{\tau^{\text{ee}}_{k}}^{\text{ee}}=\hat{a}_{k},\hat{a}_{k}=\arg\max_{i\in B}\hat{\mu}_{i,N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)}+U(N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})$
-$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}$
-$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1,\frac{\delta_{k}}{K})<\mu_{0}.$
+For simplicity, we denote $\tau_{k}^{\text{ee}}$ as $\tau_{k}^{\text{ee}}(B)$.
+From the outputting rule, we have
+$\begin{split}&A_{\tau^{\text{ee}}_{k}}^{\text{ee}}=\hat{a}_{k},\hat{a}_{k}=\arg\max_{i\in B}\hat{\mu}_{i,N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)}+U(N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\\ &\hat{\mu}_{\hat{a}_{k},N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}\\ &\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1,\frac{\delta_{k}}{K})<\mu_{0}.\end{split}$ (25)
 where $\hat{\mu}_{i,N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)}$ denotes the empirical mean value of arm i of the first $N_{i}^{\text{ee}}(\tau^{\text{ee}}_{k}-1)$ samples.
 The third line is by the Lemma 23, suggesting $\text{LCB}_{\hat{a}_{k}}^{\text{ee}}<\mu_{0}$ must hold before the latest collection of the sample.
+We consider two cases, $N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})=1$ or $N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})\geq 2$. If $N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})=1$, we have
+$\mu_{0}\leq\hat{\mu}_{\hat{a}_{k},1}-C\cdot U(1,\frac{\delta_{k}}{K})$
+$\Rightarrow$ $\mu_{0}\leq\mu_{\hat{a}_{k}}-(C-1)\cdot U(1,\frac{\delta_{k}}{K})$
+$\Rightarrow$ $\mu_{0}\leq\mu_{\hat{a}_{k}}-2(C-1)$
+$\Leftrightarrow$ $\mu_{0}+2(C-1)\leq\mu_{\hat{a}_{k}}.$
+The second last line is from the definition of (22), suggesting that $U(1,\frac{\delta_{k}}{K})\geq 2$.
+From the assumption $\{\mu_{a}\}_{a=1}^{K}\in[0,1]^{K}$ and $C+1$, we know $\omega\mu_{j}+(1-\omega)\mu_{0}=\frac{C-1}{C+3}\mu_{j}+(1-\frac{C-1}{C+3})\mu_{0}<\frac{C-1}{4}+\mu_{0}<\mu_{0}+2(C-1)$.
+We complete the proof.
+If $N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})\geq 2$, we argue by contradiction.
+Assume $\mu_{\hat{a}_{k}}<\omega\mu_{j}+(1-\omega)\mu_{0}$.
 As we take $U(t,\frac{\delta}{K})=\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}t\rceil,1\}}\log\frac{2K(\lceil\log_{2}t\rceil)^{2}}{\delta}}}{t}$, we get
@@
 $\Leftrightarrow$ $\frac{2\omega(\mu_{a^{*}(B)}-\mu_{0})}{C-1}>\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}.$
-By $\omega=\frac{C-1}{C+3}$, we have $\frac{(1-\omega)}{2}=\frac{1-\frac{C-1}{C+3}}{2}=\frac{4}{2(C+3)}=\frac{2\omega}{C-1}$, which means
-$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}>\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}.$
+By $\omega=\frac{C-1}{C+3}$, we have $\frac{(1-\omega)}{2}=\frac{1-\frac{C-1}{C+3}}{2}=\frac{4}{2(C+3)}=\frac{2\omega}{C-1}$, we have
+$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}>\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})}.$ (26)
 Notice that
@@
 $\leq$ $\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})},$
-we can conclude
-$\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}\geq\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}.$
+Together with (26), we can conclude
+$\hat{\mu}_{\hat{a}_{k},N_{\hat{a}}^{\text{ee}}(\tau^{\text{ee}}_{k})}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}$
+$\Rightarrow$ $\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}\geq\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}.$ (27)
 On the other hands,
@@
 […]
-$\Leftarrow$ $\omega\mu_{1}+(1-\omega)\mu_{0}+2U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
+$\Leftarrow$ $\omega\mu_{a^{*}(B)}+(1-\omega)\mu_{0}+2U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}-1),\frac{\delta_{k}}{K})\leq\mu_{a^{*}(B)}$
 $\Leftrightarrow$ $\frac{2\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\leq(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})$
-$\Leftrightarrow$ $\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\leq\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}.$
-In summary, we can conclude
+$\Leftrightarrow$ $\frac{\sqrt{2\cdot 2^{\max\{\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil,1\}}\log\frac{2K(\lceil\log_{2}N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1\rceil)^{2}}{\delta}}}{N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k})-1}\leq\frac{(1-\omega)(\mu_{a^{*}(B)}-\mu_{0})}{2}.$ (28)
+Combining (27) and (28), we can conclude
 $\hat{\mu}_{\hat{a}_{k},N_{\hat{a}_{k}}^{\text{ee}}(\tau)}-C\cdot U(N_{\hat{a}_{k}}^{\text{ee}}(\tau^{\text{ee}}_{k}),\frac{\delta_{k}}{K})>\mu_{0}$
@@
 $\Rightarrow$ $A_{\tau^{\text{ee}}_{k}}^{\text{ee}}\neq\hat{a}_{k}.$
-The last line suggests a contradiction. We can conclude $\mu_{\hat{a}}\geq\frac{C-1}{C+3}\mu_{j}+\frac{4}{C+3}\mu_{0}$. ∎
+[…]
 Lemma 25.
@@
 $\leq$ $\max\left\{N_{a}^{\text{ee},B}\Big(\tau_{k-1}^{\text{ee}}(B)\Big),\frac{29(C+1)^{2}\left(\log\frac{2K}{\delta_{k}}+\log\log\frac{24(C+1)^{2}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right)}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\right\}$
-hols for for any phase index $k\geq\kappa^{\text{ee}}_{B}$. If the above conclusion holds, we can use induction to conclude
+holds for for any phase index $k\geq\kappa^{\text{ee}}_{B}$. If (29) holds, we can use induction to conclude
 $N_{a}^{\text{ee},B}\Big(\tau_{k}^{\text{ee}}(B)\Big)$
@@
 Proof of Lemma 28.
-This is a direct conclusion from Lemma 25, 26. ∎
+This is a direct conclusion from Lemma 26, 27. ∎
 Lemma 29.
@@
 ∎
-C.3 Proof of Theorem 2 and 3
-The full proof requires multiple lemmas to support, which motivates us to use figure 1 to summarize the overall ideas.
-An arrow from A to B means Lemma A is used in the proof of Lemma/Theorem B.
-All the Lemmas inside the orange box are the properties of Algorithm 2, which are mainly adapted from Li and Cheung (2025).
-Figure 1: Overall Idea of Full Proof of Theorem 2 and 3
-Before illustrating the proof of Theorem 2 and 3, we first introduce a Lemma for the random permutation, Line 1 of Algorithm 1.
-Lemma 31.
+C.3 Sketch Proof of Theorem 2, 3
+The full proof of Theorem 2, 3 requires multiple lemmas to support.
+Sketch Proof of Theorem 2.
+The first step is prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)=0$, whose main idea is very similar to the proof of Theorem 5.2 in Li and Cheung [2025].
+As in Li and Cheung [2025], denote $\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b}$ as the minimum phase indices after which the exploration and exploitation concentration events of copy $b$ hold, we can prove $\Pr(\tau(B_{\lceil\log_{2}K\rceil+1})=+\infty)\leq\Pr(\max\{\kappa^{\text{ee}}_{\lceil\log_{2}K\rceil+1},\kappa^{\text{et}}_{\lceil\log_{2}K\rceil+1}\}=+\infty)=0$.
+Given the first step, we are able to conclude $\Pr(\tau<+\infty)=1$, by the fact that $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$.
+Then, to prove Algorithm 1 is $\delta$-PAC, we suffice to show $\Pr_{\nu}\left(\hat{a}=\textsf{None}\text{ or }\{\hat{a}\in[K],\mu_{\hat{a}}<\mu_{0}\}\right)<\delta$ for a positive instance $\nu$, and $\Pr_{\nu}\left(\hat{a}\in[K]\right)<\delta$ for a negative instance $\nu$.
+The two cases are proved similarly.
+In Line 1 of Algorithm 1, we assign $\delta/(\lceil\log_{2}K\rceil+1)$ to each copy.
+We are able to prove $\Pr(\text{alg}_{b}\text{ outputs incorrect }\hat{a})<\frac{O(1)\delta}{\lceil\log_{2}K\rceil+1}$ holds for all $b\in[\lceil\log_{2}K\rceil+1]$.
+A union bound over all copies completes the proof.
+∎
+Sketch Proof of Theorem 3.
+We consider negative and positive instances separately.
+If the instance $\nu$ is negative, we focus on analyzing $\tau(B_{\lceil\log_{2}K\rceil+1})$, which is very similar to the negative case of Theorem 5.3 in Li and Cheung [2025].
+By the fact that $\mathbb{E}\tau\leq[\lceil\log_{2}K\rceil+1]\mathbb{E}\tau(B_{\lceil\log_{2}K\rceil+1})$, we complete the proof.
+If the instance $\nu$ is positive, denote $\mu_{1}\geq\cdots\geq\mu_{m}>\mu_{0}\geq\mu_{m+1}\geq\cdots\geq\mu_{K}$.
+For $j\in[m]$ and $a\in[K]$, define $\tilde{T}_{j}(a)=\frac{\log(4K)+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$, $\tilde{T}_{j}(a;k)=\frac{\log\frac{4K}{\delta_{k}}+\log_{2}\frac{1}{\Delta_{j,0}^{2}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}$.
+We show that for every $j\in[m]$, $\mathbb{E}\tau\leq O(\log K)\cdot O\left(\frac{\log\frac{1}{\delta}}{\Delta_{j,0}^{2}}+\frac{\log^{3}K}{j}\sum_{a=1}^{K}\tilde{T}_{j}(a)\right)$.
+In addition, the constant hidden by the Big O notation is independent of $j$.
+For any given $j\in[m]$, we introduce $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$, the smallest bracket index that contains at least one arm among the top-$j$ arms.
+As we adopt a random permutation to create brackets, $b_{j}$ is a random variable, and we have $\mathbb{E}\tau=\sum_{b=1}^{\lceil\log_{2}K\rceil+1}\mathbb{E}\tau\mathds{1}(b_{j}=b)$.
+Using $\tau\leq(\lceil\log_{2}K\rceil+1)\tau(B_{b})$ for each $b\in[\lceil\log_{2}K\rceil+1]$, it suffices to bound $O(\log K)\,\mathbb{E}_{\nu}[\tau(B_{b})\mathds{1}(b_{j}=b)]$.
+Let $\tau^{\text{ee}}(B_{b}),\tau^{\text{et}}(B_{b})$ denote the total length of exploration and exploitation periods of algorithm copy $\text{alg}_{b}$ respectively.
+For any arm $a\in B_{b}$, denote $N_{a}^{\text{ee},B_{b}}$ as the pulling times of arm $a$ in the exploration period.
+It is evident that $\tau(B_{b})=\tau^{\text{ee}}(B_{b})+\tau^{\text{et}}(B_{b})$.
+We assert that
+$\mathbb{E}\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)\Pr(b_{j}=b),$ (32)
+$\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)\leq\sum_{a=1}^{K}O(1)\tilde{T}_{j}(a)\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}+O(\log K)|B_{b}|\Pr(b_{j}=b).$ (33)
+Given (32)(33), we can sum up (32)(33) over $b=1,2,\cdots,\lceil\log_{2}K\rceil+1$.
+Lemma 7(section C.4), 33, 34( section D) provides upper bounds for $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}|B_{b}|\Pr(b_{j}=b)$ and $\sum\limits_{b=1}^{\lceil\log_{2}K\rceil+1}\min\{\frac{|B_{b}|}{K},\Pr(b_{j}=b)\}$, which complete the proof of Theorem 3.
+The remaining sketch proof reduces to (32)(33) for each $b\in[\lceil\log_{2}K\rceil+1]$.
+To prove (32), we first show the algorithm must terminate no later than the end of phase $L_{b}^{\text{pos}}=\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{O(1)(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}$, if $b_{j}=b$.
+Notice that $\beta_{k}\log\frac{4K(\lceil\log_{2}K\rceil+1)\alpha_{k}}{\delta}|_{k=\lceil\log_{2}\frac{O(1)(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil}=O\left(\frac{(C+3)^{2}}{(C-1)^{2}}\frac{\log\frac{4K}{\delta}+\log_{2}\frac{(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}}{\Delta_{j,0}^{2}}\right)$, and Lemma 18 shows $\Pr(\kappa^{\text{ee}}_{b}\geq k)\leq\frac{\pi^{2}\delta_{k-1}}{6}$, $\Pr(\kappa^{\text{et}}_{b}\geq k)\leq\frac{\pi^{2}\delta}{6\alpha_{k-1}}$, we can take expectation on both sides of $\tau^{\text{et}}(B_{b})\mathds{1}(b_{j}=b)\leq\frac{(C+3)^{2}}{(C-1)^{2}}\beta_{k}\log\frac{4K\alpha_{k}}{\delta}|_{k=L_{b}^{\text{pos}}}\mathds{1}(b_{j}=b)$ to complete the proof of (32).
+To prove (33), we first show for each $a\in B_{b}$, we have
+$N_{a}^{\text{ee},B_{b}}\mathds{1}(b_{j}=b)\leq O(\tilde{T}_{j}(a;L_{b}^{\text{pos}}))(C+1)^{2}\mathds{1}(b_{j}=b).$ (34)
+The intuition is as follows.
+Given the fact that $j\in B_{b}$, once the pulling times of arm $a$ is of order $O(1)\tilde{T}_{j}(a;L_{b}^{\text{pos}})(C+1)^{2}$, copy $\text{alg}_{b}$ can assert(with some probability, see Lemma 18, 25) either $\text{UCB}_{a}<\mu_{j}$ or $\text{LCB}_{a}>\mu_{0}$.
+In both cases, $\text{alg}_{b}$ stops pulling arm $a$ in the exploration period.
+By the facts that elements in $B_{b_{j}}$ are independent of $\kappa^{\text{ee}}_{b_{j}},\kappa^{\text{et}}_{b_{j}}$, we can sum up (34) over $a\in B_{b}$ and derive
+$\mathbb{E}\tau^{\text{ee}}(B_{b})\mathds{1}(b_{j}=b)\leq\sum_{a=1}^{K}O(1)\tilde{T}_{j}(a)\Pr(a\in B_{b},b_{j}=b)+O(\log K)|B_{b}|\Pr(b_{j}=b).$
+Since $\Pr(a\in B_{b},b_{j}=b)\leq\min\{\Pr(a\in B_{b}),\Pr(b_{j}=b)\}$, and $\Pr(a\in B_{b})=\frac{|B_{b}|}{K}$, (33) is shown. ∎
+C.4 Proof of Theorem 2 and 3
+In this subsection, We present the rigorous proof of Theorem 2 and 3 and the required lemmas.
+Before illustrating the proof of Theorem 2 and 3, we first prove Lemma 7 for the random permutation, Line 1 of Algorithm 1.
+Lemma (restatement of Lemma 7).
 Given a qualified arm index $j\in[m]$, denote bracket index $b_{j}=\min\{b\in[\lceil\log_{2}K\rceil+1]:B_{b}\cap[j]\neq\emptyset\}$. We have
@@
 $\Pr(b_{j}\geq\tilde{b})\leq\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{\tilde{b}-2}\rfloor}{K}\right)},\forall\tilde{b}\in[\lceil\log_{2}K\rceil+1].$
-Proof.
+Proof of Lemma 7.
 If $\tilde{b}=1$, the conclusion is $\Pr(b_{j}\geq 1)\leq 1$, which is obviously true. We focus on the case that $2\leq\tilde{b}\leq\lceil\log_{2}K\rceil+1$. We have
@@
 $\leq$ $\delta.$
-The second last line is by Lemma 18 and 19.
+The second last line is by Lemma 19 and 20.
 If $\nu$ is negative, we get
@@
 $\leq$ $O(\log K)\frac{1}{\exp\left(\frac{j\cdot\lfloor 2^{b-2}\rfloor}{K}\right)}\min\{2^{b-1},K\}.$
-The second line is by the fact that $\kappa^{\text{ee}}_{b}$ is independent with $b_{j}$. The third line is by the Lemma 17, 31. For the term $\ddagger$, we have
+The second line is by the fact that $\kappa^{\text{ee}}_{b}$ is independent with $b_{j}$. The third line is by the Lemma 18, 7. For the term $\ddagger$, we have
 $\mathbb{E}\left[\sum_{a=1}^{K}\frac{\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{b},\kappa^{\text{et}}_{b},\lceil\log_{2}\frac{448(C+3)^{2}}{(C-1)^{2}\Delta_{j,0}^{2}}\rceil\}}}+\log\log\frac{1}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}}{\max\{\Delta_{a,j}^{2},\Delta_{a,0}^{2}\}}\mathds{1}(b_{j}=b)\mathds{1}(a\in B_{b})\right]$
@@
 Similar to the proof of positive case, we split $\tau(B)=\tau^{\text{ee}}(B)+\tau^{\text{et}}(B)$ and prove an upper bound for $\mathbb{E}\tau^{\text{ee}}(B)$ and $\mathbb{E}\tau^{\text{et}}(B)$ separately.
-From Lemma 14, 29 and Lemma 30, we know
+From Lemma 15, 30 and Lemma 31, we know
 $\tau^{\text{ee}}(B)\leq$ $|Q|+\sum_{a=1}^{K}N_{a}^{\text{ee},B}\Big(\tau_{\kappa^{\text{ee}}_{B}-1}^{\text{ee}}(B)\Big)+\frac{O\left(\log\frac{4K}{\delta_{\max\{\kappa^{\text{ee}}_{B},\kappa^{\text{et}}_{B},\lceil\log_{2}\frac{226}{\Delta_{1,0}^{2}}\rceil,\lceil\log_{3}\frac{3}{\delta}\rceil\}}}+\log\log\frac{1}{\Delta_{a,0}^{2}}\right)}{\Delta_{a,0}^{2}}$
@@
 Notice that for $\lambda=\frac{z}{n\Delta^{2}}$, we have
-$\Pr\big(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}\geq s\big)$
-$=$ $\Pr\big(\max_{1\leq t\leq n}\lambda\big(\sum_{s=1}^{t}Z_{s}\big)\geq\lambda s\big)$
-$=$ $\Pr\left(\max_{1\leq t\leq n}\prod_{s=1}^{t}\exp\left(\lambda\big(Z_{s}\big)\right)\geq\exp\left(\lambda s\right)\right).$
+$\Pr\big(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}\geq z\big)$
+$=$ $\Pr\big(\max_{1\leq t\leq n}\lambda\big(\sum_{s=1}^{t}Z_{s}\big)\geq\lambda z\big)$
+$=$ $\Pr\left(\max_{1\leq t\leq n}\prod_{s=1}^{t}\exp\left(\lambda\big(Z_{s}\big)\right)\geq\exp\left(\lambda z\right)\right).$
 Easy to see $\{\prod_{s=1}^{t}\exp\left(\lambda\big(Z_{s}\big)\right)\}_{t=1}^{+\infty}$ is a submartingale, as
@@
 $\Pr(\max_{1\leq t\leq n}\sum_{s=1}^{t}Z_{s}>z)$
-$\leq$ $\frac{\mathbb{E}_{\nu^{\prime}}\prod_{t=1}^{n}\exp(\lambda(Z_{t}))}{\exp(\lambda s)}$
-$=$ $\frac{\exp(\frac{n\lambda^{2}\Delta^{2}}{2})}{\exp(\lambda s)}$
-$=$ $\exp(-\frac{s^{2}}{2n\Delta^{2}}).$
+$\leq$ $\frac{\mathbb{E}_{\nu^{\prime}}\prod_{t=1}^{n}\exp(\lambda(Z_{t}))}{\exp(\lambda z)}$
+$=$ $\frac{\exp(\frac{n\lambda^{2}\Delta^{2}}{2})}{\exp(\lambda z)}$
+$=$ $\exp(-\frac{z^{2}}{2n\Delta^{2}}).$
 Lemma 33.
@@
 $\leq$ $O\Big(\frac{\log j}{j}\Big).$
-Lemma 35 (Adapted Lemma 3 in Jamieson et al. (2014), Lemma D.3 in Li and Cheung (2025)).
+Lemma 35 (Adapted Lemma 3 in Jamieson et al. [2014], Lemma D.3 in Li and Cheung [2025]).
 Denote $\{X_{i}\}_{i=1}^{+\infty}$ as i.i.d $\sigma^{2}$-subgaussian random variable with true mean reward $\mu=0$. For any $\delta\in(0,1)$, we have
@@
     $\geq$ $\Pr\left(\exists t,|\sum_{s=1}^{t}X_{s}|\geq\sqrt{4\sigma^{2}t\log\frac{2(\log_{2}2t)^{2}}{\delta}}\right)$
-Lemma 36 (Lemma D.2 in Li and Cheung (2025)).
+Lemma 36 (Lemma D.2 in Li and Cheung [2025]).
 For any $\Delta\in(0,1],K\geq 2,\delta\in(0,\frac{1}{2}],C\geq 1$, we can conclude
@@
 $\Leftrightarrow$ $C\sqrt{\frac{4\log\frac{2K(\log_{2}2t)^{2}}{\delta}}{t}}<\Delta.$
+The last lemma is to solve inequality.
 Lemma 37.
@@
 The second inequality also implies $x\geq a\log\log(x)+b\Rightarrow x\geq b+a\log\log b$.
-Proof.
+Proof of Lemma 37.
 We prove the first claim.